A Complete Bibliography of Publications in Statistics and Probability Letters: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 December 2024 Version 1.38

Title word cross-reference

[YL22]. (2 + ι) [Sun20]. (p ∈ (0, 1)) [LXZ20]. (s^{2}) [SDH20]. +1 [RK23]. 0 [MZ21, Riv21]. 1 [MZ21, Riv21, WDG23, YS24]. 2 [MX22, ZZ24c, ZWW22]. 2HK ≤ 1 [Tal20]. 2 × 2 [BAB24]. 3 [FEVC21, PG20]. ^{⋆} [LHW24]. ^{X} [SP24a]. α [BKZ21, SXY22, SXX20, WW23, ZYS20, ZX20, ZH24]. β [LL21a]. c [Ses24]. S′ [Vil24]. D [LYC20, Shp24]. δ [BKB24, CE24, KS20c]. E [SB21]. ℓ_{1} [XX20a]. ϵ [Buo22]. f [GL21a]. G [He24, HD23b, MLZ23, PZ20, XC22b, Xu23, YY22, YWY23, ZJ21, ZHD24, JSWZ22, SWZ21]. H > 1 [Tal20]. H ∈ (0, 1/2) [MR24]. H → 0 [FFGS22]. I [Hah22]. k [Ahm20, Bas24, CD22, DS21, LL23a, NAV21, PG21, Wan22b]. κ = 8 [LMN24]. L [BKM20b, JB25]. L^{∞} [AJO20]. L^{p} [Cha24, Osȩ24]. L_{1} [GYZ22, SMS20]. L_{2} [HNP20, NP24, Vog20]. L_{F}^{∞}(μ) [SJK22]. L_{k} [Jur23]. L_{p} [ALO22, HT23a]. log GARCH [GHH24]. M [VY20, BS24, Gir24, QHK25, SD22b, XK24]. R [TT23]. C^{2} [EK22]. N [XZYC20, FV25, Fri22, LL23a, Vir21]. p [BTSL21, CK23, LXZ20, Vir21, XH22, Yu22]. Φ [HSZ21, Tan24]. Q [Myr21, HC23, MGMSP24, XK24]. R [HH20, HA24, PC24]. R^{n} [PG21]. s [LKA20, ZCL22]. σ [VM21]. SLE_{κ} [LMN24]. SO(3) [WL21a]. T [FFS24, Nov22, GH23, Tre20, WLS20]. T^{2} [AA21]. τ [SS25]. U [AKW22, CLP22, DP23, Gir24]. W_{2} [GW23]. Z^{d} [Tzi20].

-algebras [VM21]. -ARCH [LL21a]. -Brownian [HD23b, He24, XC22b, Xu23, YY22, ZJ21, ZHD24]. -Class [Bas24]. -contaminated [FFS24]. -convex [Ses24]. -death [ZZ24c]. -dependent [YL22]. -Dimensional [WDG23]. -distance [GYZ22]. -ergodicity [GL21a]. -expectation [Xu23]. -expectations [JSWZ22, SWZ21]. -functionals [HT23a]. -FWER [Wan22b]. -independent [Myr21]. -isotropic [HC23]. -level [LKA20, ZCL22]. -Lévy [MLZ23, YWY23]. -linear [Buo22]. -means [CD22]. -mixture [BKZ21]. -mixtures [BKB24]. -moments [JB25]. -norm [SMS20]. -normal [PZ20]. -norms [Cha24]. -optimal [HH20, HA24, LYC20, SB21, Shp24, SDH20]. -order [Yu22]. -orthogonal [QHK25]. -out-of- [LL23a]. -process [WLS20]. -quantiles [ALO22]. -records [Ahm20]. -rings [KS20c]. -scale [MGMSP24]. -shock [CE24]. -simplex [Fri22]. -small [NP24]. -stable [SXY22, WW23, ZYS20, ZH24, SXX20, ZX20]. -statistics [AKW22, BKM20b, CLP22, DP23, Gir24]. -subgaussian [Tan24]. -test [Nov22]. -th [BTSL21, BS24, LXZ20, PG21, SD22b, XK24, Sun20]. -type [MX22, HNP20]. -urn [XZYC20]. -value [XH22]. -values [CK23]. -variation [HSZ21]. -widely [XK24].

/G/1 [SP24a].

1 [LZ20, MP24, SP24a]. 137 [Suz20]. 145 [Sko20]. 155 [BBM22]. 167 [ZCL22]. 198 [ZLMT24].

2020 [Ano20a, Ano20b, Ano20c, Ano20p, Ano20q, Ano20s, Ano20r, Ano20t, Ano20u, Ano20v, Ano20w, Ano20y]. 2021 [Ano21a, Ano21b, Ano21c, Ano21p, Ano21q, Ano21s, Ano21r, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x]. 2022 [Ano22a, Ano22b, Ano22c, Ano22p, Ano22q, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x]. 2023 [Ano23a, Ano23b, Ano23c, Ano23p, Ano23q, Ano23s, Ano23r, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x]. 2024 [Ano24a, Ano24b, Ano24c, Ano24p, Ano24q, Ano24s, Ano24r, Ano24u, Ano24t, Ano24v, Ano24w, Ano24x]. 2025 [Ano25d, Ano25e, Ano25f]. 2BSDEs [KKH24]. 2SGMM [Bas24].

8/109836 [ZLMT24]. 87 [BRBB23].

A. [Sze22]. absolute [Bar23, Che24c, DSH24, Egh21, JB25, Kap22, Sar21a]. absorbing [HZY21, ZM23]. absorption [IM23, LZ22a]. abstract [LZZ24b]. accelerated [DLZZ22, FHSP21, Sai20]. acceleration [HC22]. acceptable [Egh21, XK24]. acceptance [SMRH24]. Accounting [Gir20]. accuracy [CBK22]. action [LG24]. active [DPSR20]. activity [PSX21]. Adapted [WFJ24]. Adaptive [MS23b, Pen23, ZF24, LZZ24a, May23]. ADCINAR [ZK24]. Addelman [JWW21]. Addition [CML20, GCW24]. additive [Fak20, HW23a, SMRH24]. Adjusted [ZLWZ20, BSB20, DH20, PC24]. adjusting [ZW23]. admits [XZ20]. admixture [CS24]. after [HFC22]. against [BKM20b]. age [AD20, Che24d, FC23]. age-structured [Che24d]. ageing [BK24]. aggregate [GLL23a]. agreement [WDP^{+}20]. AIC [DT21b]. algebras [VM21]. Algorithm [SH22, AH23, RT22, SB21, Wri20]. Algorithms [BPB24, Hah20]. Alhakim [Wei23]. alias [CLLP23]. alleles [Ful24]. allocation [AY20, MDS25, SN20, Wri20]. allowing [ZX25]. Almost [DM21, GM22, LL24d, Haf23, LXZ20, LEJT22, LZ21, PR21]. alpha [Ais23]. alternative [Liu20b, MO22]. alternatives [BKM20b]. always [FV25]. amalgam [YM24]. American [LW24b]. among [JZ23]. amount [GLL23a]. amplitude [Nał21]. Analysis [Mod23, AF22, BAB24, FEVC21, FRR^{+}20, For23, Lee21b, LGHG23, LL24b, LXC20, Lv22, MLZ23, MPD20, Mao22, NAV21, Tak24, WZGY21]. analytic [Bat21]. analytical [Las23]. analyzing [JS20]. and/or [LM21]. Anderson [BN24, Lyu20]. anisotropic [CWW20, CX24, SQT24]. anonymised [Kro22]. ANOVA [ZCP21]. Anticipated [XX20b, KIMG24, WS20, WL21b]. anticipative [Lar22]. any [Arn21, yCLjL23, PZL24]. anytime [TG23]. anytime-valid [TG23]. appear [MU23]. Application [HL21a, SN20, BBL22, BCDR23, BU24, Bro22, DT21a, GM22, Kub20, LPŠ21, LZH23, Nał21, DMPP23, Ost22a, SBK20, SH22]. applications [Ber23a, BRBB14, BRBB23, Bud22b, EH22, FY24, HLL20b, JB20, Kay24, LBB20, LL20a, LL21b, LT22c, MK24, NN21, Pas22, QW21, RT21, VY20, Xu23]. approach [BDK24, BVG20, BT25, CD22, CLZZ21, FV25, GS22a, GI24, HSZ21, HA24, Kaw21, LM24b, MST20, MP24, NS21, WZ24c, YZ23b]. approximate [ZG24]. Approximating [GK20, Lu22]. approximation [Are24, BCM21, BU24, Bat21, BTS23, EN24, Ful24, Hug21, KP22b, KK24, LX24b, Pin20, RS24, RB20, SS23b, Sha25]. approximations [CLP22, Jou22, KP24, Sun20]. April [Ano20a, Ano21a, Ano22a, Ano23a, Ano24a]. arbitrarily [BDK21]. arbitrary [FLW22, GLL23a, Sze20]. Arboreal [Hua24b]. ARCH [LL21a]. area [FK21]. ARFIMA [ZLWZ20]. argmaximum [DR24]. arising [Jas20]. ARMA [ZLWZ20]. armed [AY20]. arrays [DDG22, DS20b, JWW21, PZL24]. artificial [HNS23]. aspects [KSW25]. assignment [LT22b, MS21]. associated [Gho23a, KD24, ZZ24d]. association [ATT21, WK21]. assumption [Sai21]. assumptions [LTZ22, ZLMT23, ZLMT24]. asymmetric [Bez21, CM21b, CH23a, PZL24, WK21]. asymmetrical [WZQ25]. Asymptotic [AF22, Aug20, BNO21, BL20, BS24, CBK22, DNTV22, DFG23a, FLH23, Gir20, HR20, HL21a, HNS23, HLL20b, IM20, IM23, KMS21, MN20, MR23, NS22, NNO20, NW24, Pac20, RB24, SM22, SMS20, TLX22, WWL23, WSM24, XP20, Yua24, ADM24, DLS20, FHSP21, HPH24, HA25, HNP20, Izm20, KC22, LT22a, LT23, LT22c, ML24, RK22, Rz21, Sai20, Sko19, Sko20, Vir21, VIVL23, Zha23]. Asymptotically [LL21a, KS20b]. Asymptotics [BDPR21, HW20, LCF21, BW23, BGV20, FS21b, FS24, Häg20, HY23, IMT24, KM20, KSX24, LCCZ24, Lyu20, NP24, Roz21]. atomic [YY20]. atomless [KS20c]. attempt [PD20]. attraction [GM24, ZM23]. attributes [MD24]. augmented [GYZ22]. August [Ano20b, Ano21b, Ano22b, Ano23b, Ano24b]. autocorrelation [CP22]. autocovariance [VIVL23]. autonomous [TL22, WXY22]. autoregression [CP22, Che24c]. autoregressive [HT23a, LL21a, LZ23a, MY24, PS21, Tra21b, WFGR24, XXZ23, ZLS24]. average [NN21, TVY20, XK24]. averaged [KR24]. averages [BSST20, GM24, Rz21, Sol22]. Averaging [CLR20, CK23, CB20, Rad23, KKC^{+}24, MYX22, PD21, WXY22, WW21b, YWY23]. Azuma [Das24].

Bablok [BH20]. Backward [YZ20, BBM22, DL21, FF25, HL21a, KIMG24, LMN24, Mar20b, SRK22, WS20, WL21b, XX20b]. Bacry [FS20a]. Bahadur [Pyc21]. balanced [GCW24]. Balancing [LEJT22]. ball [NP24]. balls [LLM20]. Banach [Myr21]. band [MS20b]. bandits [AY20]. bands [Sub22, ZYLB22]. bankruptcy [HCW22]. Barak [MT22]. Barrier [KKS24, HCW22, IM23]. Bartlett [BCDR23]. Bartlett-type [BCDR23]. barycenters [GW23]. base [BL20]. based [Ahm20, BS24, Bro22, CHHS25, DLLY25, Din20, FY24, FHM21, GVL20, HLS24, JWW21, JB25, KKS24, KR24, Lan24, LR22, LZZ24a, Li24a, LD21b, LNN22, LWX21, RG20, SJSL21, Sin23, SD22b, VY20, VBS24, WX21, WW21b, WH24, XC22a, ZF24, Zha25, vNBR21]. baseline [YZ23a]. bases [HK24]. basis [LW24a]. batch [CBK22]. Baum [Fen23]. Bayes [JW24, RB24]. Bayesian [Bab20, Bic20, DR21, FRR^{+}20, GH23, HT21, Hug24, Kaw23, KP23, LZZ24a, Li24b, MST20, DML24, PD21, SN24, SL25, ZCP21]. BDS [LBZH20]. Becker [HA24]. behavior [Aug20, Che25, Gra22, HA25, JX22, KSW22, LM24a, ML24, VIVL23, WWL23, WSM24, Zha22c]. behaviour [JSV20, KMS21, MR23, NNO20, SM22]. being [JZ23]. belong [Jur23]. Benford [BP20]. Benjamini [Izm20]. Bergomi [FFGS22]. Berkson [SBS20]. Bernoulli [BSB22, BM20, GR21, RB20, XZ22, Zha20b]. Bernstein [AKW22, WG23, Zha20b]. Bernstein-type [Zha20b]. Berry [CQ25, DT23, Döb24, HLX22, KS22]. Besov [Cha24]. Bessel [MR24]. Best [JQS20, Pin21a]. Beta [Ful24, BJ22, HD23a, JB21, Jon22]. between [BTSL21, BP24, BKB24, BPF20, DM21, DSH24, EMS21, FLW22, GLL23a, GP23, MR22, Vog20]. beyond [TG23]. bi [CD21]. bi-clustering [CD21]. Bias [SPG22, ZK22, AZvdH23, DSH24, RA23]. Bias-correction [ZK22]. biased [BMT23, BPB24, RGS22, YDB23]. BIC [Bar23]. Bickel [AAD25]. bidimensional [FLW22, LCCZ24, Sha25]. bidirectional [Nis24].

Bifractional [Tal20]. bilinear [HA25, WK21]. bimodal [Arn21]. binary [Hwa21, KC22, MD24, Ost22a, RZ22, SPG22, SCGW20]. binomial [Ahl22, BAB24, BRBB14, BRBB23, DZ21, DW20, GTH24, Hug24, Jan21, KK24, Mah23, Pin20, Pin21a, SP24b, XX20a]. birth [Dal24, Vid20, VK24]. birth-and-death [Vid20]. bivariate [Che20a, FS21b, HPN22, KM22, PvdHZ23, Sep20]. Black [RN24]. Bloch [YS24]. Block [CHHS25, CML20, JH22, Lia20, WH24, ZY20]. Blow [LZZ24b]. Blow-up [LZZ24b]. Blowup [LW20]. BMO [KO21a]. Board [Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano25a, Ano25b, Ano25c]. Boltzmann [LMZ21]. bond [JLL20]. Boolean [Bez21, Fak20]. Boosting [Vog20]. bootstrap [BDK24, BCHL24, KKI21, Moj21, SM24, ZCP21]. Borel [EH22]. both [BH20, LLWZ20]. Bound [DB21, Che20b, DNH24, Döb24, FJ24, HLX22, JS20, KS22, LQ22, MKF20, Mui20, Pin21a, VBS24]. boundaries [HMRT21]. boundary [HWW23, ZX21, Zha22c]. Bounded [Zha22b, FJ24, Hah20, HS22, Wri20, ZY20]. boundedness [LZ23b]. Bounds [KZ21, Zaj20, Zei21, Ahl22, CW20, Cha24, CD21, CY25, Das24, DSH24, Fro21, Gho23a, HS22, HNP20, JB25, KP21, Kri22, LSS22, LT22a, MW20, Pek22, Pin22, QX23, YKK20, ZSK20]. brain [FEVC21]. Branching [Kev24, BG23, Che24a, Che24d, CH23a, DFG23a, Gao21, HLX22, Le22, LZLX25, Liu24, MX22, Mit21, RS24, RK22, SP22, Tzi20, WWL23, Ye24, ZH22, ZZ24d]. Brascamp [BU22]. bridge [Lee23]. broad [WDP^{+}20]. broken [Ver22, ZX21]. Brownian [Bai20, Bib20, BV20, CÇ20, Che24a, Daw21, Fal22, GS22a, GWWY22, He24, HD23b, HCW22, HL21b, Hua24a, IM20, IO21, KN22, Kry22, Lee23, LZ22a, Lou23, LZ22b, MT21b, Mar22, MO23b, MMB20, NWY20, NS22, PCPS24, Sha25, Shi20, Tak24, TT23, Tal20, Tou21, Val23, WSM24, XC22b, Xu23, XY23, YY22, YB20a, ZJ21, ZHD24, ZX21, ZL25]. Brox [GPP20, GPP22]. BSDEs [CZ23, HWX22, He24, KO21b, KKO22, Luo20, MT21a, Mar21b, ZJ21, ZDF24]. building [CX23]. Burgers [HW20, ZWW22].

calculation [CLLP23]. call [KKS24]. can [Arn21, BDK21]. canonical [Suz18, Suz20]. Cantelli [EH22, Tem23]. capacity [CM21b, YJT20]. capital [YMWW25]. Caputo [AM21, HT23b, HA25, LHW24, XFJN22]. Carlo [CJ23, ZYLB22]. case [FS21b, Gra21, MG20, TP22]. catastrophes [Zei21]. categories [ATT21]. Cauchy [CJM20, FA24, Oka20, Sol22, XCDS22]. Causal [LSFZ24, Mer23, DK22, Ken21, Pas22, ZLMT23, ZLMT24]. causality [VP20, VM21, WZGY21]. cause [Kel24]. censored [Efr24, FE21, KDN21, LSFZ24, LD21b, RRL20, Sub21, XOC21, Yu20, YZ23b].

Censoring [DMPP23, HC20, JB20, LLM24, MP21]. censorship [Sub22]. centered [BP24, LT23, NN21]. centers [AAD24]. central [Aug20, BM22, Chi22, DS20b, FM20, Gao21, Jan24, LZ21, LZLX25, MS23a, Nas20, TP22, WL21a]. centralized [CC24]. centre [Sin23]. Centroid [HA24]. centrosymmetry [BFK21]. certain [BDPR21, EA20, Gho23b, Has21, Kre22, Sko19, Sko20, WW21a]. chain [AFSW20, CH23b, HC22, Syl22, Tra21a]. Chainability [BLL25]. chains [CM21b, Col23, Cru20, Dal24, HX23, LNYZ22, LM23, LNN22, PSZ24, QK22, Rie22, XFZ^{+}21, XCD20, ZSK20]. chance [NLS24]. Change [YZ22, CE24, DK22, HT23a, Mar21a, RZ22, Zha24]. Change-point [YZ22, DK22, RZ22]. changed [KMS21, Vid23, Wan22c]. changepoint [Hah22]. changes [Bib20, Gzy21a]. Chao [FD21]. chaos [BC24, KP22b]. Chapman [LN20]. characterisation [GLS25]. characteristic [Che24b, Jur23, Nał21, WW21a]. Characterization [Ahm20, Myr21, Fak22, JK22, JP20, LYC20, VBS24]. Characterizations [SBK20, YLP23, KM22]. Chatterjee [WFJ24, Zha23]. Chebyshev [Jak21]. Chervonenkis [CD21]. chi [KSW25]. chi-squared [KSW25]. Choice [SZ24, WG23, MD24]. Cholesky [BVG20, LZZ24a]. Choosing [Lee20]. Chvátal [BPR23]. Chvátal’s [GTH24, LXH23]. Ciesielski [KN22]. CIR [GWWY22]. circle [HL21b, JGA24, Łuc21]. CKLS [MRD22]. claim [FLW22, GLL23a]. claim-size [FLW22]. claims [GLL23a, LBZ20, LCF21, YL22]. Clarke [AZ23]. Class [Bas24, CH23b, CN24, Das24, Häg20, HWX22, Jur23, KSMF20, LMSY22, LG24, MLZ23, MGMSP24, NNO20, SQT24, XNZL24, Yu22, ZAP24, ZSK20, ZYL23]. classes [Haf23, KP21]. classical [HSZ21, HY23, KM20, SY24, Vid21a]. classification [FEVC21, MK24]. clique [Zha20a]. closed [LMN24]. closed-form [LMN24]. Closure [BJ22]. CLT [Döb24, Kub20, LBL^{+}22, MP20a]. cluster [AAD24, BFK21, GS20]. cluster-weighted [GS20]. clustered [HPH24]. clustering [Bro22, CPD20, CD21, DM24, LZ25, ZS24]. clusters [BL20, JLL20]. Coalescence [LRW24]. coarse [Yua24]. coefficient [Bud22a, HH20, HT23a, Hwa21, LWZ25, LYC20, TLW22, Tra21b, WFJ24, XX20b, Zha23]. coefficients [AGP20, CB20, CZ23, DMS20a, EMS21, FF25, HMRT21, He24, JLZ24, LZMZ20, PS21, WXY22, Yan25a, YZ20, ZJ21, Zha21a, ZH24]. cold [LL23a]. collecting [Dou22]. collective [KRN23]. COM [BRBB23, BRBB14]. COM-Poisson [BRBB23]. combination [KSW25]. combinations [Fro21]. Combinatorial [Kop24]. Comments [PS21, Wei23]. commodity [KSAA24]. common [CCY20, Lev21, WjWbK22, ZX25]. comonotonic [SMRH24]. compact [MP21]. compactness [Kar24]. Comparing [Wal23, BSB22, LYY21]. Comparison [ALO22, YY22, ASKI24, BH20, BKM20a, GN23, KO21b, LDM22, Sto20, WL21b]. comparisons [BKZ21]. compatible [Gho23b]. compensator [SS24]. competing [CZLC20, LPY24, TLW22]. competition [Le22, RK23]. complement [Jak21]. Complete [Vid23, XK24, Fen23, WWZ20, WYWS21]. complex [MLC25, TP22]. complexity [BGV20, CD22]. Component [GCW24]. components [ASKI24, Bur21, GVL20, JA21, LL23a, WC21]. composite [LH25]. composition [FC23]. compositional [LW24a, LWM22]. Compound [KdS23, ASS24, BM22, GLL23a, LCCZ24, SY24]. compression [JQWD23]. computation [Nał21, Per24, Pri21]. computational [KSW25]. computer [LLWZ20]. Computing [BMP20]. concave [Ari23, Arn21, FY24, MP25, SMS20]. concavity [BdCP23, LLM24]. Concentration [Tan24, BK22, CD21, KXBS21, LS22, LSS22, LT23, Pin22, ST21]. concept [Jas20]. concepts [HSWZ21]. condition [Ahm22, Cha21, CT21, GM21, HWW23, HT21, HJKL21, RT21]. Conditional [CÇ20, GN23, BAB24, BHS24, yCLjL23, Cin22, DS25, DS24, FS20a, Gir20, JP20, Lau24, MC21, Nie22, Nog21, PP23, Xie23, XOC21, XJG21, ZK24, ZYY22, ZYLB22]. conditionally [BDPR21]. conditioned [Sun25]. Conditions [BN24, CD20b, HIS24, KKC^{+}24, Lee21b, NV24, ZDF24]. conductivity [MRZ20]. conference [HLS24, WLL22]. Confidence [WML22, BSB22, BLK20, Bur21, Kab24, Sub22, TG23, Wri25, ZYLB22]. confirmatory [Lee21b]. conformable [Ahm22, NNO20]. conformal [SD22a]. confounder [Mao22, SPG22]. confounding [LKA20, ZCL22]. conjecture [BPR23, LXH23]. Conjugate [RA23]. Connection [CX23, MGM21]. connections [AIJD21, KB21]. connectivity [HNS23]. consecutive [DS21]. conservative [FD21]. Consistency [RZ22, XX20a, YLW22, CD22, CHW20, DP23, FMR22, GS20, Izm20, Kaw23, Kro22, Lar22, WYWS21]. Consistent [DS25, ZLS24, GLLOS24, Ken21, MN20, NOH21, PSP^{+}25]. constant [CH23b, Hah22, KM22, LHW24, MZ20, MRZ20, Naa21, Tak24, XFJN22]. constant-differentials [MZ20]. constrained [BDK24, JQS20, KP23]. constraints [LLM24, NLS24, SLA21, TVY20]. Construction [Ahm22, CML20, LKA20, LYL22, LLL^{+}25, MPD20, ZCL22, ZYLB22, KN22]. contact [Sto20]. Contaminated [DM24, FFS24]. Contamination [HIS24]. contextual [AY20]. contingency [ATT21, TG23, WK21]. Continuity [Yan25b, Döb22, DM20, HW24, HC23, NV24]. continuous [Ahm20, AGP20, BDK21, Col23, Le22, MC21, MST20, MRD22, XFZ^{+}21, Yan25a, Yu22, ZSK20, Zha21a]. continuous-time [BDK21, Col23, MST20, XFZ^{+}21, ZSK20]. continuum [LX24b]. contoured [JW24, LDM22]. contraction [Tra21a]. control [yCLjL23, Hah22, RK22, Wan22b]. controllability [Ahm22, AZ23]. controlled [Liu24, RK22]. Controlling [DZZ21, RK23]. conventional [BSB22]. Convergence [AVA23, FFS24, KS20a, LMN24, MGD20, PP23, QHK25, Rie22, Sar20, CP22, CM21a, DS20b, Fen23, GM22, Gao21, Gho23b, GL21a, HW23a, ITM24, JL25, JH22, KP21, KKG20, KZ21, Lau24, LMSY22, LZLX25, MS23a, MS20b, PR21, RN24, Roz21, Sar21b, Wal23, Wan22c, Xie23, XK24, ZSK20, Zei21, ZH22]. convergent [IK21]. Converse [WL21b, Vid21a]. Convex [WWL20, CLZZ21, FK21, HD23b, JK22, Ses24, SJK22]. Convexity [JSWZ22]. Convolution [Arn21, Fak20, Tor24]. convolutions [Liu20a]. Conway [BK23, GX22]. coordinate [DL24, Fri22]. coordinates [JZ23]. Copas [AZvdH23]. copula [CY25, DS24, Hug21, LD21a, LNN22, Sep20, SS25, WK21]. copula-based [LNN22]. copulas [CDP24, Nas20, Tre20, ZAP24]. corrected [DT21b]. Correcting [DT21b]. Correction [Suz20, ZCL22, BCDR23, FD21, ZK22]. corrections [HLL20a]. correlated [CCY20, DB21, DB23, FS20a, GLS25, Lyu20, MLC25, Sel24, VBS24]. Correlation [TVY20, DB22, EMS21, Gan21b, Hua23, Hua24b, MGM21, NP22, PS21, WFJ24, Zha23, ZJD20]. correlations [ASV24, XCDS22]. Corrigendum [ZLMT24]. cosine [Ma21]. Coskewness [BCRV23]. cost [BBM22, BS22b, DL21, LL20b, Wri20]. costs [Ren24]. count [KD24, DMPP23, PvdHZ23, VY20, Wei21]. countable [XZ20]. counter [CCY20]. counts [CLP23, DT21a, GT20]. coupled [KIMG24, SRK22]. Couplings [Kop24]. Covariance [FHM21, BVG20, CHHS25, Din20, GZH20, HS22, KZB23, KP23, LW24a, LM24b, LBL^{+}22, LXC20, OK21, QTLT20, WX21, WZ24c, ZPC21, ZG24]. covariate [DH20, LSFZ24, OY23, ZW23]. covariate-adjusted [DH20]. covariates [Che21, LP20]. cover [DIY23, Syl22]. covering [nHL21]. Cox [JX22, YZ23b]. CP [Zha20a]. Cramér [CQ25, CL23, DFG23b, Sun25]. Cramér-type [CQ25, CL23]. Cramér’s [MY24]. credible [SL25]. criteria [Buo22]. criterion [Bar23, DMN24, Kha22, NW24, Oka20, WZ24a, Zha24]. critical [CS20, Che24b, HSZ21, LRW24, Mit21, Mui20, RS24, WFGR24]. Cross [WW21b]. Cross-validation-based [WW21b]. crossing [IM20]. crossings [AVA23]. Cumulants [MG20, Pri21]. cumulative [AIJD21, KP22a]. Curie [Lu22]. curing [MDS25]. current [AD20, HC20]. curvature [AGP20, Yua24]. curves [LYY21]. Cusum [Bib20]. Cutoff [Bar21]. cycle [GI24]. cycles [GI24, Yua21]. cyclic [MKF20].

D [FEVC21, YS24, ZWW22]. Data [BSB20, AA21, BH20, BGV20, BAB24, CZLC20, CA20, DDR20, Din20, DNH24, DM24, Efr24, FE21, FEVC21, FHM21, GZ24, HC20, JGA24, KD24, KC22, KS20b, LGHG23, LW24a, LD21b, LWM22, LZ23a, LXC20, LZ25, LP20, LWX21, MEA21, MHHJC22, MGP20, MLC25, NOH21, NAV21, PF23, PvdHZ23, RRL20, RT22, RZ22, SN24, Sub21, SCGW20, TSG20, TLX22, VY20, WX21, Wan22b, WLZ24, Yu20, YZ23b, ZZF22, ZJD20]. data-based [FHM21]. data-row [ZJD20]. death [Sto20, Vid20, VK24, ZZ24c]. decay [MGM21]. December [Ano20c, Ano21c, Ano22c, Ano23c, Ano24c]. decision [BS22b]. decomposable [LBB20]. Decomposing [CCY20]. decomposition [BDBB20, CHW20, CHHS25, LZZ24a, MRG22, Val23, YY20, Zha20a]. decompositions [Ber23a]. Decorrelated [GZ24]. decoupling [QK22]. decreasing [Gra21]. deep [DNH24, SSL22]. Defaultable [LW24b]. definite [AGP20, FPBK20, MO23b]. definitive [WLL22]. deflation [DZZ21]. Degasperis [CZG24]. degenerate [CLP22, HLL20b, LG24, QW21]. Degree [DY24, FLH23, Shp24, WG23, ZY20]. degrees [Oga23]. delay [BTSL21, HT23b, IKCSKH25, KKS24, KIMG24, Liu20a, LZ23b, MLZ23]. delayed [AY20, BB23, LCF21, LZ23b, Luo20]. delays [GS21, HFC22]. delta [Pas22]. demeaned [NP24]. demimartingales [FZ23, FY24]. denseness [Kut21]. densities [CW20, Cha24]. density [AM19, CLR20, DS24, DW20, Efr24, GS22a, GPP20, HX23, JSV20, KP21, Lar22, Lev21, Lou23, MC21, Mar22, MG20, Moj21, MP20b, MP20c, Wal23, Wei23, WYWS21, XOC21]. density-dependent [HX23]. denumerable [Cru20]. Departing [Bic20]. departures [Lan23]. Dependence [LNN22, OY23, BCRV23, DMS20b, FLW22, GLL23a, HPN22, HPH24, JL25, KKI21, LL21b, MM22, MSM20, PSP^{+}25, Sep20, XH22, ZAP24, Zha22b, Zou23]. dependent [ART20, ASKI24, BB22, Bou20, CY21, Che25, DNH24, ERR23, GK20, GR21, HX23, HC20, HP25, KP24, LBZ20, LCF21, MYX22, Mau24, NOH21, RS24, RT22, RB24, Sze20, Tan24, YJWY20, YL22, Zaj20, ZZ24b]. depth [Zuo20]. Derivative [SXX20, CX24]. Derivatives [HX21a, MP20b, MP20c]. derived [HSWZ21]. Deriving [Chi22]. descending [FP22]. design [HZZZ21, HT21, OY23, SB21, Sai25, XLNQ20]. designs [BFK21, CML20, CLLP23, CA20, DMS20a, DMS21, GYZ22, GCW24, HH20, HLS24, Ida20, LKA20, LYL22, LLL^{+}25, Lia20, LYC20, LYY21, MPD20, MD24, Shp24, SDH20, SWZ20, Wai24, WLL22, WZ24a, WZQ25, ZCL22, ZGQC20]. detection [CS24, DK22, Hah22, HT23a, Mar21a, OGR24, YZ22, Zha24]. Determinacy [WM21]. determinantal [Hir21]. determinants [DT23]. Determination [OH21]. determined [Aug20, DAA21]. Deterministic [RTT21, HL21a, RS24]. Deviation [Gan21a, AAD25, Che24a, CQ25, DSH24, EH22, FWY20, HD23b, Kap22, LMS22, MT21c, Sun25, WZ24b, Xu23, YS24, YL22, ZL20]. deviations [BM22, Cha21, CC24, CL23, DFG23b, Fan24, FLW22, GLL23a, Gao24, HX23, LBZ20, LM23, MY24, Nak22, Sel24, SYM25, WFGR24, XCD20, ZZ24d]. diagonal [ATT21, DD23, SD20, San24]. diagonalizable [PG20]. Dickman [Wei23]. diffeomorphic [CS20]. difference [CDP24, Che25, DS22a, DLLY25, For24, VBS24, Zha22b]. differences [Vog20]. different [AZvdH23, Lee20, Oga23, Vir21]. differentiability [Suz18, Suz20]. Differential [HD23b, AM21, AZ23, BBM22, BTSL21, BT25, CB20, DL21, FLLW24, FF25, HSS22, HJO20, HT23b, HA25, IKCSKH25, JL25, KKS24, KIMG24, LXZ20, LHW24, Liu20a, LZ23b, LL24d, LG24, Lv22, Mar22, Mer23, NS22, QW21, RTT21, SRK22, SQT24, TL22, WXY22, WW23, WL21b, XX20b, XFJN22, YY22, YZ20, YWY23, Zha21a, ZHD24, ZZ24b]. differentials [MZ20]. diffuse [BL20]. Diffusion [SS23b, TT23, BDY22, Buo22, EN24, GPP20, GPP22, HSS22, JX22, JLZ24, LW24c, MR24, QX23, Sai20, SXY22, SW20, SCMZ24, XFJN22]. diffusions [CN24, FHSP21, Sar20, Tra21a]. diffusive [KSX24]. dimension [Che21, CQ25, Daw21, DB22, DL24, Fal22, Hua24a, San24, Wan22c].

Dimensional [WDG23, AA21, BP24, Bez21, CZLC20, Che21, CLP22, Din20, DLZZ22, GZ24, GT20, HY23, HSS22, He24, Hua24a, HNP20, JLZ24, KK25, KP24, Kry22, LT23, LW24a, LWM22, LM24b, LZMZ20, LBL^{+}22, LZ25, MGM21, QTLT20, TLX22, Tem23, Tre20, WX21, ZY20, ZPC21, Zha21a, ZF24, ZG24, ZGZF21]. dimensions [Vog21]. direct [Hug21]. directional [DM24]. directions [Gat24]. Dirichlet [BJ22, BDY22, HW23b, Itk24, ZH24]. disconnected [EA20, TT23]. disconnection [CX23]. discontinuous [PSX21]. discounted [BS22b, IK21]. discounting [Gap20]. discovery [DZZ21, Izm20, RK23]. discrepancies [JQWD23]. discrepancy [BNO21, MN20, RG20]. discrete [ADM24, Ari23, BdCP23, BJ22, Gho23b, Hil20, HX21b, Hug21, IW20, Jas20, Kha22, KD24, KM22, KMP25, KKT24, LPY24, MP25, NS24, NMR25, SS24, SN20, SL25, WML22, Wan22a, Wan22b, YJWY20, ZS24]. discrete-time [SS24, Wan22a, YJWY20]. disintegrations [BHS24]. disorder [Buo22]. Dispersion [BW23, BK23, GSL20]. dispersive [AEAC21]. displacement [BH23]. distance [BP24, DM21, EMS21, FM20, GYZ22, Mod23, MGD20, PG21, Sar20, SWZ20, TLX22, Ye24, Yu22, vNBR21]. distance-based [vNBR21]. distances [MP25, Rie22]. distorted [Shi20]. distortion [Sep20]. distortions [CDP24]. distributed [BDPR21, Fri22, Jas20, NLS24]. distribution [AA21, Arn21, BB23, BU22, BK23, Bic20, BRBB14, BRBB23, Bro22, CE24, Che24c, Che24b, DR24, FLH23, FRR^{+}20, FK21, For24, FS21b, Gau23, GLS25, GX22, Gho23b, GM24, GTH24, HNP20, JB20, JLL20, JW24, Kap22, KW21, KK22b, KKG20, KKT24, KK24, KSW25, LPY24, LMN24, MYX22, Mao22, MG20, MR23, Mau24, MP24, MGD20, NW24, Oga23, PZ20, Pin20, RRL20, Rz21, SBS22, Sol22, TLX22, TP22, WL21a, WFZ25, WM21, Wei23, WDG23, XFZ^{+}21, Xie23, XZYC20, Yu20, Yu22, Yua21, Zha23, ZM23]. distributional [Hug21]. Distributions [MZ20, Ahl22, Ahm20, Aly20, Ari23, BJ22, BSB22, BDK24, Ber23b, Cha21, Che20a, CJM20, CFH23, Cru20, DM21, DM20, DW20, DH22, ERR23, EA21, EK20, GW23, GP23, Gra21, GZH20, GPP22, Häg20, HY23, HMRT21, HR22, HD23a, Itk24, JP20, JB21, KK20, KD24, KKG20, KMP25, KP23, Kut21, Lan24, LRS20, LXH23, LNYZ22, MBR22, MA22, NS23b, Nał21, DMPP23, Ost22b, PG21, Pin21c, Pyc21, SN20, Tem23, Tor24, WW23, Yan25b, Yan25c, ZS24]. divergence [Gho23b, LV23]. dividends [Ren24, YMWW25]. divisibility [Kha22]. divisible [BLL25, CCH20, GX22, Kaw21, KKT24, Kut21, MR23, NS23b]. DMTTF [BKM20b]. do [Jur23]. does [WML22]. Domain [Ber23b, ZM23, CZ24, GM24, LZ24a, WZGY21, ZL25]. domains [HD23b, LZ22b]. dominance [LBB20, MSM20]. domination [HR20, RT21]. Doob [Kaw23]. Double [HCW22, NP20, ZGZF21]. Double-barrier [HCW22]. doubly [MYX22]. down [ASV24]. down-correlations [ASV24]. drift [FP22, FHSP21, IM20, LZ24b, Lou23, PSX21, Sai20, Tak24, YSZ20, ZX21]. drifts [LW24d]. driven [GWWY22, Gok25, He24, HJO20, HD23b, HW23a, LL20b, LZ24b, Lin24, MLZ23, Mar22, NS22, NMR25, Rad23, SXY22, SXX20, WW23, WZZ23, XY23, YY22, YZ20, YWY23, ZYS20, ZJ21, Zha21a, ZYL23, ZHD24]. driving [Mer23]. dry [GS22a]. dual [Rie22, SBK20, Zha24]. dual-criterion [Zha24]. duality [Ses24]. Duals [Jon22]. duration [Dem21]. Dvoretzky [Naa21, XZ23]. Dynamic [FM21, DIY23, FC23, LH22, ZLZK20]. dynamical [GGX21, Kre22]. dynamics [TWW21]. Dynkin [Mar23].

edge [GT20]. Edgeworth [HLL20a, Wei21]. edit [Gan21a]. Editorial [Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano25a, Ano25b, Ano25c]. effect [BFK21, yCLjL23, LSFZ24, ZGZF21]. Effective [Egi22, FV25]. effects [CX23, IMT24, LZ23a, LL24d, MD24, MHHJC22, Sep20, YDB23]. Efficiency [KR24, DNTV22, Gir20, LLL^{+}25, LQZ24]. efficient [Pen23, WLZ24, ZZZD21]. Efron [Oud21]. Eggenberger [FM21]. Ehrenfest [XZYC20]. eigenvalue [BDY22, KK22a, MS20b, SM22]. eigenvalues [HW23b, LBL^{+}22]. eigenvector [KK22a, LZH23]. elementary [Kre22]. elephant [GM22, GS21, GS22b]. ellipsoid [Val23]. elliptical [Bro22, GZH20, LRS20]. elliptically [JW24, LDM22]. embedding [NP20]. Empirical [BC24, LD21b, SD20, Are24, BSST20, DLLY25, DM21, GZ24, Gir20, LT23, Li24b, MPK22, Rie22, ZLWZ20]. empirical-variance [Gir20]. END [CC24]. Endogenous [yCLjL23, ZGZF21]. ends [XZ22]. energy [Gir20, Pir24]. Ensemble [DS22b]. entrance [FLZ20]. entries [DD23, Mau24]. entropic [MT21a]. Entropy [ITM24, LT23, ZS24]. enumerator [WZQ25]. envelopes [Mar20b]. environment [Gao21, HLX22, Kev24, LZLX25, SS23b, SP22, Ye24, ZH22, ZZ24d]. environmental [TWW21]. environments [Mit21, TT23]. Epidemic [DK22, LX24a, MDS25, WFZ25]. epoch [FP22]. equality [BCDR23, BN24, HNP20, MEA21]. equation [CZG24, Gok25, HW24, HSS22, Hil20, HW20, IKCSKH25, KKS24, Kum22, LZ21, Li23, LZZ24b, Mar22, NNO20, Rad23, TWW21, YSZ20, YS24]. equations [AM21, BBM22, BTSL21, BT25, Bra23, CFH23, CB20, DL21, DP22, FLLW24, FF25, HJO20, HCW22, HT23b, HA25, JL25, JLZ24, KKC^{+}24, KIMG24, LXZ20, LHW24, Liu20a, LZ23b, LG24, LW20, Lv22, MO23a, Mer23, MRZ20, NS22, OGR24, QW21, RTT21, SRK22, Sol22, SQT24, TL22, WXY22, WW23, WZZ23, WS20, WL21b, XX20b, XFJN22, YY22, Yan25a, YZ20, YWY23, Zha21a, ZYL23, ZHD24, ZWW22, HD23b]. equi [FS21b]. equi-skew [FS21b]. equicontinuity [HSWZ21]. equicorrelational [MGM21]. equilibrium [CDLZ25, Sar20, Sar21b]. equiprobable [Riv21]. Equivalence [BTSL21, HSWZ21, BDK24, Oka20, Ost22a, Ost22b, Tor24]. equivalent [SLA21]. Erdos [BV23, FWY20, Gao24, MT22, XZ23]. Ergodic [Wan22c, Gir24, KS20a, Tem23]. ergodic-martingale [KS20a]. Ergodicity [SQT24, CZ24, GL21a, LNYZ22, Łuc21]. Errata [BBM22]. Erratum [BRBB23, Sko20]. error [BKK20, Che21, Che24c, FE21, Hah20, Hah22, HNP20, LT22a, MW20, Sai25, SBS20]. errors [ART20, BH20, Din21, Gir20, LL21a, LM21, Mar21a]. errors-in-variables [Mar21a]. escape [Tzi20]. Esseen [CQ25, Döb24, HLX22, KS22]. Essén [DT23]. Estimate [GL21a, LLMB23, Osȩ24, SBS20]. estimates [EH22, Gan21a, GGX21, KO21a, PK24]. Estimating [AEMZ24, BGV20, CBK22, CX23, MHHJC22, SY24, Bra23, DMS20a, DMS21, Gir20, KD24, OGR24]. Estimation [HX21b, WX21, ZP20, AY20, BVG20, BPB24, BHS21, Bra23, CL20, DLS20, DS24, DH22, Efr24, FE21, FHM21, GLLOS24, GZH20, HC20, HPH24, Hwa21, KDN21, KP22a, KK20, KY22, Kub20, KP23, LWZ25, LPY24, Lee20, Lee21a, LD21a, LTZ22, LZZ24a, LSFZ24, LW24a, LWM22, LM24b, LH25, LQ22, LZ23a, LWX21, MC21, Mar20a, MR23, May23, MST20, MO22, MRD22, Moj21, MP20b, MLC25, NMR25, DMPP23, Pen23, QP23, RGS22, RZ22, Sai25, SXY22, SJZ23, SCMZ24, Tre20, WLS20, WZ24c, XXZ23, Xie23, XOC21, YB20b, ZK24, ZYS20, ZPC21, ZHD24, ZG24, ZLS24, ZGZF21]. estimations [LP20]. estimator [AAE22, BNO21, CBK22, Che24c, DMN24, FD21, GS20, HLL20a, JSV20, JW24, KKI21, KjL24, KR24, Lar22, Las23, MN20, MT21c, MY24, MK24, MP21, NOH21, PvdHZ23, RL20, SCGW20, SJK22, WFGR24, WYWS21, Yu20, ZL20, Zou23]. estimators [Bas24, BCHL24, BFMS22, BS22c, BFK21, CBK22, CLR20, JQS20, Mar22, MP20c, NS22, PF23, Pin21b, SJSL21, VY20, ZK22]. Euler [Che24b, GWWY22, HW23a, LZ24b]. Eulerian [Gan21a]. events [Fro21, LSS22]. Every [XZ20]. evolution [Ahm22, DY24]. evolving [FLH23]. Exact [BAB24, DLS20, DDR20, Gao21, HC23, Lee23, Lef24, LZLX25, PC24, TG23, Yan24, Gra21, LXZ20, MEA21, MA22, MMB20, SL25, Wri20]. exactness [Hug21]. examples [Hua23]. exceedance [Bat21]. exceedances [LT22c]. exceeding [Pin21a]. exchangeability [CK23]. exchangeable [CLP22, DDG22, RB24]. excitation [BHS21]. exciting [SPT23]. exist [WML22]. Existence [KIMG24, PSX21, SRK22, Ber23b, KKO22, NNO20, XNZL24]. exit [HW23b, LWP21, LZ22b, ZL25]. expansion [CQ25, SY24, YB20a]. expansions [DFG23a, XP20]. expectancy [CPL20]. expectation [GL21b, GLL23b, Jan21, Liu20b, Nog21, Pin21a, Xu23, Zha22a]. expectations [Fen23, JSWZ22, PP23, SWZ21, XK24, Zha21b]. Expected [Dem21, Buo22, Che24b, Hah20, MU23, Vid20, YKK20]. expectile [LLMB23]. experiment [SB21]. experiments [GCW24, LLWZ20, LLL^{+}25, Lia20, XLNQ20, HA24]. expert [KVW24]. Explicit [ZZ24c, NP20]. Exploration [AZ23]. exponent [Bib20].

Exponential [BV23, CZ24, HR22, HZY21, MGM21, Sol22, Tra21a, ASKI24, FLLW24, GL21a, Hwa20, IK21, LXZ20, Ma23, NP20, Pin22, Sar21b, ST21, TL22, WC21, Yu20, YL22]. exponentiality [BKM20b]. exponentially [nHL21]. Extended [ATT21, KK22a, HCW22, SPT23, WLS20]. Extending [Lee22]. Extension [KK25, BW23, Osȩ21, Rib20, WK21]. extent [Bic20]. extinction [Le22]. extrema [Rat21]. Extremal [Gan20]. extreme [ADM24, LT22c, Yan25b, Yan25c, YZ23b]. Extremes [Bai20, CD20b, GM24, XP20, Yan24]. extropy [BBL22, KP22a, Kay24, NS23a, RGS22].

factor [BVG20, LM21, Lee21b, LGHG23, MD24, ZG24]. factorial [Lia20]. factorials [YZ23a]. Factorization [Bab20]. factors [BU24, DZZ21, LM21, LLWZ20]. failure [DLZZ22, HFC22]. falling [Aug20]. false [DZZ21, Izm20, RK23]. families [FA24, LBB20, PZL24]. family [AIJD21, Bur21, HX23, Lan21, Pin21c, Pyc21, Vil24]. fast [BDK21, LMSY22, RL20]. fat [ZZ25]. feasible [Egh21]. feature [CZLC20, DLZZ22, TLW22]. February [Ano20p, Ano21p, Ano22p, Ano23p, Ano24p, Ano25d]. Feller [CN24, FLZ20, Zha21a]. few [FS20b]. Feynman [SXX20]. FGM [ZAP24]. FGM-type [ZAP24]. field [FFGS22, FF25, Gao24, HWX22, IW20, KSX24, Lyu20, MFF22, MS21, SRK22]. fields [CWW20, CX24, CS20, Che24b, FWY20, FS20a, HC23, HX21a, Lee21a, LT22c, LLM20, LDM22, Lu24, Ma21, Ma23, Mui20, NV24, Tem23]. filling [LYL22, WZ24a]. filtration [SWZ21]. financial [YJWY20]. Finding [SN20, Gho23b]. Finite [Hir21, BKZ21, BKB24, Col23, FD21, Gra21, LPY24, LCCZ24, Pas22, PSX21, Tem23]. finite-dimensional [Tem23]. finite-time [LCCZ24]. finitely [WM21]. First [JLL20, Che24c, CFH23, EH22, FP22, KR24, Lee23, Lef24, LZ22b, Rat21, Vid23, ZX21, ZL25]. first-order [Che24c]. first-passage [Lef24]. Firth [BSB20]. Fisher [WL21a]. fit [AAD25, BKK20, EA21, FLW23, SJSL21, Sin23, WH24]. Fitting [HJKL21, Gir20]. fixed [GYZ22, LZ23a, SS25]. flatness [AM19, Wei23]. flattening [FFGS22]. flow [CY21]. Fluctuation [LL24b, NP22]. Fluctuations [DD23, DS22a, GT20, MS20b]. foldover [ZGQC20]. Föllmer [EN24]. follow [JB21]. Forced [Gzy21a]. form [KK25, LMN24]. forming [Ver22]. forms [MS20a, Zaj20, ZZ25]. formula [Are24, CM21b, EK22, IKCSKH25, KKS24, LHW24, RN24, SXX20, XFJN22, Yak21]. formulae [Liu20b, SBS22]. Formulas [LV23, MRG22]. formulations [LM21]. forward [HL21a, KIMG24, SRK22]. forward/backward [HL21a]. Fourier [MST20]. fractal [HSZ21, TT23]. Fractional [MR24, AM21, Ahm22, AZ23, BM22, Bib20, CML20, Daw21, Fal22, FS20a, GWWY22, HW23a, HT23b, HA25, KKC^{+}24, Kre22, Kum22, KIMG24, LHW24, LZ22b, Mah23, Mar20a, Mar22, NS22, NNO20, PCPS24, SRK22, SP24b, Tak24, WSM24, XFJN22, XY23, YB20a, YZ20, ZL25]. fractions [YZ23a]. frailty [AIJD21]. frame [Hir21]. frames [Hir21]. framework [Ais23, BT25, Gho23a, Zha24]. Frank [SS25]. Fréchet [MGD20]. Fredholm [CFH23]. free [CCH20, CZLC20, HWW23, LBL^{+}22, MS23a]. freedom [Oga23]. frequencies [EH22, LLM24, LSS22, WjWbK22]. Frequency [Lar22, Tak24, WX21, WZGY21, Zha25]. friction [GS22a]. Frisch [Bas24, Din21]. frog [BDK21]. Fubini [ZX20]. full [JL25, KjL24, YZ23b]. fully [Hil20, Klo21, KIMG24]. function [BMT23, BS22c, BSB20, CL20, CLP23, DZ21, Fak20, FWY20, FE21, HC20, Hwa20, KZB23, KBM21, KB21, Łuc21, MGMSP24, Nał21, Pek22, Pir24, RGS22, RK22, XOC21, YZ22, Zha20b, Zou23]. Functional [IK21, MP20a, AAE22, BGV20, CP22, CB20, DMN24, FEVC21, KMS21, KS20b, LLMB23, LG24, MC21, MGP20, NAV21, RZ22, WDP^{+}20, XC22a, YY22, ZP20]. functionals [BV20, HK21, HT23a, NWY20, XY23]. functions [AIJD21, AEMZ24, BPF20, CR25, EK22, FPBK20, FY24, Gir20, HSZ21, HPN22, JB21, Jou22, Jur23, KXBS21, KP21, LLM24, Lou23, Nie22, SPT23, Sol22, Sub22, Suz18, Suz20, Wal23]. Further [Kay24]. FWER [DB21, DB23, Wan22b].

G [SP24a]. Galton [DFG23b, FJ24, KW21, LRW24, LL23b, TL23]. game [ZZ24b]. games [Mar23, NLS24, SLA21]. Gamma [BB23, CPL20, CS24, ERR23, For24, Gau23, Jon22, KKG20, MR23, NW24, Pin21c, FS21b, KSW20]. gamma-Gompertz [CPL20]. gap [GL21a, LMZ21]. GARCH [SW20]. Gas [Hua24b]. Gauss [PG21]. gaussian [Zaj20, ASS24, BCM21, BCDR23, BP24, CWW20, CX24, CY25, CS20, CLP22, Che24b, ERR23, FS20a, GT20, Gzy21b, HW24, HC23, HX21a, Hua23, Hug21, IMT24, KZB23, KK25, KP24, KKG20, Kub20, Kum22, LV23, LT22a, LT22c, Lyu20, MGM21, MS21, Mui20, NN21, NP24, NP22, Pak20, Pen23, RS22, Sel24, SV20, Sun20, TZT21, XNZL24, Yan24, Yan25b, YB20b, ZYL23]. Gaussianity [LT24]. GE [Gir20, Gir20]. GEE [LXC20]. General [FF25, DY24, BS22b, Bra23, CZ24, DK22, DS22b, JZ23, Kaw21, Ken21, KO21b, KKO22, Kum22, LKA20, LBZ20, LMS22, Liu20b, ML24, SPT23, SWZ21, Wri20, YB20b, ZCL22, ZDF24]. Generalization [Tem23, BRBB14, BRBB23, DNH24, LZ20]. Generalizations [BK22, Oud21]. Generalized [CDP24, CM23, CH23b, MT21a, SP24b, WZQ25, ZGQC20, AF22, BNO21, BU22, BW23, FS21b, GZ24, GSL20, IO21, Ida20, Jur23, KKG20, LD20, OGR24, Ost22a, PG21, Pyc21, SBK20, TWW21, WFZ25, ZS24, ZZ24a, BP20, Itk24]. generated [FK21]. generating [BS22c, Hil20, KB21]. generation [Kaw21]. generators [KKO22, Luo20, ZDF24]. Geo [SP24a]. Geometric [LYC20, BV20, GT20, HCW22, LXH23, SD22b, WDG23]. GFF [HWW23]. Gibbs [Gir20, JH22]. Gibbs-energy [Gir20]. Gilbert [Gok25]. Gini [CDP24, KDN21, VBS24]. given [MU23]. Glivenko [Tem23]. Global [ZWW22, AAD25, CZG24, NNO20, XC22a, YLW22]. GMC [FS20a, FFGS22]. goal [HP25]. goal-reaching [HP25]. goes [DB23]. Gompertz [CPL20]. goodness [AAD25, EA21, SJSL21, Sin23, WH24]. goodness-of-fit [AAD25, EA21, SJSL21, WH24]. Goovaerts [LPŠ21, XJG21]. grade [KP22b]. gradient [OGR24, RT22]. gradually [GS22b]. graph [BV23, DIY23, Gao24, MT22, Ste23, WWZ20, WDG23, Yua24, Zha22b]. graph-dependence [Zha22b]. graphical [BCDR23, LZZ24a, MGM21, ZLZK20, ZZZD21]. graphs [Gan20, Gan21a, GT20, KSH24, LD20, SX24, XZ22]. greater [KP22b]. Green [KdS23]. Greenwood [AF22]. Grey [Kev24]. Grincevicius [Kev24]. group [CHHS25, SMS20, XZ20]. grouped [BH20]. groups [EA20, MEA21]. growing [DY24]. growth [Cha21, HWX22, KKO22, SX24, ZDF24]. guessing [Gho23a]. guesswork [WZ24b]. Gumbel [GM24].

Haar [AJO20]. Hadamard [LHW24]. Haezendonck [LPŠ21, XJG21]. half [Bez21, Sar21b]. half-line [Bez21, Sar21b]. Hamburger [ITM24]. Hammersley [LN20]. Hankel [Mau24, SM22]. hard [LW24a]. Hardy [YY20, YM24]. Harmonic [Gzy21b, CY21, CR25, Gzy21a]. Harnack [HLL20b, YSZ20]. having [ZAP24]. Hawkes [BHS21, Pri21, SS24, SPT23, Wan22a, ZX25]. hazard [AIJD21, CR25, Fri22, HFC22, Lan23, MP21]. heads [FFS24]. heat [CY21, HW24, Hil20, Kum22, LZ21, Li23, LZZ24b, MRZ20, YSZ20, ZYL23]. heatmap [ZS24]. heavily [YZ23b]. heavy [AF22, CC24, Che25, DM24, GW20, Häg20, KK20, LPŠ21, LMS22, Nak22, DMPP23, SYM25, XCDS22]. heavy-tail [Che25, DMPP23]. heavy-tailed [CC24, Häg20, KK20, LPŠ21]. height [Che24b, Mui20]. helicity [Gok25]. heterogeneous [ASKI24, Bat21, CX23, PF23, WC21, ZLS24]. heteroscedastic [ZCP21]. heteroscedasticity [Klo21]. Heyde [Zha21b]. Hidden [MGP20, LH25, Tra21a]. Hierarchical [CPD20, LYY21]. High [MW20, AA21, BP24, CZLC20, CLP22, Din20, DLZZ22, GZ24, GT20, HY23, HNP20, KP24, LT23, LW24a, LWM22, LZMZ20, LBL^{+}22, LZ25, Lyu20, QTLT20, San24, Sel24, Tak24, TLX22, Tre20, WX21, WjWbK22, ZY20, ZPC21, ZF24, ZG24, ZGZF21, ZGQC20]. high-dimensional [AA21, BP24, CLP22, DLZZ22, GZ24, GT20, HY23, HNP20, KP24, LT23, LWM22, LBL^{+}22, LZ25, QTLT20, TLX22, Tre20, WX21, ZY20, ZPC21, ZGZF21]. high-frequency [Tak24, WX21]. high-level [ZGQC20]. High-probability [MW20]. Higher [CX24, WL21a, LPŠ21, Lig23, MO23a, Wei24]. Higher-order [CX24, WL21a, LPŠ21, MO23a, Wei24]. Hilbert [CM23, DMN24, FS20a, MN20]. Hilfer [AZ23, KKC^{+}24]. histograms [Kro22]. Hitting [Col23, HX23, Lee23, Nis24, PCPS24, XFZ^{+}21, XZYC20, YKK20, ZX21]. Hochberg [Izm20]. Hoeffding [AKW22, CL22, Das24, Lig23, MRG22, SŠ23a]. Hoff [Kab24]. Hogg [CL20]. Hölder [HW24, JLZ24, KP21, LZ24b, Yan25a, Zha21a]. holes [AAD24]. homogeneous [BFMS22, FLH23, HK21, Kre22, LL23a, LH22, Tem23, ZYL23, ZL20].

Homogenization [Yan25a]. Hopfield [Pir24]. Huber [XXZ23]. hulls [FK21]. Hunt [Nis24]. Hurst [Bib20, Lee21a, MR24]. Hüsler [HPN22]. Hyers [AM21]. Hyperbolic [Ma21, FS21b, Jur23]. hypercontractive [NV24]. hypercube [LYL22, SWZ20]. hypergeometric [Nał21]. hypergraph [Yua21]. hyperplane [MBR22]. hyperplane-truncated [MBR22]. hyperspheres [DNTV22]. Hypothesis [DPSR20, QTLT20, Ais23, LT22a, MLC25, NS23b, PZ20, SZ24]. hypothesis-testing [PZ20].

i.i.d [GM21, KSH24, LMS22, ML24]. identically [BDPR21, Jas20]. identifiability [Lee21b, TSG20]. Identification [KK22b, Klo21, Mao22, CHHS25, CT21, HJKL21]. identify [Ste23]. identity [GW21, QTLT20]. II [Gir20, HX21a]. iid [JZ23]. images [Fal22]. imitation [Li24a]. immigration [HLX22, KW21, Kev24, LRW24, LL23b, MX22, Mit21, RS24]. implicit [RTT21]. impropriety [DR21]. Improved [Bud22a, Pin22, CL22, KC22]. improvement [LM23]. Improving [AAD25, BP21, FC23, Lig23, LQZ24]. impulse [Ahm22]. impulses [CB20, LZ23b]. impulsive [LZ23b, LL24d]. inactivity [KBM21, LL21b, MPK22]. INAR [ZK22]. inclusions [AZ23]. Incomplete [MPD20, DP23, GM24, KSW22]. Incorporating [Bab20]. increasing [GS22b, Lan21, Lan23, LL20a]. increment [Xu23]. increments [LPŠ21, NP20, Sel24, XC22b]. Independence [LRS20, BHS24, CD20b, DS25, DMN24, GLL23b, Kel24, Lau24, RG20, Sai21, ZZ24a]. independent [CWW20, CL22, Dou22, GL21b, Jas20, JB21, KSW25, Lev21, LL23a, Li24b, Myr21, NLS24, Pin22, WWL20, ZZ25]. Indeterminate [ITM24]. index [BPB24, BGV20, KDN21, KR24, KC22, Kub20, LWZ25, MR24, YZ22]. indexed [YB20a]. indicator [KK24, Suz18, Suz20]. individual [DMS20a, HH20]. individualized [CX23]. individuals [TL23]. induced [Zuo20]. induction [SMRH24]. inequalities [AKW22, BU22, BK22, Bud22a, BV23, ERR23, Egh21, FZ23, FY24, HD23a, Hua23, KXBS21, LL20b, MM22, MT21c, ST21, YSZ20, Zha22b, ZZ25, Zha20b]. inequality [AJO20, BBM22, BP24, BM20, CL22, DL21, EK20, HLL20b, Jak21, KK25, LS22, LZ24a, LW24d, Lig23, LN20, Naa21, Ols24, Osȩ21, RS22, SŠ23a, SWZ21, Tan24, Xu23]. Inference [WjWbK22, Bic20, BTS23, DH20, Hug24, KR24, Lan24, LCH20, Pas22, PC24, YDB23, ZLMT23, ZLMT24, Zha25, ZZZD21]. infinite [BB22, DP23, Hwa21, Kha22, LL24a, LZ24a, Oka20, SS23b, SYM25, XZ20]. infinite-server [SS23b]. Infinitely [MR23, BLL25, CCH20, GX22, Kaw21, KKT24, Kut21, NS23b]. infinity [FLZ20]. inflated [LP20, PvdHZ23]. inflation [CLP23]. Information [CY25, Bab20, Dem21, DPSR20, KB21, LR22, LZH23, Lig23, LQZ24, NW24, SS25, ZZ24a]. INGARCH [Wei24]. Ingersoll [JX22]. inhibition [BHS21]. inhomogeneous [BG23, CL23, Gan20, Gan21b, Yua21, Yua24]. injections [YMWW25]. Innovated [ZLZK20, ZZZD21]. Input [NMR25]. inspection [KR23]. instantaneous [MX22]. Instrument [KjL24, Mao22]. Instrument-residual [KjL24]. instrumental [CT21]. instruments [KjL24, ZGZF21]. insurance [YJWY20]. Integer [CLP23, BMP20]. Integer-valued [CLP23, BMP20]. integrability [RT21]. integral [BP21, BV20, CFH23, Jak21, Lef24, Sol22, VY20, WS20, XY23]. integrals [FS20b, FS24, VK24]. integrated [CL20, Che24c, MP20b, MP20c, SJZ23, WX21]. Integro [SP22]. Integro-local [SP22]. intensity [LCH20]. interacting [Pir24]. interaction [CZ24, KSX24, MD24, Mao22]. intercept [DMS20a, Mar21a, Shp24]. Intermediate [Fal22]. Intermittency [Gra22]. Interplay [YJWY20]. interpoint [TLX22]. interpretation [NWY20]. Interpreting [Ais23]. interrelationship [NS23a]. intersection [XNZL24]. Intersections [LL24c, CWW20, Vog21]. intershock [CE24]. interval [BSB22, BM20, Hil20, LLM24, ZDF24]. intervals [Bur21, Kab24, TG23, WML22]. intraday [WjWbK22]. intransitivity [Oak25]. Intrinsic [GZH20, KP21]. invariance [CJM20, DM21, Haf23]. Invariant [ZDF24, BT25, SXX20, ZZ24c]. Inverse [GK22, BB23, KKG20, PD21, SPT23]. inverse-Gamma [BB23]. inversion [DS22b]. investment [ZZ24b]. involving [Jon22, Kay24, MSM20]. IPW [LWX21]. IPW-based [LWX21]. irreducible [SŠ23a]. irreversible [BDY22]. isolated [WDG23]. isometric [XZ22]. isonormal [LV23]. isotonic [CLZZ21]. Isotropic [Lu24, AGP20, HC23, LLM20]. issue [Gho23b]. Itô [EK22, Mus22]. items [FC23]. iterated [MP21, Pak20, Ye24, Zha22a].

Jaccard [KC22]. Jack [SH22]. jackknife [MPK22, Per24, ZLWZ20]. Jajte [CS21]. James [BFMS22]. Janson [BPR23, Zha22b]. January [Ano20q, Ano21q, Ano22q, Ano23q, Ano24q, Ano25e]. Jensen [KB21, SWZ21]. Jensen-information [KB21]. Joint [GS22a, FK21, LXC20, NLS24, Wri25]. Jong [Döb24]. July [Ano20r, Ano21r, Ano22r, Ano23r, Ano24r]. jump [CM21a, Gat24, Gok25, JX22, LW24c, Sar20, SCMZ24, Tra21a, WZZ23, XFJN22]. jump-diffusion [JX22, SCMZ24]. jumps [ASS24, Col23, EK22, FLZ20, HW20, Lef24, PSX21, QW21]. June [Ano20s, Ano21s, Ano22s, Ano23s, Ano24s].

K-sample [PSP^{+}25]. Kac [FWY20, SXX20]. Kalman [DS22b]. Kantorovich [DM21]. Karhunen [YB20a]. Katz [Fen23]. Katz-type [Fen23]. Kawada [Mus22]. Kemeny [CH23b]. Kempthorne [JWW21]. Kendall [FS21a, Per24, PvdHZ23, SS25]. Kent [DM24]. Kernel [RGS22, AAE22, BP21, BS22c, CD22, Che24c, CLR20, DLS20, DS24, JSV20, JA21, KD24, Moj21, MK24, Pin21b, FA24]. kernel-smoothed [JSV20]. kernel-type [BS22c, MK24]. kernels [AGP20, HL21a, Nog21]. Kesten [Che20b]. Kesten-type [Che20b]. Khintchine [Yak21]. Kiefer [Naa21]. kind [Dou22, ZJ21]. kindness [LM24a]. known [KB21]. Kolmogorov [Vid21a]. Kotlarski [LZ20]. Kullback [NW24]. Kummer [KKG20]. Kuo [MRG22]. kurtosis [Bud22a]. KWW [Wri25].

labelled [AVA23]. Lack [BKK20, FLW23]. Lack-of-fit [BKK20, FLW23]. Laguerre [SY24]. Lahiri [Sai25]. Lancaster [Che20a]. Landau [Gok25, YS24]. Langevin [BDY22, RT22]. Laplace [Aly20, Bar23, BU24, Ber23b, BTS23, Chi22, SBS20, SBS22]. Laplace-type [Aly20]. Laplacian [MO23b]. Large [Cha21, Cha20, FWY20, GLL23a, Gao24, HD23b, KSX24, LBZ20, MT21c, Nak22, Sel24, WFGR24, ZZ24d, Zuo20, BHS24, Bou20, CM23, CD20a, Che20a, CS21, CC24, CQ25, DDG22, DD23, Fan24, FZ23, FLW22, GL21b, KZ21, LMS22, LZZ24a, LM23, MA22, RT21, SYM25, SM24, Tan24, YS24, YL22, ZL20, ZZZD21]. Large-population [KSX24]. larger [ZAP24]. Lasso [LZZ24a, Vog20, SCGW20]. Last [WWZ20, BG23, IM20, LWP21, LW24b, Rib20]. Last-Success-Problem [Rib20]. latent [DZZ21, Ols24, TSG20]. Latin [LYL22, SWZ20]. lattice [BH23, HK24]. lattices [MFF22]. law [BHS24, Bou20, CÇ20, CD20a, Che20a, CS21, DDG22, FZ23, FS20a, MZ21, MP21, Ost22b, Pak20, RT21, SV20, Tan24, Vid21a, Ye24, BP20]. laws [BP24, CCH20, CM23, GL21b, Kha22, KZ21, LD20, MA22, SM24, Sze20, Zha22a]. layer [HNS23]. Learning [SSL22, DR21, May23, WLZ24, ZLZK20]. Least [GM21, SXY22, Bar23, Hwa21, NS22, SJK22, ZK24]. Leffler [Lev21]. left [LPY24, XX20b]. left-Lipschitz [XX20b]. left-truncation [LPY24]. Leibler [NW24]. Lemma [EH22, LZ20]. length [KM20, MT22, RGS22, Vid21b, YDB23]. length-biased [RGS22, YDB23]. leptokurtic [Bro22]. leptokurtic-normal [Bro22]. less [DAA21, FV25]. Lett [BBM22, BRBB23, Sko20, Suz20, ZCL22]. Letters [ZLMT24]. Level [GYZ22, CLLP23, DLS20, Daw21, HLS24, LKA20, Lia20, LQZ24, YZ23a, ZCL22, ZGQC20]. Level-augmented [GYZ22]. levels [Ais23]. Lévy [MLZ23, Suz20, BCM21, BM20, DMS20b, Döb22, GB21, Gra22, KN22, LWP21, LZ24b, LZZ24b, Lin24, Ma23, MGMSP24, NMR25, Ren24, SJZ23, Suz18, Vid23, Vil24, Wan22c, Yak21, YWY23, Zha21a]. Lévy-driven [NMR25]. Lévy’s [HL21b]. Lieb [BU22]. life [AD20, CPL20, EA21]. lifetime [LL21b, Pek22]. lifetimes [TL23]. Lifshitz [Gok25, YS24]. like [HSWZ21]. likelihood [BCDR23, BHS21, DF24, CL23, DLLY25, Fri21, GS20, GZ24, HY23, IMT24, JK22, LPY24, LLM24, LD21b, LH25, MPK22, Ols24, SD20, SW20, Tak24, Tre20, WC21, WM24, Yu20, YZ23b, ZLWZ20, ZL20, Zha20b]. likelihoods [DS22b]. Limit [BB22, LD20, SD22b, TZT21, Wan22a, BC24, Chi22, DLS20, DFG23a, DS20b, FM20, Gao21, GR21, IK21, Jan24, Jur21, KMS21, KRN23, LZ21, LZLX25, Liu24, MS23a, Mar21b, PSZ24, QHK25, SW20, SP22, SPT23]. Limiting [LM24a, Mau24, Riv21, Yua21]. line [Bez21, Sar21b, Tor24]. Linear [MHHJC22, BW23, Buo22, CA20, DT21a, DMS20b, Fen23, FE21, For23, GH23, GS20, GZ24, HT21, HJKL21, Hwa20, Ida20, JQS20, Jur21, KSW25, LL21a, Las23, LZH23, LLMB23, LZMZ20, Mar21a, Mar23, MS20b, May23, NLS24, Ost22a, PSZ24, SM22, SXY22, Sko19, Sko20, SY25, WML22, WW21b, Xie23, XC22a, XK24, Yu20, Zha21b, Zha22a, ZP20]. Linear-in-Means [MHHJC22]. link [MR22, OY23, YZ22]. Linking [MRG22]. Liouville [FS20a, FFGS22, Shi20]. Lipschitz [CB20, CZ23, He24, KKC^{+}24, KXBS21, LX24a, WXY22, XX20b, YZ20, ZJ21]. lives [BKM20a]. Local [AFSW20, Bra23, DB22, KKO22, KRN23, LLMB23, CX24, DFG23a, FE21, GPP20, GPP22, HX21a, IMT24, Kar24, LZZ24b, PSZ24, Pyc21, RK22, SP22, Xie23, XNZL24]. locality [SX24]. locally [FLLW24, FF25, Ida20, LW24d, WXY22, YSZ20]. location [CL20, DNTV22, EK20, PF23, Pyc21, QP23]. location-scale [EK20]. locations [LT22c]. Loève [YB20a]. Log [LLM24, Ari23, Arn21, BdCP23, BVG20, CL23, For23, FS20a, HLL20b, Lan21, Lyu20, MP25, WML22, Zha20b]. log-Cholesky [BVG20]. log-concave [Ari23, Arn21, MP25]. Log-concavity [LLM24, BdCP23]. log-correlated [Lyu20]. log-Harnack [HLL20b]. log-likelihood [CL23, Zha20b]. log-linear [For23, WML22]. log-odds [Lan21]. logarithm [MP21, Pak20, Ye24, Zha22a]. logarithmic [Hua24a, ML24]. logistic [Jur23, MLC25, Ost22a, Pyc21, RA23]. Lomax [FRR^{+}20]. Long [KSW22, Zha22c, KKI21, XZ20, ZPC21]. long-memory [ZPC21]. long-range [KKI21]. Long-time [Zha22c]. longest [FFS24, LM23]. longitudinal [GZ24, LXC20, TSG20]. Lookback [IKCSKH25]. Lorentz [KO21a]. Lorentz-norm [KO21a]. loss [FHM21, GVL20, Hwa20, Lee22, Nie22]. loss-based [GVL20]. Lovell [Bas24, Din21]. low [Zha25]. low-frequency [Zha25]. Lower [Che24a, DSH24, YKK20, Fro21, HS22, KP21, LKA20, LQ22, PSZ24, Pin21a, ZCL22]. lower-order [LKA20, ZCL22]. LS [MT21c, MY24]. LSD [LZH23]. LSE [YLW22]. Lugannani [Are24]. Lyapunov [Pir24].

m [YL22, BDBB20]. M-Vine [BDBB20]. MA [YLW22]. Machlup [LG24]. madogram [Zou23]. madogram-type [Zou23]. magnetization [Lu22]. main [LCF21, MD24]. majority [Li24a]. make [FD21]. Malliavin [Suz20, Suz18]. Mallows [MGD20]. Manhattan [BH23]. manifolds [BT25, CY21, KP21, KY22]. MANOVA [LZ25]. manufactured [FC23]. many [WM21]. map [CY21, NS21]. maps [DS20a, GW21]. March [Ano20t, Ano21t, Ano22t, Ano23t, Ano24t, Ano25f]. Marchenko [Fak22]. Marcinkiewics [Bou20]. Marcinkiewicz [DDG22, Sze20]. Mardia [Bud22a]. Marginal [For23, DF24, CJ23, DM21, DL24, DH22, Ses24, TLW22]. margins [Hug21]. marked [Wan22a]. Markov [AFSW20, BS22b, CY25, CM21b, CH23b, Col23, Cru20, EA20, FMR22, HK24, HZY21, HX23, HL21a, HSWZ21, HC22, HW23b, IW20, Jas20, LH25, LNYZ22, LM23, LNN22, MGP20, MX22, Nog21, PSZ24, QK22, Rie22, SŠ23a, Syl22, Tra21a, XFZ^{+}21, XCD20, YL22, ZSK20, ZM23]. Markov-modulated [QK22]. Markov-switching [FMR22]. Markovian [CH23b, SS23b, TL22, Zei21]. martingale [AJO20, FM20, FYY24, KS20a, MT21b, NS21, Ses24, YY20]. martingales [KO21a, KS22, YM24, ZC23]. Maruyama [GWWY22, LZ24b]. mass [Jou22, LT24, WM21]. matching [BCHL24, Has21]. Matchings [Kop24]. Matérn [AAD24, Che24b, Gir20]. mathematical [SLA21]. matrices [CH23b, Din20, HR22, HLS24, Lee22, LZH23, LBL^{+}22, MO23b, MS20b, Mau24, Mod23, NP22, PG20, QTLT20, Riv21, SM22, TVY20, VIVL23, WLL22, ZPC21]. Matrix [Bab20, LGHG23, Nał21, BVG20, CHHS25, DD23, FHM21, GM21, GZH20, LR22, LW24a, LWM22, LM24b, MR23, SZ24, WX21, WL21a, WZ24c, ZG24]. Matrix-variate [LGHG23]. max [HK21]. max-stable [HK21]. Maxima [MS21, BK24, HR20, KK24, Sun20, TZT21]. Maximal [FZ23, FY24, LL24c, Fri22, MM22, ST21]. Maximin [GYZ22, SWZ20]. Maximum [BHS21, Tre20, BNO21, Bar21, Che24a, Fen23, FYY24, GS20, GPP22, JZ23, KK22b, KSX24, LPY24, LT22b, LZ22b, MN20, RG20, TLX22, Vid20, WZ24a, WWL20, Yu20, YZ23b, Zha22a, ZL20]. Maxwell [BK23, GX22]. May [Ano20u, Ano21u, Ano22u, Ano23u, Ano24u]. McDiarmid [Zha22b]. McKean [JLZ24, LMSY22]. MCMC [CBK22]. Mean [Sai25, BNO21, BH23, BKM20a, CDP24, CS24, CZ23, DL21, DSH24, EH22, FF25, Gao24, Gir20, HWX22, He24, JW24, KBM21, KSX24, KP23, LXC20, MN20, MPK22, NLS24, Pek22, RG20, SRK22, VBS24, YDB23, ZF24]. mean-covariance [KP23, LXC20]. mean-field [FF25, HWX22, KSX24, SRK22]. mean-reflected [He24]. meander [IO21]. Means [MHHJC22, BLK20, BB22, CD22, CBK22, CD20a, HNP20, Kab24, SM24]. measurable [VM21]. Measure [WK21, AAD25, BL20, Fak22, GSL20, Gzy21a, HDC24, LPŠ21, Rad23, SY24, XJG21]. measurement [BKK20, Che21, Gir20, SBS20]. measurements [BH20]. measures [ALO22, BLL25, Gho23b, HSWZ21, Kar24, KB21, KdS23, KS20c, LMZ21, MT21a, MP25, MR22, NS23a, Oka20, PSP^{+}25, Rie22, SMRH24, ZZ24c]. mechanisms [AZvdH23]. Median [Sub21, Zuo20]. mediation [For23]. meets [Zha22b]. Mellin [GK22]. memory [GS22b, Hah22, WW23, ZPC21]. Method [BH20, AZvdH23, Bar23, BS22c, Dal24, FR20, Ful24, GWWY22, GW20, Gra21, JWW21, KK22a, KKG20, LXZ20, MYX22, MEA21, MO22, Pas22, ZGQC20]. Methods [FD24, BTSL21, CJ23, Lee20, MST20, RTT21]. metric [BVG20, CY21, nHL21]. Middle [JB20]. Midzuno [Cha20]. mild [MRZ20]. milder [HIS24]. Milstein [RTT21]. minima [Gap20, Sel24]. minimal [DS22a, HD22]. Minimax [GP23, ZPC21, JW24, Li24b]. minimaxity [Bic20]. Minimizing [HP25, PK24, CDP24]. minimum [BLK20, Dem21, Dou22, GPP22, KK22b, LKA20, LZ22b, SS25, TZT21, ZCL22]. Minkowski [Hua24a]. mismeasurement [SPG22]. misrepresentation [ZW23]. miss [Buo22]. missing [LT24, LP20, LWX21, WW21b, ZP20]. Mittag [Lev21].

Mittag-Leffler [Lev21]. Mixed [HLS24, YY20, BW23, IMT24, KZ21, RK22, Sai21, SV20, TSG20, WXY22, ZYL23]. mixed-effects [IMT24]. Mixed-level [HLS24]. mixing [BJ22, DM21, Haf23, HZY21, nHL21, LNN22, MP20a, PSZ24, WYWS21]. Mixture [HA24, ASV24, BKZ21, DM24, Egi22, GVL20, HZZZ21, LL24a, NLS24]. mixtures [BKB24, LLL^{+}25, XLNQ20]. MLE [FMR22, WML22]. Möbius [MRG22]. modal [AAE22]. mode [RL20]. Model [DF24, CZLC20, ART20, ATT21, BKK20, Bar23, BDK21, Bez21, BW23, Bro22, CE24, Che24c, CDLZ25, DH20, DM24, FLH23, FD24, FFGS22, FLW22, Ful24, GWWY22, GLL23a, Gao24, GLLOS24, GHH24, HZZZ21, Has21, HT21, HPN22, HJKL21, Hwa21, IKCSKH25, IMT24, KSAA24, Ken21, Klo21, Kri22, Kry22, LWZ25, LL21a, LM24a, LBZ20, LZH23, LGHG23, LZZ24a, LW24b, Li24a, LW24c, LX24a, LCCZ24, LCF21, LZ23a, LH22, LXC20, Lu22, LP20, LWX21, Lyu20, MD24, MHHJC22, MDS25, MRD22, MLC25, Pen23, PD21, Ren24, Sai21, Sha25, Shp24, SW20, SY24, TLW22, Tra21b, WLS20, WG23, WFZ25, WW21b, XXZ23, XZYC20, XOC21, XH22, XC22a, YZ22, YZ23b, ZY20, ZZF22, ZW23, ZZ24b, ZLS24, vNBR21]. model-based [Bro22]. Model-free [CZLC20]. modelling [DML24]. Models [MGP20, ASV24, Ari23, BCDR23, BKZ21, BDBB20, CLP23, CY25, Das24, DR21, DLZZ22, FLW23, For23, GS20, GM22, GZ24, Gir20, GVL20, HIS24, HH20, HT23a, HJKL21, Hug21, HA24, Ida20, IW20, JQS20, KM22, KRN23, LCH20, LL24a, LM21, LTZ22, LH25, LYC20, LZMZ20, LYY21, MEA21, MGM21, Mar21a, MST20, MT21c, OK21, Ols24, Ost22a, PD20, RA23, SŠ23a, ST21, SCMZ24, SY25, TSG20, WML22, WW21b, Wei21, WH24, YJWY20, YLW22, Zei21, ZLWZ20, ZG24, ZLZK20, ZP20, ZZZD21]. Moderate [HX23, WZ24b, XCD20, BM22, CL23, DFG23b, Fan24, FWY20, MY24, Sun25]. modes [Arn21, SN20]. modified [BG23, BS22a, LXZ20, LZZ24a, LBZH20, MSM20]. modulated [QK22]. modulus [HC23]. Moivre [Chi22, Sze22]. Molchanov [Wei23]. Moment [EH22, BTSL21, BS22c, Fen23, Gho23a, ITM24, JB25, Kap22, KP24, LXZ20, LTZ22, Lyu20, MP24, Mui20, XK24, ZZ25]. moment-based [JB25]. moment-dependent [KP24]. moment-recovered [MP24]. Moments [FP22, KW21, WW21a, Ahl22, DP23, Egh21, Fri22, Gau23, HW23b, JLL20, JB25, KSW20, LPŠ21, Lig23, NWY20, Sun20, WL21a, WM21, Wei24, YMWW25, Zha22a]. monetary [MR22]. monitored [LN20]. monitoring [AA21]. monotone [Dal24, FLLW24, FF25, LW24d]. Monotonic [Mar21b, SB21]. Monotonicity [DP22, Pin20, Pin21c, Pin21b, Tzi20, DZ21, KKO22, KjL24, PD20, Vid23, Yan25c]. Monte [CJ23, ZYLB22]. Moran [Ful24]. Morrey [YY20]. most [BG23, BK23, Gho23b, Hua24a, Jan21]. motion [Bai20, Bib20, BV20, CÇ20, Che24a, Daw21, Fal22, GS22a, GWWY22, He24, HD23b, HCW22, HL21b, Hua24a, IM20, KN22, Lee23, LZ22a, LZ22b, MT21b, Mar22, MO23b, MMB20, Shi20, Tak24, Tal20, Tou21, Val23, WSM24, XC22b, Xu23, XY23, YY22, YB20a, ZJ21, ZHD24, ZX21, ZL25]. motions [Lou23, NS22, PCPS24, SXY22]. motivated [GTH24, LXH23]. movements [CCY20]. moving [BSST20, GM24, XK24]. MSO [MZ21]. Multi [He24, SD22a, AY20, JL25, KjL24, Ses24, WLZ24]. multi-armed [AY20]. Multi-dimensional [He24]. multi-marginal [Ses24]. multi-scale [JL25]. multi-task [WLZ24]. multi-valued [KjL24]. multidimensional [CZ24, CN24, DB22, Kri22, MT21b, RL20, Shp24, WL21b]. multifractal [KN22]. multifractional [Bai20]. multiinformation [MG20]. multilayered [MHHJC22]. multinomial [BDK24, Ber23b, BAB24, JB20, LLM24, SN24]. multiparametric [LBB20]. multiple [Ais23, BLK20, BKB24, CD20a, DDR20, Hah20, HC22, HP25, RZ22, Syl22, WX21]. multiple-outlier [BKB24]. multiplicative [HLL20b, Jon22, LCH20, ZWW22]. multiscale [GGX21]. multitype [KW21, Sto20]. Multivariate [HK21, Ma23, ZW23, ZX25, AH23, BLK20, Bro22, Cha24, DT21a, DMS20b, ERR23, GS20, GZH20, Häg20, Hah22, IW20, LRS20, MBR22, MG20, MGP20, Naa21, PF23, RB24, SN20, SBS22, WM24]. mutual [ZZ24a]. Muzy [FS20a]. Myersonian [CLZZ21].

Nadaraya [XOC21]. Narayana [FR20]. NBUE [Lan24]. near [ART20, Gir20, MS21, WFGR24]. near-efficiency [Gir20]. near-maxima [MS21]. nearest [RG20]. Nearly [Li24b, Gho23b, RS24, SWZ20]. necessarily [Gir24]. Necessary [NV24, CT21, HJKL21]. needed [DS21, RK23]. negative [Ber23b, FP22, FLZ20, GTH24, Hua24b, KK24, LWP21, Mah23, MM22, MGMSP24, Ren24, SP24b, XX20a, YLP23]. negatively [Tan24]. neighborhood [MY24]. neighbors [RG20]. nested [PZL24]. network [DY24, FLH23, LH22, MHHJC22, SJSL21, XXZ23, ZLS24]. networks [BS22a, DNH24, HNS23, HD22, SSL22, WW21a]. neural [DNH24, HNS23, HD22, SJSL21, SSL22, WW21a]. neutral [AM21, CB20, YY22]. Neyman [Ais23]. NMAR [MK24]. NN [NAV21]. no [Dem21, DMS20a, FLZ20, Shp24]. no-information [Dem21]. nodal [Vid21b]. node [LD20]. nodes [Ste23, WDG23]. noise [AZ23, Cha21, Gok25, HW24, HLL20b, HL21b, IMT24, Kaw21, Kum22, LL20b, LZH23, LZZ24b, Liu20a, MS23b, TWW21, WZZ23, YZ20, ZYL23, ZWW22]. noises [HW23a, KSX24, SJZ23, ZYS20]. noisy [WX21]. Nominal [Gan21b]. Non [BM22, Bez21, CLZZ21, KDN21, KP22a, LT22a, Mar23, TLW22, AJO20, BSB22, BL20, CC24, CLP22, CH23b, CB20, CZ23, DT21a, DMS20b, DM24, EK22, Haf23, Has21, He24, HD23b, HT21, IM20, IM23, Jas20, Jur21, KKC^{+}24, Kaw23, Kre22, Kum22, LL21a, Lar22, Lee22, LT23, LX24a, LZZ24b, MS20a, NNO20, Nas20, PG20, SŠ23a, TP22, TL22, WXY22, YLP23, YZ20, ZYS20, ZJ21, ZL20]. non- [EK22]. non-anticipative [Lar22]. Non-asymptotic [LT22a, LT23]. non-autonomous [TL22, WXY22]. non-Bayesian [Kaw23]. Non-central [BM22, Nas20, TP22]. non-centralized [CC24]. non-conventional [BSB22]. Non-convex [CLZZ21, HD23b]. non-degenerate [CLP22]. non-diagonalizable [PG20]. non-diffuse [BL20]. non-existence [NNO20]. non-Gaussian [Kum22]. non-homogeneous [Kre22, ZL20]. non-identically [Jas20]. non-irreducible [SŠ23a]. Non-linear [Mar23, DT21a, DMS20b, HT21, Jur21, LL21a]. non-Lipschitz [CB20, CZ23, He24, KKC^{+}24, LX24a, YZ20, ZJ21]. non-local [LZZ24b]. Non-marginal [TLW22]. non-Markovian [CH23b]. non-negative [YLP23]. non-null [IM20]. Non-parametric [KDN21, KP22a]. non-regular [Has21]. non-regularity [HT21]. non-spherical [DM24]. non-square [Lee22]. non-standard [IM23]. non-stationary [Haf23, ZYS20]. non-stochastic [MS20a]. non-symmetric [AJO20]. Non-triviality [Bez21]. noncentral [KSW25]. nonclassical [Vid21a]. nondifferential [SPG22]. nonextensive [Gho23a]. nonhomogeneous [XCD20]. noninstantaneous [Ahm22]. nonlinear [AZ23, CN24, FMR22, Luo20, MLZ23, MT21c, WZZ23]. nonlocal [Ahm22, AZ23, CZG24]. nonnegatively [Sel24]. Nonparametric [BS22a, BCHL24, BS22c, DH20, DS24, DH22, Efr24, FE21, KY22, Lan24, LD21a, Mar20a, Mar21a, SBS20, ZS24, ZHD24, AY20, BKM20a, Bra23, DNTV22, Jou22, Klo21, OY23]. Nonreconstruction [ZY20]. nonstationary [MP20a]. Nonsymmetric [Hua23]. norm [KO21a, Lee22, SMS20, WZLL22]. Normal [KP22b, ASV24, BNS23, BLK20, Bro22, CD20a, DB21, DB23, FS24, GLS25, HY23, KP23, LX24b, MBR22, MG20, NLS24, PZ20, RB24, XP20, XCDS22, ZZ24a]. normality [BNO21, CBK22, Izm20, JK22, MN20, NS22, Pac20, RK22, Sko19, Sko20]. normalized [Zha22a]. norms [Cha24, LG24]. Note [Ano20x, Pas22, SX24, VIVL23, Bat21, BT25, CJM20, Cin22, DT21a, DMS21, DL24, FHSP21, GS20, HY23, Ida20, IW20, JSV20, JP20, KKI21, Lee21b, Lin24, LH22, LP20, MBR22, MP25, MGMSP24, Mar20b, MA22, MMB20, Moj21, Mus22, Nog21, PD20, RK22, Ren24, RT21, SP24a, SMRH24, Sep20, TSG20, Tap21, Vid21b, Vog21, XZ23, Zha20b]. Notes [GLL23b, Gir24, Kop24]. novel [Zha24]. November [Ano20v, Ano21v, Ano22v, Ano23v, Ano24v]. nuclear [WZLL22]. null [AD20, CS24, IM20, Zha23]. number [AVA23, Arn21, BL20, BV23, Che24b, Col23, DS21, FD21, FK21, GVL20, GM21, GS21, LH25, LD20, LQ22, NW24, OH21, WDG23]. numbers [BHS24, Bou20, CM23, CD20a, Che20a, CS21, DDG22, DAA21, Dou22, FZ23, FR20, Gan21a, GL21b, KZ21, MA22, RT21, SM24, Tan24]. numerical [BTSL21].

Objective [FRR^{+}20, ZCP21]. observables [Li24b]. observations [Aug20, DAA21, HX21b, KDN21, LGHG23, MRD22, NMR25, Tak24, WX21, Zha25]. observed [KS20b, MO22]. obtain [DS21]. obtained [Gat24]. obtaining [ZSK20]. occasional [Hug21]. occasions [FD21]. Occupation [Che24d, LWP21, CÇ20, Col23]. occurrences [DS20a]. Ocone [ZC23]. October [Ano20w, Ano21w, Ano22w, Ano23w, Ano24w]. odds [Lan21]. OFBM [Lee20]. off [DSH24]. offspring [FJ24]. omega [Pyc21]. omega-square-type [Pyc21]. One [HSS22, JLZ24, Roz21, Bez21, CQ25, HWW23, HNS23, KK20, LL23a, Lin24, LZ25, MGM21, Vid21a, Wan22c, WM24, ZCP21, Zha21a]. One-dimensional [HSS22, JLZ24, Bez21, MGM21, Zha21a]. one-layer [HNS23]. one-sided [KK20, Lin24, WM24]. one-way [LZ25, ZCP21]. Online [Hah22, MP20b, MP21, NN21]. Onsager [LG24]. operator [CP22, Lee21a, Lee22, LV23, Li23]. operators [EA20, FPBK20, HSWZ21, SJK22]. opinions [KVW24]. opposite [RS22]. Optimal [CS24, CA20, DMS20a, Gap20, Gho23a, KVW24, Lee22, LMSY22, LT23, LYY21, LQ22, MD24, MDS25, Sar21a, Wri25, YZ23a, BPB24, CPD20, CML20, CT21, DMS21, GW21, GP23, HZZZ21, HH20, HD22, HT21, HA24, Ida20, KSAA24, KD24, KS20b, LL21a, LLL^{+}25, Lin24, LYC20, NS21, OH21, Ren24, SB21, Ses24, Shp24, SDH20, Wri20]. optimality [NW24, Pyc21, RB24]. Option [LW24c, HCW22, IKCSKH25, KKS24, LW24b]. optional [Ber23a, Mar20b]. Order [CML20, AEAC21, Aly20, AH23, Aug20, BdCP23, CX24, Che24c, DS22a, DP23, Fri21, Fri22, FS21a, Gir24, GCW24, HZZZ21, HA24, Jas20, JH22, Kap22, KR24, LBB20, LPŠ21, Lev21, LKA20, LT22c, MO23a, MPK22, PG20, SBK20, Tra21b, Vir21, WL21a, Wei24, WWL20, Yu22, ZCL22]. Order-of-Addition [CML20]. Ordering [BKB24, AEMZ24, MS20a, PCPS24, WC21, YLP23]. orderings [AEAC21]. orders [GN23]. Ordinal [FEVC21, ATT21, WK21]. ordinate [ART20]. Orey [Kub20]. origin [AGP20, IM23]. Orlicz [YM24]. Orlicz-amalgam [YM24]. Ornstein [Bar21, CL23, HX21b, Lef24, Li23, SJZ23, YB20b, ZYS20, ZZ24b, Zha25, ZL20]. orthant [ERR23]. orthogonal [Ber23a, JWW21, LYL22, PZL24, QHK25, SWZ20]. oscillating [LMSY22]. oscillator [CR25]. oscillators [Gzy21a, Gzy21b]. other [SM24]. Ottaviani [BM20]. out-of-order [JH22]. outcome [LSFZ24, Mao22]. outlier [BKB24, OGR24]. outward [Sin23]. overall [NP22].

packing [LQ22, XLNQ20]. pair [JQS20]. pairs [CLP22, DH22]. pairwise [MGM21]. Paley [WLL22]. panel [CA20, Hwa21, LZ23a, MEA21, ZZF22]. pantograph [KKC^{+}24]. paper [ZCL22]. parabolic [LZ22b, LW20, Lyu20]. parallel [ASKI24, WC21]. Parameter [KK20, MRD22, MLC25, SJZ23, ZYS20, BK23, BW23, CL20, CY25, Lee20, LTZ22, MST20, DMPP23, SCMZ24]. parameterization [YZ23a]. parameters [Che20a, HX21b, Hua24b, KK22b, MR23, QP23, SL25]. Parametric [ZCP21, BKK20, KDN21, KP22a, LZ23a]. paraproducts [KS20a]. Pareto [MA22]. Pareto-type [MA22]. Parisian [Kry22, Ren24, YMWW25]. Parseval [GW21]. Part [MP24]. partial [BT25, CC24, FLLW24, Fen23, LLM24, LMS22, ML24, NOH21, RB20, SQT24, YKK20, Zha22a, ZP20]. Partially [Lia20, BAB24, KS20b, LZMZ20, MO22, SY25]. particles [CQ25, KSX24]. partitions [DS22a, LL24a]. partly [Tan24]. Pascal [LXH23]. passage [CFH23, Lef24, LW24b, Rat21, Vid23, WWZ20, ZX21]. Passing [BH20]. past [BKM20a, Kay24]. Pastur [Fak22]. path [Kaw21, MT22]. paths [Gan20, JLL20]. pathwise [DF24]. Pattern [DS20a, BBL22, CLLP23, MGM21]. PCA [MW20]. Pearson [Ais23, EMS21]. Penalized [LH25, HP25, XX20a, YZ23b]. penalties [SMS20]. Peng [GLL23b]. percolation [Bez21, JLL20, LX24b, WWZ20, XZ20, XZ22]. perfectly [LEJT22]. performance [Moj21]. periodic [ZHD24]. periodicity [WjWbK22]. Periodogram [ART20]. permutations [GI24]. perpetual [LW24b]. perpetuity [IK21]. perspective [PZ20]. perturbation [LMSY22]. perturbations [LX24a, LNN22, PG20]. perturbed [Kum22, Liu20a, MYX22, Yan25a]. phase [JZ23, KP24, LMN24, Nał21, XZ20]. phase-type [Nał21]. phenomenon [AM19, Bar21, JQWD23, Wei23]. Pickands [Zou23]. piecewise [MRZ20]. piecewise-constant [MRZ20]. planar [DS20a, Gat24, Mui20, Vid21b]. plot [MPD20]. plus [JWW21, LZH23, MD24]. Poincaré [BU22]. Point [GI24, LT22c, CE24, DK22, FK21, Hir21, HT23a, Liu20a, RZ22, YZ22, Zha24]. points [CS20, Che24b, Mui20, WM21, Yu22]. pointwise [TZT21]. Poisson [BRBB23, Ahl22, ASS24, BB23, BK23, BM22, BPB24, BRBB14, DW20, FK21, GX22, GLLOS24, Hwa20, KK22b, KdS23, KZ21, Kre22, LXH23, Li24b, LX24b, LP20, PG21, Pin20, RB20, SY24, ZW23]. Pólya [FM21, MZ20]. Polygon [Lar22]. polygons [HMRT21, Ver22]. Polynomial [CM21a, Bra23, DMS20a, DMS21, LZLX25, LL24d, Shp24, WG23]. polynomials [SH22, WW21a]. pooling [KVW24]. Population [LQZ24, KSX24, LYY21, Pas22, TWW21]. Population-level [LQZ24]. populations [Bat21, DAA21, FC23]. portfolio [Sar21a]. posedness [CZG24, CZ23, Gok25, HT23b, ZWW22]. position [BG23]. positive [AGP20, FPBK20, Gap20, Kar24, MO23b, Sun25]. positively [ERR23]. post [PC24]. post-selection [PC24]. Posterior [DR21, HIS24, RRL20, Bic20, DM20]. Potentials [KdS23, Hir21]. Power [Pak20, Bud22b, JA21, Ma23, Ost22b, Sar21a, WZGY21]. powered [XP20]. powerful [BK23]. Prabhakar [GB21]. Pratelli [Osȩ21]. precedence [GN23]. Precise [CC24, FLW22, Häg20, Lyu20, YL22, BW23, Roz21]. precision [DD23, LWM22, ZPC21]. predictability [Mer23]. Prediction [ASS24, LLWZ20, SV20, HH20, Li24b, SBK20, SD22a]. Predictive [Has21, JA21]. predictors [JQS20]. prequential [EA21]. presence [PS21]. Preservation [LL20a, BK24, Pin21b]. preserves [Lau24]. prevalence [PK24]. prices [KSAA24]. Pricing [IKCSKH25, KKS24, CDLZ25, HCW22, LW24c]. principal [JA21, Yu22]. principle [Cha21, CB20, DM21, Haf23, HD23b, KKC^{+}24, QK22, Rad23, Vid21b, WZ24b, YWY23]. principled [vNBR21]. principles [FWY20, YS24]. prior [Egi22, GVL20, LR22, SSL22, vNBR21]. priors [Has21, OK21, RA23]. Probab [BBM22, BRBB23, Sko20, Suz20, ZCL22]. Probabilistic [CZG24, CY21, HSZ21]. probabilities [BMP20, Bat21, Cha20, Cin22, Fro21, Häg20, KSMF20, LCCZ24, MKF20, Nie22, Syl22, Tzi20, Zaj20]. Probability [QX23, ZLMT24, ZZ25, Ari23, BP21, Bud22a, Fak22, FK21, Fri22, Gho23b, Has21, HR22, HP25, JZ23, Jan21, JP20, Jou22, Kri22, Kry22, LZ22a, MW20, Nis24, Pin21a, Sha25, SN20, VY20, Ver22, WYWS21, ZS24]. problem [Buo22, DR24, Dem21, DIY23, ITM24, Lef24, NS21, NP20, Ren24, SB21, SBK20, SN20, WM24, Zha20a, Rib20]. problems [BDY22, BS22b, CDP24, Gap20, Lin24, YDB23, ZF24]. procedure [Hwa20, Izm20, NAV21]. procedures [Wan22b]. Procesi [CZG24]. Process [MS20b, AA21, AAD24, BB23, BM20, BSST20, Cin22, DFG23a, DB22, DFG23b, FLW23, FK21, FYY24, Gao21, HFC22, HX21b, Hwa20, KR23, Kaw23, KZB23, KN22, KMS21, Lef24, LZ24b, LT22b, LL23b, Mah23, MZ20, MO22, Mit21, Osȩ24, PG21, Pek22, PS21, RK22, SS24, SPT23, SP24b, Sto20, SP22, SXX20, Vid20, WLS20, Wan22a, Yan24, Yan25c, YB20b, YWY23, ZK22, ZK24, ZH22, ZZ24b, ZX25, ZX20, ZZ24d, MLZ23]. processes [ASS24, AD20, BCM21, Bar21, BM22, BHS21, Buo22, CM21a, Che24d, CL23, DMS20b, Dou22, EA21, EA20, FMR22, FJ24, FLZ20, GB21, GI24, Gat24, Gra22, GK22, Gzy21b, HK21, HZY21, Hir21, HLX22, HW23b, IM23, JX22, KSMF20, Kaw21, KW21, Kev24, KM20, KdS23, Kre22, Kub20, Lar22, Le22, LWP21, LZLX25, Lin24, LT22c, LRW24, Liu24, Ma23, MC21, MGMSP24, MX22, MP20a, MY24, MR24, NN21, NP24, Nis24, NV24, Pir24, QX23, RS24, Rat21, RB20, Sar21b, Sel24, SJZ23, SV20, TT23, TL23, TZT21, VP20, Vid23, Vil24, VK24, Wan22c, WWL23, WW23, WFGR24, Wei24, XNZL24, XK24, Ye24, ZYS20, Zha21a, ZM23, ZZ24c, Zha25, ZL20]. Product [BHS24, ERR23, GLS25, KK25, KSW22, RS22, SH22, Val23]. products [GK20, HL21a, Oka20]. progeny [BG23]. program [SLA21]. Projection [Mar22, GCW24]. projective [WZ24a]. proof [Döb22, FMR22, Kre22, Yak21]. propensity [BCHL24, yCLjL23, OH21]. proper [Nie22]. Properties [CJ23, Ahm20, ADM24, BS24, BLK20, BK24, BRBB14, BRBB23, Cha20, DP22, FA24, FYY24, GH23, HNS23, KBM21, LRS20, MX22, Pin20, Pin21c, Pin21b, RRL20, Riv21, SMS20, ZYY22, Zuo20]. property [CN24, HZY21, HSWZ21, nHL21, Jas20, KR23, Lan23, TL23, Zha21a, ZC23]. proportional [Cha20]. proportions [Aug20]. Proximal [Pac20, ZLMT23, ZLMT24]. pseudo [AH23, GL21b]. pseudo-independent [GL21b]. Pseudodifferential [EA20]. Pseudoprocesses [MO23a]. psi [PSZ24]. psi-mixing [PSZ24]. publication [AZvdH23]. Publisher [Ano20x]. pure [WZZ23]. put [LW24b].

quadratic [HWX22, KKO22, KKH24, Kub20, LL20b, MS20a, WSM24, Zaj20, ZZ25, ZDF24]. qualitative [LLWZ20]. Quantifying [LL24a]. Quantile [Bab20, WZLL22, DLLY25, EK20, FLW23, LQZ24, SCMZ24, SY25]. quantile-based [DLLY25]. quantiles [ALO22, Are24, DLLY25, Fan24, Yan24]. Quantitative [GGX21, LLWZ20]. quantities [Izm20]. Quantization [HMRT21]. quantum [LN20]. Quasi [RRPB21, Tak24, Kha22, Kut21, SW20, ZM23]. quasi-infinite [Kha22]. quasi-infinitely [Kut21]. Quasi-likelihood [Tak24, SW20]. Quasi-random [RRPB21]. quasi-stationary [ZM23]. Quenched [GPP22, ZH22]. queue [GW20, SS23b]. queueing [SP24a]. queuing [Zei21].

R [HZZZ21]. R-optimal [HZZZ21]. Rademacher [CD22]. radial [FPBK20]. radially [GW23]. Ramanujan [DZ21]. Random [nHL21, KSH24, NLS24, ZH24, ADM24, AEAC21, AVA23, Aug20, BMP20, BG23, BU24, BH23, BB22, BMT23, BK24, Bou20, Bud22b, BV23, CCY20, Che20b, CX24, CC24, Che25, CD20b, CS20, CL22, Che24b, CH23a, DAA21, DIY23, DT23, DZ21, DS20a, DS20b, Egh21, FD24, FP22, Fri22, FYY24, GK20, GI24, Gan20, Gan21a, Gan21b, GM22, Gao21, GLS25, GR21, GN23, GM24, GM21, GT20, GL21b, GS21, GS22b, HH20, HL21a, HC23, HT23a, HS22, HLX22, Hwa21, IK21, IW20, JZ23, JB21, Jon22, JS20, KR23, Kev24, KK22b, KdS23, KZ21, KS20c, KK24, KSW25, Lee21a, LMS22, LZLX25, LT22b, LL24c, LYC20, LD20, LT22c, LLM20, LDM22, Lu24, LMN24, Ma21, Ma23, MU23, MZ21, MFF22, ML24, MP21, MS21, Myr21, NWY20, Nak22, Pin22, PS21]. random [QHK25, RK22, RRPB21, Rz21, SYM25, SLA21, Sol22, Ste23, SP22, Sun20, SJK22, Sun25, Tan24, Tem23, Tra21b, Tzi20, Vid21b, VBS24, VK24, Vog21, Wai24, WW21a, WW21b, WDG23, WWL20, XK24, YLP23, Ye24, Yua21, Yua24, Zaj20, ZH22, ZZ25, ZX21, ZZ24d]. randomised [Kro22]. Randomized [AY20, BFK21, CHW20, GM22, LS22, Sai21]. randomly [Che20b, Rz21, SM24]. Range [GYZ22, CH23a, KKI21, MA22, XZ20, ZAP24]. Range-fixed [GYZ22]. Rank [Itk24, BP21, Dem21, DZZ21, HDC24, KR24, Sko19, Sko20, ZF24]. rank-based [KR24, ZF24]. ranked [RRPB21]. ranking [Wri25]. ranks [RG20]. rapid [SS23b]. rapidly [Gra21]. rare [LSS22]. Rate [KKG20, RN24, AVA23, CP22, DZZ21, DS20b, FFS24, Fri22, Gao21, GP23, GL21a, HFC22, HW23a, Izm20, JH22, KZ21, Lan21, Lan23, LMSY22, LZLX25, LX24b, MDS25, MP21, RK23, RB20, Sar20, Sar21b, SX24, WYWS21, XFZ^{+}21, ZSK20, Zei21]. Rates [MS23a, Gap20, KP21, QHK25, Rie22, Roz21, Sto20, Wan22c, ZH22]. ratings [KSW22]. ratio [AD20, BCDR23, Bur21, CL23, Din20, DW20, Fri21, HY23, Ma21, MPK22, Oak25, Sai25, WC21, WM24]. rational [Syl22]. ratios [JK22, SW20]. raw [Ahl22]. RDS [LBL^{+}22]. re [Moj21]. re-sample [Moj21]. reaching [HP25]. Real [EA21, Tor24]. Real-time [EA21]. recognition [BBL22]. reconstruction [KS20b, MW20]. record [KR23]. record-values [KR23]. records [Ahm20, BS24, DAA21, SD22b]. recovered [MP24]. recovering [MP24]. recovery [MS23b]. Recurrence [TL23, BMP20]. recurrent [AD20]. Recursive [MP20c, Pri21, Sol22, BS22c, KY22, LD21a, MZ21, WYWS21]. reduction [Che21, DL24, RA23, Vid21b]. redundancy [LL23a]. refined [LS22]. Reflected [Luo20, CZ24, CZ23, FF25, He24, HJO20, HX21b, KO21b, LL20b, LZ24a, Mar20a, Mar20b, Sar20, ZYS20, ZX21, HD23b]. reflecting [QX23]. reflection [DL21, LNYZ22]. reflection-type [LNYZ22]. regime [JX22, LT23]. regime-switching [JX22]. regimes [CT21]. region [Wri25]. regions [Aug20, BLK20, DAA21]. regression [AAE22, BH20, BS22a, CLZZ21, DLS20, DMS20a, DMS21, FLW23, FE21, FV25, GH23, GP23, HH20, JQS20, KY22, Las23, LBZ20, LTZ22, LZH23, LLMB23, LYC20, LZMZ20, LP20, LQZ24, MT21c, MK24, DML24, OY23, Ost22a, Pin21b, PD21, RA23, SBS20, SCMZ24, SY25, Sub21, SCGW20, WLS20, WZLL22, Xie23, Yu20, Zuo20, Bab20]. regressions [CA20, May23, XX20a]. Regular [CCH20, Nie22, CLLP23, HMRT21, Has21, LKA20, ZCL22]. regularity [HT21, HT23b, Liu20a]. regularization [BVG20, WZLL22]. Regularized [ZG24]. regulated [HJO20, Mar21b]. reinfection [WFZ25]. reinforced [Pir24]. reinsurance [CDLZ25, ZZ24b]. Reiss [HPN22]. relapse [WFZ25]. related [ALO22, BFK21, CDP24, CD20a, CJM20, GW23, Gho23b, GK22, HR22, Izm20, MT21a, MO23a, NS23a, NS23b, Zha22a]. relations [BMP20]. relationships [EMS21, Jon22]. relative [FE21, HDC24, LT23, MP24, Sar21a]. relative-rank [HDC24]. relevance [LL24c]. ReLU [BS22a]. remark [KP24, Sai20]. renewal [AD20, DP22, EA21, Jan24, KR23, KSMF20, LBZ20, LCF21]. Rényi [BV23, Gao24]. repeated [BH20, PD20]. repeatedly [LN20]. repelling [SXY22]. replacement [AKW22]. replicated [LGHG23, Lia20]. Replication [Wai24]. Representation [KKH24, YJT20, CY21, Kaw21, KMP25, KSW25, MGMSP24, WDP^{+}20, ZZ24c, ZDF24]. representations [LLM20]. Reproducible [DLZZ22, WLZ24]. residual [KP22a, Kay24, KjL24, LL21b, Pek22]. resistance [Luo20]. resource [SN20]. respect [Yu22, YJT20]. response [Ari23, yCLjL23, Kro22, MK24, Sai21, SCGW20, WW21b]. responses [ZP20]. restricted [Ida20, YDB23]. restriction [AH23]. restrictiveness [Fri21, Fri22]. result [Daw21, Kut21, LMS22, Sto20, ZL20]. results [Ahm20, Bas24, BKB24, HPH24, IM20, IM23, Kay24, Liu20a, NS23a, SPG22, Xie23]. resurrection [MX22]. returns [LCCZ24]. reverse [FY24]. reversible [CM21a, HC22, XFZ^{+}21]. revisited [HIS24, MO23b, Sze22, ZK24]. Revisiting [BV20, JA21]. Revitalizing [Bro22]. rewards [AY20]. Reweighted [XOC21, Zou23]. Ricci [Yua24]. Rice [Are24]. Ridge [KS20b]. Riemann [FS20a, FFGS22, NS23b]. Riemannian [CY21, KY22, KP21]. Right [BG23, Efr24, FJ24, KDN21, LD21b, XOC21, Yu20]. right-censored [XOC21, Yu20]. Right-most [BG23]. rings [JLL20, KS20c]. Risk [BS22b, FLW22, GLL23a, HCW22, KRN23, Kri22, Kry22, LPŠ21, LBZ20, LCCZ24, LCF21, MT21a, MR22, SMRH24, Sha25, SY24, XJG21, YJWY20, ZZ24b]. Risk-sensitive [BS22b]. risks [CZLC20, HP25, LPY24, TLW22, YJWY20]. RKHS [XC22a]. RKHS-based [XC22a]. RMLE [AH23]. Robbins [LN20]. robin [MM22]. Robust [CL20, LTZ22, LWM22, QP23, ST21, ZZF22, Hwa20, LWX21, PF23, WLS20, Bab20]. robustification [AAE22]. robustness [HIS24]. rods [MO23a]. root [KVW24, MY24, NW24]. roots [ART20].

Rosenblatt [AAD25]. Ross [JX22]. rotationally [SXX20]. rotor [KSH24]. rough [FFGS22, HW24, HW23a]. roughly [XZ22]. roughness [HSZ21]. round [MM22]. round-robin [MM22]. rounded [JGA24]. row [ZJD20]. rows [GM21]. Ruin [Sha25, Häg20, Kri22, Kry22, LCCZ24, Ren24, SY24, YMWW25]. rules [CX23]. run [CQ25]. running [Gap20, Rat21]. runs [FFS24, LM23]. runtime [Hah20].

Saddlepoint [Jou22, Are24]. SAI [LL20a]. sample [Cha20, Din20, DD23, Egi22, Fan24, FV25, HR20, Kaw21, LEJT22, MA22, Moj21, PSP^{+}25, Sai25, Wri20, YDB23, Zuo20]. sampler [JH22]. samples [BPB24, GM24, Hil20, WYWS21, YDB23]. sampling [Bar21, BL20, BPB24, Cha20, FD21, Lee20, RGS22, Riv21, RRPB21]. Samuel [XH22]. sandpiles [KSH24]. SAR [LWX21]. Scalable [WLZ24, ZLZK20, ZZZD21]. scale [CS24, EK20, Hua24a, JL25, Lee20, MGMSP24]. scaling [Jur21, Kre22, Lee21a, Liu24]. scan [SN24]. scatter [DM24]. scattering [KM20]. sceneries [CD20b]. Scheffé [HZZZ21]. scheme [Cha20, HW23a, LZ24b, Riv21]. Schmidt [DMN24]. Schoenberg [AGP20]. Scholes [RN24]. Schrödinger [EN24, WZZ23]. Schur [HSWZ21, KM22]. Schur-constant [KM22]. Schur-like [HSWZ21]. score [BCHL24, BSB20, yCLjL23, OH21]. scores [MM22]. screening [CZLC20, HLS24, TLW22, WLL22]. SDE [LZ24b, Mar20a, Vid21a]. SDEs [CZ24, HW23a, LMSY22, MYX22, Mar20b, PSX21, SXX20]. Searching [HD22]. Second [LBB20, DS22a, Dou22, DP23, HZZZ21, JLL20, Mui20, Shp24]. second-degree [Shp24]. Second-order [LBB20, DP23, HZZZ21]. seemingly [JQS20]. segmentation [RZ22]. selecting [WZ24a]. selection [Bar23, DF24, DLZZ22, Ken21, LSFZ24, PD20, PC24, WZ24c, ZGZF21]. Self [SCMZ24, BHS21, SXY22, XNZL24]. self-excitation [BHS21]. self-intersection [XNZL24]. self-repelling [SXY22]. Self-weighted [SCMZ24]. selfsimilar [TT23]. Semi [KM20, LWZ25, BS22b, LZ23a, YL22]. Semi-classical [KM20]. semi-Markov [BS22b, YL22]. semi-parametric [LZ23a]. Semi-supervised [LWZ25]. semigroup [SXX20]. semilinear [Lv22]. semimartingales [EK22, HJO20, Mer23]. Semiparametric [YDB23, CY25, DS25, NAV21]. sense [WDP^{+}20]. sensitive [AZvdH23, BS22b]. Sensitivity [Cru20, Mao22]. separability [VM21]. separable [VP20]. September [Ano20y, Ano21x, Ano22x, Ano23x, Ano24x]. sequence [MU23, Pen23]. sequences [BL20, BDPR21, DM21, Fal22, Gan21b, Gir24, Haf23, PSZ24, Sze20]. Sequential [BSB22, CJ23, MO22]. serial [PS21]. Series [LLM20, DT21a, DK22, Hah22, HPH24, Ken21, MGMSP24, PP23, SY24, Wei21, ZPC21, ZH24]. server [GW20, SS23b]. set [BSB22, Gho23b, NS21, RRPB21, YB20a]. set-up [BSB22, Gho23b]. set-valued [NS21]. SETAR [MO22]. sets [DLS20, Daw21, nHL21, LL24c, SMRH24, SL25, TT23]. setup [Kev24]. several [AEMZ24, MEA21]. Shannon [ZS24]. shape [DML24, vNBR21]. shaped [MR22]. Sharp [Ahl22, Fan24, HS22, KO21a, Osȩ24, ZSK20]. sharpening [BSB20]. shifted [Bud22b]. shock [CE24]. shocks [HFC22, ZX25]. Short [GI24, Döb22, Yua21]. shortest [MT22]. shot [Kaw21]. show [CLP23]. Shrinkage [LZ23a, OK21, Las23, Lee22, WZ24c]. shrinking [Mui20]. side [Bab20]. sided [CÇ20, KK20, Lin24, WM24]. Sierpiński [KSH24]. sieve [SJSL21]. Sign [GLLOS24, VIVL23]. signal [MS23b]. signed [BP21]. Simes [XH22]. similarity [KC22, WZ24c]. Simple [JB21, Ahl22, AH23, FYY24, Yak21]. Simplex [HA24, Fri22]. simulating [Gra21, MBR22]. simulation [Lee23, MMB20, WZGY21]. simulations [Ber23b]. Simultaneous [Per24, Sub22, QP23, ZYLB22]. simultaneously [Che20a]. sine [Ma21]. single [Dal24, GW20, LWZ25, OK21, YZ22]. single-birth [Dal24]. single-index [LWZ25, YZ22]. single-spiked [OK21]. singular [CHW20, GM21, Lee22]. singularly [Yan25a]. SIR [LX24a]. SIRI [WFZ25]. SIS [MDS25]. sites [Hua24a]. size [BPB24, Cha20, Egi22, FV25, FLW22, JLL20, DML24]. size-and-shape [DML24]. size-biased [BPB24]. sizes [GLL23a, Moj21, Wri20]. Skew [GS22a, Val23, FFGS22, FS21b, FS24, HLS24, KZB23, LM21, Lou23, Tou21, XP20]. skew-Gaussian [KZB23]. skew-normal [FS24, XP20]. Skew-product [Val23]. skew-symmetric [HLS24]. skewness [ALO22, NOH21]. Skorokhod [NP20]. slab [SSL22]. sliced [PD21]. SLLN [XZ23]. Sloan [MRG22]. slope [DMS21, Mar21a]. slow [Zha22c]. small [BPB24, Gra22, Moj21, NP24, SJZ23, XC22b, YS24]. small-time [Gra22]. Smooth [Che24b, BP21, CW20]. smoothed [JSV20]. smoothness [XNZL24, XY23]. Snell [Mar20b]. Sobol [KR24, MRG22]. Sobolev [FS20a]. social [MHHJC22]. soft [WDG23]. sojourn [IO21]. solution [KIMG24, Lef24, LZZ24b, NNO20, NP20, SRK22, Vid21a]. Solutions [ZJ21, DP22, HT23b, HA25, KKO22, KKH24, LXZ20, Lin24, LW20, Mer23, MRZ20, PSX21]. Solvability [HWX22, WS20]. Some [BGV20, Din20, Egh21, FA24, FYY24, GR21, Gir24, HD23a, HPH24, Jas20, Liu20a, MX22, NS23a, PZL24, Pek22, RG20, Sai20, Sin23, Wei23, BLK20, BK22, DNTV22, DR21, Haf23, HX21a, Jou22, KSW20, KBM21, KB21, LKA20, LL21b, LMN24, MS20a, NP24, SN20, ZK22, ZCL22]. space [CWW20, CX24, nHL21, HL21b, LV23, LL20b, LYL22, Mah23, MN20, Myr21, RL20, SJK22, WZ24a]. space-filling [LYL22, WZ24a]. spaces [Cha24, CM23, FS20a, HSWZ21, Kar24, MC21, SWZ21, Suz18, Suz20, YY20, YM24]. spacing [NS24, SD22b]. spacings [Ahm20, BNS23, BS24, JGA24, Sin23, ZYY22]. Sparse [LW24a, CS24, DR21, GLLOS24, HNS23, HJKL21, LWM22, MT22, SSL22]. sparsity [BPB24, RB24]. Spatial [ART20, FV25, KSMF20, KZB23, LZ24a, LZ23a, SN24, VIVL23]. spatially [ZYL23]. SPDE [HLL20b]. SPDEs [LL20b, LZ24a, LW24d]. special [FS21b]. species [BL20]. specific [Hua24b]. specification [SY25]. specified [KZB23, MD24]. spectral [GL21a, LMZ21, Mau24, WH24]. spectrally [LWP21, MGMSP24, Ren24]. Speed [BMT23, Che24a]. Sphere [XLNQ20, Hir21, LMZ21]. spheres [AGP20, Lu24]. spherical [DM24, dMGP24, MMB20]. spherically [Nał21]. sphericity [Din20, SH22, ZZF22]. spike [SSL22]. spike-and-slab [SSL22]. spiked [LBL^{+}22, OK21]. split [MPD20, Mar23, SD22a]. split-plot [MPD20]. splitting [DDR20, WLZ24]. spot [HLL20a]. spread [BDK21]. square [ATT21, BH23, GM21, KVW24, Lee22, NW24, Pyc21, Sai25]. square-root [KVW24, NW24]. squared [KSW25, MP20b, MP20c, Nas20, PC24]. squares [Hwa21, NS22, SXY22, SJK22, ZK24]. SSEP [Zha22c]. Stability [Lv22, MLZ23, AFSW20, AM21, BTSL21, DT21a, FLLW24, GGX21, LXZ20, LL24d, NS21, RT22, TL22]. stabilization [LL24d]. stable [BU22, Gra21, GK22, HK21, KM20, Lar22, DMPP23, SXY22, SXX20, Tor24, WW23, ZYS20, ZX20, ZH24]. Stackelberg [ZZ24b]. stage [Hwa20, ZLS24]. standard [Bai20, Din21, IM23]. standby [LL23a]. star [AEAC21, MR22]. star-shaped [MR22]. stars [JLL20]. state [Col23, Le22, MO22]. states [LH25]. stationarity [Tra21b]. Stationary [Gat24, LNYZ22, WW23, WFZ25, CY25, Cru20, Gir24, Haf23, KKI21, KW21, Mar22, ZLWZ20, ZYS20, ZM23]. Statist [BBM22, BRBB23, Sko20, Suz20, ZCL22]. statistic [AA21, HNP20, OGR24, Pac20, SZ24, SN24, TP22, ZJD20]. Statistical [VP20, VM21, Zha25, AA21, BTS23, MP25, SBK20]. Statistics [ZLMT24, AKW22, AF22, AEAC21, Aly20, Aug20, BdCP23, BKM20b, CLP22, DP23, FS21a, Gir24, Jas20, Jou22, Kap22, LBB20, LT22c, MS20b, Pyc21, RB24, SM22, SBK20, Sko19, Sko20]. status [HC20]. stay [Sun25]. Stein [BU24, BVG20, BFMS22, Dal24, FR20, Ful24, GW20, GLS25, KKG20]. step [May23, RTT21, ZK24]. stepped [BFK21]. steps [DIY23]. stereographic [dMGP24]. stick [Ver22]. sticky [CÇ20, Tou21]. Stieltjes [FA24]. Stirling [Dou22]. Stochastic [ASKI24, HD23b, Kum22, Li23, LDM22, PCPS24, WXY22, AM21, Ahm22, AEMZ24, BBM22, BTSL21, BKZ21, BT25, CB20, DL21, DB22, FLLW24, Fri21, FF25, GN23, Gok25, GHH24, GGX21, HW24, HSS22, HJO20, Hil20, HW20, HT23b, HA25, IKCSKH25, JL25, JLZ24, KSAA24, KKC^{+}24, KKS24, KIMG24, LXZ20, LBB20, Lar22, LZ21, LL21b, LX24a, LCCZ24, LHW24, LZZ24b, Liu20a, LZ23b, LL24d, LG24, LW20, Lv22, MLZ23, MS20a, MST20, Mer23, MRZ20, MSM20, NS22, NNO20, PG20, Pir24, QW21, Rad23, RT22, RTT21, RT21, SRK22, Sto20, SQT24, TL22, WW23, WZZ23, WFZ25, WS20, WL21b, WH24, XX20b, XFJN22, YY22, Yan25a, Yan25c, YLP23, YSZ20, YS24, YZ20, YWY23, ZY20, Zha21a, ZYY22, ZYL23, ZHD24, ZX20, ZWW22, ZDF24]. stochastically [Dal24]. stopping [Gap20, KSAA24, Lin24, Mar23]. storable [KSAA24]. storage [NMR25]. Strassen [Kop24]. strata [OH21]. stratification [WZ24a]. stream [RT22]. strength [JWW21, PZL24, Sep20]. Strict [Yan25c, Tra21b]. Strictly [Ari23, Nie22]. string [Gzy21a]. Strong [HW23a, Izm20, JL25, JWW21, Lar22, SM24, Zha21a, Bou20, CD22, CD20a, Che20a, CS21, DM21, FZ23, KK25, LMSY22, MPK22, RT21, Sze20, Wal23, WYWS21]. Strongly [Ken21, MP20a]. structural [Mar21a, WZ24c]. structure [CHHS25, LBZ20, LH22, YL22]. structure-based [CHHS25]. structured [Che24d]. structures [FEVC21]. Student [GH23, Tre20]. Student- [GH23, Tre20]. Study [SS24, GTH24, LXH23]. Sub [Sar21b, Fen23, LT22a, LT24, Pin22, XK24, YL22, Zaj20, Zha21b, Zha22a]. Sub-exponential [Sar21b, Pin22, YL22]. sub-gaussian [Zaj20, LT22a]. sub-Gaussianity [LT24]. sub-linear [Fen23, XK24, Zha21b, Zha22a]. Subbotin [MS23b]. subclassification [OH21]. subcritical [KW21]. subdifferential [AZ23]. subexponential [Che20b, LCF21]. subgaussian [Tan24]. subgraphs [BV23]. sublinear [GL21b, GLL23b]. sublinearity [JSWZ22]. subordinated [DMS20b]. subordinators [GK22, Yak21]. subspaces [LZ25]. subsphericity [Vir21]. Success [Rib20]. successes [DS21]. SUE [SDH20]. sufficient [Che21, CT21, DL24, HJKL21, NV24]. sum [CL22, KK24, MKF20, RB20, Tan24]. sums [BU24, Che20b, CC24, Fen23, GK20, KZ21, LPŠ21, LLM24, Lev21, LBZ20, LMS22, ML24, NOH21, Pin22, PR21, QHK25, SYM25, SM24, YKK20, YL22, Zha22a, ZZ24d]. super [LMS22, LZ22a, Nak22]. super-Brownian [LZ22a]. super-heavy [LMS22, Nak22]. supercritical [DFG23b, FJ24, HLX22, SP22, WWL23, Ye24, ZH22, ZZ24d]. superiority [BPF20]. supermartingales [Ber23a]. supersaturated [SDH20]. supervised [LWZ25]. supplement [SYM25]. support [Efr24, Ida20, LPY24, Vil24]. Supports [QW21]. sure [DM21, GM22, Haf23, LXZ20, LZ21, LL24d, PR21]. surprising [FS20b]. Survival [HC20, LZ22a, AEMZ24, Sub22, YDB23, YZ23b]. Surviving [XZ22]. SV [SW20]. switching [FMR22, JX22, SS23b, Tra21a, TL22]. symmetric [Ahm20, AJO20, Arn21, HR22, HLS24, HW23b, KD24, KM20, Nał21, Nis24, Wan22c]. symmetrized [Zha23]. symmetry [ATT21, SD20, San24]. synchronization [KSMF20]. synchronized [LL24b]. synthetic [LD21b]. system [Ahm22, AJO20, IMT24, LL24b, LN20, NMR25, SP24a]. systems [ASKI24, GGX21, LL23a, LL24d, Łuc21, MLZ23, WC21].

tables [ATT21, BAB24, TG23, WK21]. Taboo [XFZ^{+}21]. Tail [FS21b, HPN22, Che25, DAA21, FJ24, JX22, JS20, DMPP23, Sep20, Zaj20]. tail-less [DAA21]. tailed [CC24, Häg20, KK20, LPŠ21, LMS22, Nak22, SYM25]. Tails [LPŠ21, AF22, DM24, XCDS22, ZZ25]. Talagrand [LL20b, LW24d]. tangential [Lu24]. Target [WZ24c]. task [WLZ24]. tau [FS21a, Per24, PvdHZ23]. tax [Sha25]. technique [Kro22]. telegraph [Cin22, IM23, Rat21]. Tempered [Mah23, BU22, BSST20, Gra21, GK22, Tor24, WSM24]. tensor [CHHS25]. term [KSW22, KIMG24]. terminal [KKH24]. terms [NWY20]. Test [San24, ZJD20, AAD25, BP21, BKM20a, BKM20b, DL24, FLW23, dMGP24, HNP20, Jou22, KC22, Lan21, LZ25, LBZH20, MPK22, Nov22, OGR24, RB24, SD20, SJSL21, SH22, Tra21b, WM24, WZGY21, WH24, XH22, ZZF22, ZZ24a]. Test-statistic [ZJD20]. Testing [BDK24, BPF20, DMN24, JGA24, Lan23, Ost22a, Ost22b, Vir21, BCDR23, CD20a, Che20a, DPSR20, EA21, FD24, Hah20, LT22a, MEA21, PZ20, QTLT20, Wan22b, XC22a, ZCP21]. Tests [LZMZ20, Ais23, BCDR23, BK23, Bib20, DNTV22, DS25, DDR20, Din20, HY23, LL21a, MLC25, PSP^{+}25, RG20, Sin23, SY25, ZF24]. th [BTSL21, BS24, LXZ20, PG21, SD22b, Sun20, XK24]. their [AIJD21, BTSL21, BU24, FS20a, Kay24, LLM24, NS23a, NWY20, NN21, Rat21, ZYY22]. theorem [BCM21, Bas24, BN24, CHW20, Chi22, CH23b, DFG23a, DT23, Din21, Döb22, DS20b, FM20, FWY20, FS24, Gao21, Gir24, GTH24, Jan24, Jur21, KO21b, Kop24, LZ21, LZLX25, MS23a, Mar21b, Mus22, Oud21, Ses24, Sze22, TZT21, Tap21, Tem23, YY22, Zha21b, ZX20]. theorem-type [Bas24]. theorems [BC24, BB22, Fen23, GR21, IK21, KMS21, KRN23, Myr21, PSZ24, QHK25, SD22b, SP22, Wan22a, WL21b, YJT20, SPT23]. Theoretical [GH23, WZGY21]. theory [CD21, GLL23b, Jan24, PZ20]. third [HA24]. three [CLLP23, KK25, YZ23a]. three-dimensional [KK25]. three-level [CLLP23, YZ23a]. threshold [LL23b]. thresholding [DF24, LW24a]. tight [Naa21]. tightening [Wri25]. tightness [BTS23, Döb22]. time [BG23, BDK21, CE24, CWW20, CY21, CX24, CFH23, Col23, DT21a, DIY23, DK22, DLZZ22, EA21, Gra22, GPP20, GPP22, Hah22, He24, HW23b, IM20, IO21, IM23, KR23, Kel24, Ken21, KBM21, KSW22, KMS21, KMP25, Lee23, LL20b, LWP21, LL21b, LW24b, LCCZ24, LCF21, LZ22b, Luo20, MU23, MC21, MPK22, MST20, NNO20, SS24, Ste23, Syl22, Vid23, VK24, Wan22c, Wan22a, Wei21, XFZ^{+}21, XNZL24, XZYC20, YJWY20, YDB23, YS24, ZSK20, ZPC21, ZX21, Zha22c, ZLZK20, ZDF24, ZL25]. time-changed [Vid23, Wan22c]. time-delayed [Luo20]. time-dependent [CY21, LCF21]. time-fractional [NNO20]. time-varying [He24, ZLZK20]. time/memory [Hah22]. times [CÇ20, CX24, Che24d, Col23, DIY23, FLW22, GLL23a, HX23, HX21a, KdS23, LWP21, LRW24, LMN24, Mar23, PCPS24, Rat21, YKK20]. Toeplitz [MS20b, Mau24]. Total [Lau24, AD20, BP24, Buo22, CW20, Osȩ24]. totally [Bez21, EA20]. tour [Hug24]. tournaments [MM22]. trace [WZLL22]. traces [Din20]. trade [DSH24]. trade-off [DSH24]. traffic [GW20]. trajectories [Mar21b]. transaction [Ren24]. transfer [CLP23]. transform [Ber23b, GK22, Hug21, Lan24, MST20, WWL20]. transformation [BP21, HDC24, Klo21, NW24, VY20]. transformations [CS20, LL20a, MT21b]. transforms [AJO20]. transient [AFSW20, CD20b]. transition [JZ23, Lou23, LMN24, Syl22, XZ20]. transitions [KP24]. transitivity [ZC23]. translation [Wai24]. transport [CPD20, GW21, GP23, NS21, Ses24]. Transportation [DL21, LL20b, LZ24a, LW24d, BBM22]. traps [BMT23]. treatment [BFK21, CX23, yCLjL23, CT21, Kab24, Liu20b, YDB23, ZGZF21]. tree [AVA23, AH23, CH23b, XZ22]. trees [JS20, MZ21]. trend [Mar20a]. trial [BFK21]. trials [DS21]. triangles [LD20]. Trimmed [SCGW20]. trinomial [RN24]. triviality [Bez21]. truncated [GWWY22, LXZ20, LRS20, MBR22]. truncation [LPY24]. Tsallis [BBL22]. TTT [Lan24]. tumble [CQ25]. Tweedie [SBS22]. Tweedie-type [SBS22]. twice [CS24, FE21, Sub21, Sub22]. Two [May23, ZK24, BSB22, BKB24, BPF20, CÇ20, Che24a, Che25, Ful24, Hua24a, Hwa20, JWW21, KP22b, Kry22, LGHG23, Lia20, Lou23, MD24, Oga23, RTT21, SQT24, Vog21, YDB23, ZLS24]. two-dimensional [Hua24a, Kry22]. two-factor [MD24]. two-level [Lia20]. two-plus [JWW21]. two-sided [CÇ20]. two-speed [Che24a]. two-stage [Hwa20, ZLS24]. Two-step [May23, ZK24, RTT21]. two-valued [Lou23]. two-way [LGHG23]. Type [XZ23, AM21, Aly20, BCDR23, BU22, Bas24, BM20, BRBB14, BRBB23, BFMS22, BS22c, Che20b, CQ25, CL23, Fen23, FS24, Hah22, HNP20, KSW20, LNYZ22, LP20, MA22, MX22, MK24, NS22, Nał21, Pak20, Pyc21, SZ24, SBS22, SY24, Sun25, YMWW25, YL22, ZAP24, Zha20b, Zou23]. typical [JQWD23].

Uhlenbeck [Bar21, CL23, HX21b, Lef24, Li23, SJZ23, YB20b, ZYS20, ZZ24b, Zha25, ZL20]. Ulam [AM21]. Ultimate [LZ23b]. Ultra [BdCP23, CZLC20, ZZZD21]. ultra-high [CZLC20]. ultra-large [ZZZD21]. Ultrahigh [Che21, LM24b]. Ultrahigh-dimensional [Che21, LM24b]. unbiased [BK23, BPB24, Bur21, JQS20]. unbiasedness [WM24]. unbounded [KKH24, YSZ20, ZL25]. uncertainty [BCRV23, LL24a, PK24]. unequal [Sto20]. unidirectional [ZW23]. Unified [SY25]. Uniform [Kri22, LSS22, LCCZ24, Xie23, ATT21, AEMZ24, CD21, Daw21, DM20, HMRT21, HC23, Kap22, Lar22, RT21, SM24, Yua21]. uniformity [dMGP24, JGA24]. Uniformly [BK23, Fri22]. Unique [Łuc21]. Uniqueness [CFH23, Yu22, KKO22, KIMG24, PSX21, SRK22, ZLMT23, ZLMT24]. unit [ART20, DZZ21, LMZ21, MY24]. unit-rank [DZZ21]. univariate [Yu22]. universal [FJ24, Kro22]. Universally [PSP^{+}25]. unknown [Efr24, LH25]. unordered [DH22]. unrelated [JQS20]. unreliable [Bic20]. until [MU23]. Up- [ASV24]. updates [Hah22]. updating [Kaw23]. upper [CW20, Cha24, ERR23, FJ24, Fro21, HS22, JB25, MKF20, VBS24]. uppermost [NS24, SD22b]. Urbanik [Jur23]. urn [Das24, GM22, XZYC20]. urns [FM21]. use [PF23]. used [AA21]. useful [NS23a]. using [BVG20, CD21, Döb22, JGA24, KN22, Lig23, LZ25, MST20, NW24, SZ24, WLL22, ZGZF21, ZYLB22]. usual [Fri21]. utilities [BS22b]. utility [FD24, Sar21a].

valid [TG23]. validation [WW21b]. value [Bar23, CHW20, GM21, Lee22, Pek22, XH22, YZ23b]. valued [BMP20, CLP23, KXBS21, KjL24, KS22, Lou23, NS21, TZT21]. values [ADM24, CK23, KR23, KKH24, Lee20, MC21, Myr21]. Vapnik [CD21]. VAR [BDBB20, SY24]. VaR-type [SY24]. variable [Bud22b, CT21, Jan21, TSG20, TL23]. variables [ADM24, AEAC21, BMP20, BH20, BB22, Bou20, Che20b, CC24, CHHS25, Che25, CL22, DMN24, DZ21, Egh21, GK20, GLS25, GR21, GN23, GM24, GL21b, HS22, Jas20, JB21, Jon22, KK22b, KK24, KSW25, LMS22, Mar21a, ML24, MK24, Myr21, Ols24, Pak20, Pin22, QHK25, SYM25, SJK22, Tan24, VBS24, WWL20, XK24, YLP23, Zaj20, ZZ25, Zha20b]. Variance [Fak20, ASV24, BPF20, Bud22b, Bur21, CBK22, FHSP21, Gau23, Gir20, Hwa21, JP20, KKI21, NLS24, OY23, Per24, Sai20, Sai25, FS21b]. variance-gamma [Gau23]. variant [YMWW25]. variate [LGHG23]. variates [Lev21]. variation [BP24, CW20, CCH20, Gra21, HSZ21, Kub20, Lau24, LHW24, Osȩ24, WSM24, XFJN22]. Variational [CM21b, LCH20, BT25]. various [NS23a]. varying [He24, KSW22, LWZ25, Mit21, TLW22, ZLZK20]. varying-coefficient [LWZ25]. vector [Fri22, JW24, KXBS21, KS22, LLM20, Lu24, TZT21]. vector-valued [KXBS21, KS22, TZT21]. vectors [DS20b, FLW22, JZ23, Rz21, Sun20]. version [MSM20]. versus [CX23, DW20, XCDS22]. vertices [FK21]. very [AF22]. via [BKM20b, BHS24, BSB20, CML20, CLP22, CFH23, CLZZ21, DDR20, DZZ21, Dou22, FS20a, Gir20, GL21a, JK22, Kaw21, KKG20, LW24a, MGD20, PSP^{+}25, Pir24, RK23, WZLL22, WK21, WLZ24, ZZ24a]. vibrations [MO23a]. Vine [BDBB20]. viscosity [SQT24]. visited [Hua24a]. Vlasov [JLZ24, LMSY22]. Volatility [YB20b, Bib20, GHH24, HLL20a, MST20]. Volterra [ASS24, JLZ24, SV20, WS20]. volume [BLK20, SX24].

Waiting [KMP25, FLW22, GLL23a, MU23]. Wald [SZ24]. Wald-type-statistic [SZ24]. walk [BG23, BH23, CH23a, FP22, FYY24, GM22, GS22b, IK21]. Walker [WFGR24]. walks [BMT23, CCY20, CD20b, GS21, KSH24, NWY20, Nak22, Sun25, Tzi20, Vog21]. Wasilkowski [MRG22]. Wasserstein [CW20, Cha24, FM20, LR22, Rie22, Sar20]. Watson [DFG23b, FJ24, KW21, LRW24, LL23b, TL23, XOC21]. Waugh [Bas24, Din21]. wave [Rad23, Yan25a]. Wavelet [Lee20]. waves [Gzy21a, Gzy21b, Vid21b]. way [LGHG23, LZ25, ZCP21]. Weak [AJO20, EN24, CM23, CD20b, KKO22, LL20a, Mao22, Mer23, MRZ20, SD22b, Tan24, Wal23, YY20]. Weak- [AJO20]. weakened [LTZ22]. wealth [Sar21a]. wedge [BFK21]. Weibull [KKT24, RRL20, WWL20, vNBR21]. Weighted [Hwa21, YM24, Che20b, GS20, LLL^{+}25, LL24c, Rz21, SYM25, SCMZ24, SM24, WSM24, Zha20a, HA24]. weighting [LP20]. weighting-type [LP20]. weights [GI24, LD20, Sol22]. Weiss [Lu22]. Weizsäcker [Tap21]. Well [CZ23, Gok25, HT23b, CZG24, KB21, ZWW22]. well-known [KB21]. Well-posedness [CZ23, Gok25, HT23b, CZG24, ZWW22]. Which [Jur23, TL23]. white [Cha21, HL21b, LL20b]. widely [XK24]. Wiener [BC24, DFG23a, KMS21]. Wilcoxon [BP21]. Wilks [TP22]. win [Oak25]. windows [Mui20]. winner [DR24]. Wishart [NP22, Oga23]. within [Ais23]. without [AKW22, Döb22, Vil24, ZLMT23, ZLMT24]. Wolfowitz [Naa21]. wordlength [WZQ25]. words [MU23]. workload [NMR25]. Woźniakowski [MRG22]. wrapping [Gat24]. WVaR [YJT20].

Young [FY24]. Yu [Kab24]. Yule [WFGR24].

Zastavnyi [FPBK20]. zero [CLP23, DB23, IM20, LP20, PvdHZ23, Vid21a]. zero-inflated [LP20, PvdHZ23]. zeros [GS21]. Zipf [LL24c]. Zolotarev [Ye24]. Zygmund [Bou20, DDG22, Sze20].

References

Ahmad:2021:DSU

[AA21] M. Rauf Ahmad and S. Ejaz Ahmed. On the distribution of the T^{2} statistic, used in statistical process monitoring, for highdimensional data. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

           (electronic). URL http://www.sciencedirect.com/science/
           article/pii/S0167715220302224.

Azimi-Abarghouyi:2024:MCP

[AAD24] Seyed Mohammad Azimi-Abarghouyi and Harpreet S. Dhillon. Matérn cluster process with holes at the cluster centers. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

           URL http://www.sciencedirect.com/science/article/pii/
           S0167715223001554.

Ahmed:2025:IBR

[AAD25] S. Ejaz Ahmed, Mohamed Amezziane, and Keshab Dahal. Improving the Bickel–Rosenblatt global measure of deviation and goodness-of-fit test. Statistics & Probability Letters, 218(??): ??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

           1879-2103 (electronic). URL http://www.sciencedirect.com/
           science/article/pii/S0167715224002797.

Amel:2022:RKE

[AAE22] Azzi Amel, Laksaci Ali, and Ould Saïd Elias. On the robustification of the kernel estimator of the functional modal regression. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002182.

Angus:2020:RCA

[AD20] John Angus and Yujia Ding. On the ratio of current age to total life for null recurrent renewal processes. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300481.

Akbash:2024:APE

[ADM24] Kateryna Akbash, Natalia Doronina, and Ivan Matsak. On the asymptotic properties of extreme values of discrete random variables. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001937.

Alimohammadi:2021:DSO

[AEAC21] Mahdi Alimohammadi, Maryam Esna-Ashari, and Erhard Cramer. On dispersive and star orderings of random variables and order statistics. Statistics & Probability Letters, 170(??): ??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220303175.

Arnold:2024:ESS

[AEMZ24] Sebastian Arnold, Hammou El Barmi, Hari Mukerjee, and Johanna Ziegel. Estimating several survival functions under uniform stochastic ordering. Statistics & Probability Letters, 208 (??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000142.

Albrecher:2022:AAG

[AF22] Hansjörg Albrecher and Brandon García Flores. Asymptotic analysis of generalized Greenwood statistics for very heavy tails.

Statistics & Probability Letters, 185(??):??, June 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000396.

Adan:2020:LST

[AFSW20] Ivo Adan, Sergey Foss, Seva Shneer, and Gideon Weiss. Local stability in a transient Markov chain. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220301589.

Arafat:2020:SCC

[AGP20] Ahmed Arafat, Pablo Gregori, and Emilio Porcu. Schoenberg coefficients and curvature at the origin of continuous isotropic positive definite kernels on spheres. Statistics & Probability Letters, 156 (??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302640.

Asfha:2023:APR

[AH23] Huruy Asfha and Xiaomi Hu. An algorithm for a pseudo RMLE under simple tree multivariate order restriction. Statistics & Probability Letters, 202(??):??, November 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001293.

Ahle:2022:SSB

[Ahl22] Thomas D. Ahle. Sharp and simple bounds for the raw moments of the binomial and Poisson distributions. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002662.

Ahmadi:2020:CRS

[Ahm20] Jafar Ahmadi. Characterization results for symmetric continuous distributions based on the properties of k-records and spacings. Statistics & Probability Letters, 162(??):??, July 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300675.

Ahmed:2022:CCC

[Ahm22] Hamdy M. Ahmed. Construction controllability for conformable fractional stochastic evolution system with noninstantaneous impulse and nonlocal condition. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001523.

Anaya-Izquierdo:2021:FCH

[AIJD21] Karim Anaya-Izquierdo, M. C. Jones, and Alice Davis. A family of cumulative hazard functions and their frailty connections. Statistics & Probability Letters, 172(??):??, May 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000213.

Aisbett:2023:ITH

[Ais23] Janet Aisbett. Interpreting tests of a hypothesis at multiple alpha levels within a Neyman–Pearson framework. Statistics & Probability Letters, 201(??):??, October 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001232.

Akboudj:2020:WIN

[AJO20] Meryem Akboudj, Yong Jiao, and Adam Osȩkowski. Weak-L^{∞} inequality for non-symmetric martingale transforms and Haar system. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030081X.

Ai:2022:HBI

[AKW22] Jianhang Ai, Ondřej Kuželka, and Yuyi Wang. Hoeffding and Bernstein inequalities for U-statistics without replacement. Statistics & Probability Letters, 187(??):??, August 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000992.

Arab:2022:CQR

[ALO22] Idir Arab, Tommaso Lando, and Paulo Eduardo Oliveira. Comparison of L_{p}-quantiles and related skewness measures. Statistics & Probability Letters, 183(??):??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522100287X.

Aly:2020:OSL

[Aly20] Emad-Eldin A. A. Aly. On order statistics from Laplace-type distributions. Statistics & Probability Letters, 160(??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300092.

Alhakim:2019:DFP

[AM19] Abbas Alhakim and S. Molchanov. The density flatness phenomenon. Statistics & Probability Letters, 152(??):156–161, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic).  URL http://www.sciencedirect.
com/science/article/pii/S016771521930135X. See comments

[Wei23].

Ahmadova:2021:UHS

[AM21] Arzu Ahmadova and Nazim I. Mahmudov. Ulam–Hyers stability of Caputo type fractional stochastic neutral differential equations. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302522.

Anonymous:2020:Aa

[Ano20a] Anonymous. April 2020. Statistics & Probability Letters, 159(??): ??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. August 2020. Statistics & Probability Letters, 163 (??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:D

[Ano20c] Anonymous. December 2020. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:EBa

[Ano20d] Anonymous. Editorial Board. Statistics & Probability Letters, 156 (??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302871.

Anonymous:2020:EBb

[Ano20e] Anonymous. Editorial Board. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219302950.

Anonymous:2020:EBc

[Ano20f] Anonymous. Editorial Board. Statistics & Probability Letters, 158 (??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303347.

Anonymous:2020:EBd

[Ano20g] Anonymous. Editorial Board. Statistics & Probability Letters, 159 (??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300134.

Anonymous:2020:EBe

[Ano20h] Anonymous. Editorial Board. Statistics & Probability Letters, 160 (??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300419.

Anonymous:2020:EBf

[Ano20i] Anonymous. Editorial Board. Statistics & Probability Letters, 161 (??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300614.

Anonymous:2020:EBg

[Ano20j] Anonymous. Editorial Board. Statistics & Probability Letters, 162 (??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300894.

Anonymous:2020:EBh

[Ano20k] Anonymous. Editorial Board. Statistics & Probability Letters, 163 (??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301255.

Anonymous:2020:EBi

[Ano20l] Anonymous. Editorial Board. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220301437.

Anonymous:2020:EBj

[Ano20m] Anonymous. Editorial Board. Statistics & Probability Letters, 165 (??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301899.

Anonymous:2020:EBk

[Ano20n] Anonymous. Editorial Board. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302108.

Anonymous:2020:EBl

[Ano20o] Anonymous. Editorial Board. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771522030242X.

Anonymous:2020:F

[Ano20p] Anonymous. February 2020. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:Ja

[Ano20q] Anonymous. January 2020. Statistics & Probability Letters, 156 (??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:Jc

[Ano20r] Anonymous. July 2020. Statistics & Probability Letters, 162(??): ??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:Jb

[Ano20s] Anonymous. June 2020. Statistics & Probability Letters, 161(??): ??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:Ma

[Ano20t] Anonymous. March 2020. Statistics & Probability Letters, 158 (??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:Mb

[Ano20u] Anonymous. May 2020. Statistics & Probability Letters, 160(??): ??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:N

[Ano20v] Anonymous. November 2020. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:O

[Ano20w] Anonymous. October 2020. Statistics & Probability Letters, 165 (??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2020:PN

[Ano20x] Anonymous. Publisher note. Statistics & Probability Letters, 158 (??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303402.

Anonymous:2020:S

[Ano20y] Anonymous. September 2020. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. Statistics & Probability Letters, 171(??): ??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:Ab

[Ano21b] Anonymous. August 2021. Statistics & Probability Letters, 175 (??):??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. Statistics & Probability Letters, 179(??):??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:EBa

[Ano21d] Anonymous. Editorial Board. Statistics & Probability Letters, 168 (??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302820.

Anonymous:2021:EBb

[Ano21e] Anonymous. Editorial Board. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220303084.

Anonymous:2021:EBc

[Ano21f] Anonymous. Editorial Board. Statistics & Probability Letters, 170 (??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220303370.

Anonymous:2021:EBd

[Ano21g] Anonymous. Editorial Board. Statistics & Probability Letters, 171 (??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000146.

Anonymous:2021:EBe

[Ano21h] Anonymous. Editorial Board. Statistics & Probability Letters, 172 (??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000390.

Anonymous:2021:EBf

[Ano21i] Anonymous. Editorial Board. Statistics & Probability Letters, 173 (??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000699.

Anonymous:2021:EBg

[Ano21j] Anonymous. Editorial Board. Statistics & Probability Letters, 174 (??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000869.

Anonymous:2021:EBh

[Ano21k] Anonymous. Editorial Board. Statistics & Probability Letters, 175 (??):??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522100105X.

Anonymous:2021:EBi

[Ano21l] Anonymous. Editorial Board. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221001334.

Anonymous:2021:EBj

[Ano21m] Anonymous. Editorial Board. Statistics & Probability Letters, 177 (??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001632.

Anonymous:2021:EBk

[Ano21n] Anonymous. Editorial Board. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221001814.

Anonymous:2021:EBl

[Ano21o] Anonymous. Editorial Board. Statistics & Probability Letters, 179(??):??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002121.

Anonymous:2021:F

[Ano21p] Anonymous. February 2021. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:Ja

[Ano21q] Anonymous. January 2021. Statistics & Probability Letters, 168 (??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:Jc

[Ano21r] Anonymous. July 2021. Statistics & Probability Letters, 174(??): ??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:Jb

[Ano21s] Anonymous. June 2021. Statistics & Probability Letters, 173(??): ??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:Ma

[Ano21t] Anonymous. March 2021. Statistics & Probability Letters, 170 (??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:Mb

[Ano21u] Anonymous. May 2021. Statistics & Probability Letters, 172(??): ??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:N

[Ano21v] Anonymous. November 2021. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:O

[Ano21w] Anonymous. October 2021. Statistics & Probability Letters, 177 (??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2021:S

[Ano21x] Anonymous. September 2021. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Statistics & Probability Letters, 183(??): ??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Statistics & Probability Letters, 187 (??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Statistics & Probability Letters, 180 (??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002455.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002704.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Statistics & Probability Letters, 182 (??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002984.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Statistics & Probability Letters, 183 (??):??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000219.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Statistics & Probability Letters, 184 (??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000487.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Statistics & Probability Letters, 185 (??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000700.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Statistics & Probability Letters, 186 (??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522200092X.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Statistics & Probability Letters, 187 (??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001134.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Statistics & Probability Letters, 188(??):??, September 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001316.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. Statistics & Probability Letters, 189 (??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001602.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001869.

Anonymous:2022:EBl

[Ano22o] Anonymous. Editorial Board. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222002000.

Anonymous:2022:F

[Ano22p] Anonymous. February 2022. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:Ja

[Ano22q] Anonymous. January 2022. Statistics & Probability Letters, 180 (??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:Jc

[Ano22r] Anonymous. July 2022. Statistics & Probability Letters, 186(??): ??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:Jb

[Ano22s] Anonymous. June 2022. Statistics & Probability Letters, 185(??): ??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:Ma

[Ano22t] Anonymous. March 2022. Statistics & Probability Letters, 182 (??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:Mb

[Ano22u] Anonymous. May 2022. Statistics & Probability Letters, 184(??): ??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:N

[Ano22v] Anonymous. November 2022. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:O

[Ano22w] Anonymous. October 2022. Statistics & Probability Letters, 189 (??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2022:S

[Ano22x] Anonymous. September 2022. Statistics & Probability Letters, 188(??):??, September 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Statistics & Probability Letters, 195(??): ??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Statistics & Probability Letters, 199 (??):??, August 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Statistics & Probability Letters, 203(??):??, December 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Statistics & Probability Letters, 192 (??):??, January 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002383.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222002656.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Statistics & Probability Letters, 194 (??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002863.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Statistics & Probability Letters, 195 (??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000196.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Statistics & Probability Letters, 196 (??):??, May 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000500.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Statistics & Probability Letters, 197 (??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000676.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Statistics & Probability Letters, 198 (??):??, July 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000871.

Anonymous:2023:EBh

[Ano23k] Anonymous. Editorial Board. Statistics & Probability Letters, 199 (??):??, August 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000998.

Anonymous:2023:EBi

[Ano23l] Anonymous. Editorial Board. Statistics & Probability Letters, 200(??):??, September 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223001177.

Anonymous:2023:EBj

[Ano23m] Anonymous. Editorial Board. Statistics & Probability Letters, 201 (??):??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001335.

Anonymous:2023:EBk

[Ano23n] Anonymous. Editorial Board. Statistics & Probability Letters, 202(??):??, November 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223001499.

Anonymous:2023:EBl

[Ano23o] Anonymous. Editorial Board. Statistics & Probability Letters, 203(??):??, December 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771522300161X.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023. Statistics & Probability Letters, 192 (??):??, January 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:Jc

[Ano23r] Anonymous. July 2023. Statistics & Probability Letters, 198(??): ??, July 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:Jb

[Ano23s] Anonymous. June 2023. Statistics & Probability Letters, 197(??): ??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:Ma

[Ano23t] Anonymous. March 2023. Statistics & Probability Letters, 194 (??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:Mb

[Ano23u] Anonymous. May 2023. Statistics & Probability Letters, 196(??): ??, May 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:N

[Ano23v] Anonymous. November 2023. Statistics & Probability Letters, 202(??):??, November 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023. Statistics & Probability Letters, 201 (??):??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2023:S

[Ano23x] Anonymous. September 2023. Statistics & Probability Letters, 200(??):??, September 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Statistics & Probability Letters, 207(??): ??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. Statistics & Probability Letters, 211 (??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:D

[Ano24c] Anonymous. December 2024. Statistics & Probability Letters, 215(??):??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:EBa

[Ano24d] Anonymous. Editorial Board. Statistics & Probability Letters, 204 (??):??, January 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001906.

Anonymous:2024:EBb

[Ano24e] Anonymous. Editorial Board. Statistics & Probability Letters, 205(??):??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223002092.

Anonymous:2024:EBc

[Ano24f] Anonymous. Editorial Board. Statistics & Probability Letters, 206 (??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522300250X.

Anonymous:2024:EBd

[Ano24g] Anonymous. Editorial Board. Statistics & Probability Letters, 207 (??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000348.

Anonymous:2024:EBe

[Ano24h] Anonymous. Editorial Board. Statistics & Probability Letters, 208 (??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000683.

Anonymous:2024:EBf

[Ano24i] Anonymous. Editorial Board. Statistics & Probability Letters, 209 (??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000944.

Anonymous:2024:EBg

[Ano24j] Anonymous. Editorial Board. Statistics & Probability Letters, 210 (??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522400110X.

Anonymous:2024:EBh

[Ano24k] Anonymous. Editorial Board. Statistics & Probability Letters, 211 (??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001305.

Anonymous:2024:EBi

[Ano24l] Anonymous. Editorial Board. Statistics & Probability Letters, 212(??):??, September 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001470.

Anonymous:2024:EBj

[Ano24m] Anonymous. Editorial Board. Statistics & Probability Letters, 213 (??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001780.

Anonymous:2024:EBk

[Ano24n] Anonymous. Editorial Board. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224002049.

Anonymous:2024:EBl

[Ano24o] Anonymous. Editorial Board. Statistics & Probability Letters, 215(??):??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224002323.

Anonymous:2024:F

[Ano24p] Anonymous. February 2024. Statistics & Probability Letters, 205(??):??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:Ja

[Ano24q] Anonymous. January 2024. Statistics & Probability Letters, 204 (??):??, January 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:Jc

[Ano24r] Anonymous. July 2024. Statistics & Probability Letters, 210(??): ??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:Jb

[Ano24s] Anonymous. June 2024. Statistics & Probability Letters, 209(??): ??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:Ma

[Ano24t] Anonymous. March 2024. Statistics & Probability Letters, 206 (??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:M

[Ano24u] Anonymous. May 2024. Statistics & Probability Letters, 208(??): ??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:N

[Ano24v] Anonymous. November 2024. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:O

[Ano24w] Anonymous. October 2024. Statistics & Probability Letters, 213 (??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2024:S

[Ano24x] Anonymous. September 2024. Statistics & Probability Letters, 212(??):??, September 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2025:EBa

[Ano25a] Anonymous. Editorial Board. Statistics & Probability Letters, 216 (??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002682.

Anonymous:2025:EBb

[Ano25b] Anonymous. Editorial Board. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224002852.

Anonymous:2025:EBc

[Ano25c] Anonymous. Editorial Board. Statistics & Probability Letters, 218 (??):??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224003031.

Anonymous:2025:F

[Ano25d] Anonymous. February 2025. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2025:Ja

[Ano25e] Anonymous. January 2025. Statistics & Probability Letters, 216 (??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Anonymous:2025:Ma

[Ano25f] Anonymous. March 2025. Statistics & Probability Letters, 218 (??):??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Arevalillo:2024:EAQ

[Are24] Jorge M. Arevalillo. On the empirical approximation to quantiles from Lugannani–Rice saddlepoint formula. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000749.

Aristodemou:2023:SLC

[Ari23] Eleni Aristodemou. Strictly log-concave probability distributions in discrete response models. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001961.

Arnal:2021:CSL

[Arn21] Charles Arnal. Convolution of a symmetric log-concave distribution and a symmetric bimodal distribution can have any number of modes. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000444.

Adu:2020:POS

[ART20] N. Adu, G. Richardson, and M. C. Tseng. Periodogram ordinate: Spatial model with near unit roots and dependent errors. Statistics & Probability Letters, 157(??):??, February 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302615.

Amini-Seresht:2024:SCP

[ASKI24] Ebrahim Amini-Seresht, Baha-Eldin Khaledi, and Salman Izadkhah. Stochastic comparison of parallel systems with heterogeneous dependent exponential components. Statistics & Probability Letters, 215(??):??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002116.

Almani:2024:PGV

[ASS24] Hamidreza Maleki Almani, Foad Shokrollahi, and Tommi Sottinen. Prediction of Gaussian Volterra processes with compound Poisson jumps. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000233.

Ansari:2024:CNV

[ASV24] Jonathan Ansari, Tomer Shushi, and Steven Vanduffel. Up- and down-correlations in normal variance mixture models. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001736.

Ando:2021:EDU

[ATT21] Shuji Ando, Kouji Tahata, and Sadao Tomizawa. Extended diagonal uniform association symmetry model for square contingency tables with ordinal categories. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302765.

Augustynowicz:2020:ABP

[Aug20] Aneta Augustynowicz. Asymptotic behavior of proportions of observations falling to random regions determined by central order statistics. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030047X.

Arenas-Velilla:2023:CRN

[AVA23] Santiago Arenas-Velilla and Octavio Arizmendi. Convergence rate for the number of crossings in a random labelled tree. Statistics & Probability Letters, 203(??):??, December 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001402.

Arya:2020:RAN

[AY20] Sakshi Arya and Yuhong Yang. Randomized allocation with nonparametric estimation for contextual multi-armed bandits with delayed rewards. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301218.

Ahmed:2023:ENC

[AZ23] Hamdy M. Ahmed and Quanxin Zhu. Exploration nonlocal controllability for Hilfer fractional differential inclusions with Clarke subdifferential and nonlinear noise. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000056.

Almalik:2023:CMS

[AZvdH23] Osama Almalik, Zhuozhao Zhan, and Edwin R. van den Heuvel. Copas’ method is sensitive to different mechanisms of publication bias. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002462.

Babkin:2020:ISI

[Bab20] Andrey Babkin. Incorporating side information into Robust Matrix Factorization with Bayesian Quantile Regression. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301504.

Boos:2024:EPC

[BAB24] Dennis D. Boos, Shannon Ari, and Roger L. Berger. Exact partially conditional binomial analysis for multinomial data in 2 × 2 tables. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001640.

Bai:2020:ESM

[Bai20] Long Bai. Extremes of standard multifractional Brownian motion. Statistics & Probability Letters, 159(??):??, April 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303438.

Barrera:2021:CPM

[Bar21] Gerardo Barrera. Cutoff phenomenon for the maximum of a sampling of Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302571.

Bardet:2023:LMB

[Bar23] Jean-Marc Bardet. Laplace’s method and BIC model selection for least absolute value criterion. Statistics & Probability Letters, 195 (??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002772.

Basu:2024:FWL

[Bas24] Deepankar Basu. Frisch–Waugh–Lovell theorem-type results for the k-class and 2SGMM estimators. Statistics & Probability Letters, 213(??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001573.

Battey:2021:NAA

[Bat21] H. S. Battey. A note on the analytic approximation of exceedance probabilities in heterogeneous populations. Statistics & Probability Letters, 179(??):??, December 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001772.

Bernou:2022:LTD

[BB22] Ismahen Bernou and Fakhreddine Boukhari. Limit theorems for dependent random variables with infinite means. Statistics & Probability Letters, 189(??):??, October 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001213.

Bareche:2023:IGD

[BB23] Aîcha Bareche and Abdelouahab Bibi. On inverse-gamma distribution delayed by Poisson process. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223000111.

Balakrishnan:2022:TEA

[BBL22] Narayanaswamy Balakrishnan, Francesco Buono, and Maria Longobardi. On Tsallis extropy with an application to pattern recognition. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002030.

Bahlali:2022:ETC

[BBM22] Khaled Bahlali, Brahim Boufoussi, and Soufiane Mouchtabih. Errata to “Transportation cost inequality for backward stochastic differential equations” [Statist. Probab. Lett. 155 (2019) 108586]. Statistics & Probability Letters, 182(??):??, March 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002649.

Bai:2024:ELT

[BC24] Shuyang Bai and Jiemiao Chen. Empirical limit theorems for Wiener chaos. Statistics & Probability Letters, 215(??):??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001913.

Banzato:2023:BTC

[BCDR23] Erika Banzato, Monica Chiogna, Vera Djordjilović, and Davide Risso. A Bartlett-type correction for likelihood ratio tests with application to testing equality of Gaussian graphical models. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002450.

Bodory:2024:NBP

[BCHL24] Hugo Bodory, Lorenzo Camponovo, Martin Huber, and Michael Lechner. Nonparametric bootstrap for propensity score matching estimators. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000385.

Bang:2021:GAT

[BCM21] David Bang, Jorge González Cázares, and Aleksandar Mijatović. A Gaussian approximation theorem for Lévy processes. Statistics & Probability Letters, 178(??):??, November 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001498.

Bernard:2023:CUD

[BCRV23] Carole Bernard, Jinghui Chen, Ludger Rüschendorf, and Steven Vanduffel. Coskewness under dependence uncertainty. Statistics & Probability Letters, 199(??):??, August 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000779.

Begin:2020:MVD

[BDBB20] Étienne Begin, Pierre Dutilleul, Carole Beaulieu, and Taoufik Bouezmarni. M-Vine decomposition and VAR(1) models. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303062.

Badiella:2023:ULC

[BdCP23] Llorenç Badiella, Joan del Castillo, and Pedro Puig. Ultra logconcavity of discrete order statistics. Statistics & Probability Letters, 201(??):??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223001244.

Bezborodov:2021:CTF

[BDK21] Viktor Bezborodov, Luca Di Persio, and Tyll Krueger. The continuous-time frog model can spread arbitrarily fast. Statistics & Probability Letters, 172(??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000080.

Bastian:2024:TEM

[BDK24] P. Bastian, H. Dette, and L. Koletzko. Testing equivalence of multinomial distributions — a constrained bootstrap approach. Statistics & Probability Letters, 206(??):??, March 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002237.

Berti:2021:ACC

[BDPR21] Patrizia Berti, Emanuela Dreassi, Luca Pratelli, and Pietro Rigo. Asymptotics of certain conditionally identically distributed sequences. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302261.

Belmabrouk:2022:DEP

[BDY22] Nadia Belmabrouk, Mondher Damak, and Nejib Yaakoubi. Dirichlet eigenvalue problems of irreversible Langevin diffusion. Statistics & Probability Letters, 180(??):??, January 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002042.

Berkaoui:2023:OOD

[Ber23a] Abdelkarem Berkaoui. On the optional and orthogonal decompositions of supermartingales and applications. Statistics & Probability Letters, 199(??):??, August 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000743.

Bernardoff:2023:DEL

[Ber23b] Philippe Bernardoff. Domain of existence of the Laplace transform of negative multinomial distributions and simulations. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200222X.

Bezborodov:2021:NTT

[Bez21] Viktor Bezborodov. Non-triviality in a totally asymmetric onedimensional Boolean percolation model on a half-line. Statistics & Probability Letters, 176(??):??, September 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001176.

Bowden:2021:CTE

[BFK21] Rhys Bowden, Andrew B. Forbes, and Jessica Kasza. On the centrosymmetry of treatment effect estimators for stepped wedge and related cluster randomized trial designs. Statistics & Probability Letters, 172(??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303254.

Boukehil:2022:HJS

[BFMS22] Djamila Boukehil, Dominique Fourdrinier, Fatiha Mezoued, and William E. Strawderman. On homogeneous James–Stein type estimators. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001675.

Bandyopadhyay:2023:RMP

[BG23] Antar Bandyopadhyay and Partha Pratim Ghosh. Right-most position of a last progeny modified time inhomogeneous branching random walk. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002103.

Bongiorno:2020:ECI

[BGV20] E. G. Bongiorno, A. Goia, and P. Vieu. Estimating the complexity index of functional data: Some asymptotics. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300341.

Baumdicker:2020:MCR

[BH20] F. Baumdicker and U. Hölker. Method comparison with repeated measurements — Passing–Bablok regression for grouped data with errors in both variables. Statistics & Probability Letters, 164(??): ??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301048.

Beaton:2023:MSD

[BH23] Nicholas R. Beaton and Mark Holmes. The mean square displacement of random walk on the Manhattan lattice. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200219X.

Bonnet:2021:MLE

[BHS21] Anna Bonnet, Miguel Martinez Herrera, and Maxime Sangnier. Maximum likelihood estimation for Hawkes processes with selfexcitation or inhibition. Statistics & Probability Letters, 179(??): ??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001760.

Borsato:2024:PDL

[BHS24] Luísa Borsato, Eduardo Horta, and Rafael Rigão Souza. Product disintegrations: a law of large numbers via conditional independence. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000257.

Bibinger:2020:CTC

[Bib20] Markus Bibinger. Cusum tests for changes in the Hurst exponent and volatility of fractional Brownian motion. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300286.

Bickel:2020:DBI

[Bic20] David R. Bickel. Departing from Bayesian inference toward minimaxity to the extent that the posterior distribution is unreliable. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030105X.

Balakrishnan:2022:CBD

[BJ22] N. Balakrishnan and M. C. Jones. Closure of beta and Dirichlet distributions under discrete mixing. Statistics & Probability Letters, 188(??):??, September 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000980.

Bhat:2022:GSC

[BK22] M. Ashraf Bhat and G. Sankara Raju Kosuru. Generalizations of some concentration inequalities. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002601.

Bedbur:2023:UMP

[BK23] S. Bedbur and U. Kamps. Uniformly most powerful unbiased tests for the dispersion parameter of the Conway–Maxwell–Poisson distribution. Statistics & Probability Letters, 196(??):??, May 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000251.

Bobotas:2024:PAP

[BK24] Panayiotis Bobotas and Markos V. Koutras. On the preservation of ageing properties under random maxima. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000804.

Bhakta:2024:ORB

[BKB24] Raju Bhakta, Suchandan Kayal, and Narayanaswamy Balakrishnan. Ordering results between two multiple-outlier finite δmixtures. Statistics & Probability Letters, 213(??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001627.

Balakrishna:2020:LFP

[BKK20] N. Balakrishna, Jiwoong Kim, and Hira L. Koul. Lack-of-fit of a parametric measurement error AR(1) model. Statistics & Probability Letters, 166(??):??, November 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301759.

Bhattacharyya:2020:NTC

[BKM20a] Dhrubasish Bhattacharyya, Ruhul Ali Khan, and Murari Mitra. A nonparametric test for comparison of mean past lives. Statistics & Probability Letters, 161(??):??, June 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300250.

Bhattacharyya:2020:TEA

[BKM20b] Dhrubasish Bhattacharyya, Ruhul Ali Khan, and Murari Mitra. A test of exponentiality against DMTTF alternatives via L-statistics. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301565.

Barmalzan:2021:SCF

[BKZ21] Ghobad Barmalzan, Sajad Kosari, and Yiying Zhang. On stochastic comparisons of finite α-mixture models. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000456.

Bassetti:2020:ANC

[BL20] Federico Bassetti and Lucia Ladelli. Asymptotic number of clusters for species sampling sequences with non-diffuse base measure. Statistics & Probability Letters, 162(??):??, July 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300523.

Bedbur:2020:MVP

[BLK20] S. Bedbur, J. M. Lennartz, and U. Kamps. On minimum volume properties of some confidence regions for multiple multivariate normal means. Statistics & Probability Letters, 158(??): ??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303220.

Bar-Lev:2025:CID

[BLL25] Shaul K. Bar-Lev and Gérard Letac. Chainability of infinitely divisible measures. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002256.

Bednorz:2020:LOT

[BM20] Witold Bednorz and Rafał Martynek. A Lévy–Ottaviani type inequality for the Bernoulli process on an interval. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522030050X.

Beghin:2022:NCM

[BM22] Luisa Beghin and Claudio Macci. Non-central moderate deviations for compound fractional Poisson processes. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000360.

Baena-Mirabete:2020:CPI

[BMP20] S. Baena-Mirabete and P. Puig. Computing probabilities of integer-valued random variables by recurrence relations. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300225.

Betz:2023:SFB

[BMT23] Volker Betz, Matthias Meiners, and Ivana Tomic. Speed function for biased random walks with traps. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222002784.

Bocinec:2024:CEA

[BN24] Filip Bočinec and Stanislav Nagy. Conditions for equality in Anderson’s theorem. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000634.

Balogoun:2021:ANG

[BNO21] Armando Sosthène Kali Balogoun, Guy Martial Nkiet, and Carlos Ogouyandjou. Asymptotic normality of a generalized maximum mean discrepancy estimator. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302649.

Balakrishnan:2023:NS

[BNS23] N. Balakrishnan, V. B. Nevzorov, and A. Stepanov. On normal spacings. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002267.

Boukhari:2020:MZS

[Bou20] Fakhreddine Boukhari. The Marcinkiewics–Zygmund strong law of large numbers for dependent random variables. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300304.

Barabesi:2020:GBL

[BP20] Lucio Barabesi and Luca Pratelli. On the Generalized Benford law. Statistics & Probability Letters, 160(??):??, May 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300055.

Bagkavos:2021:IWS

[BP21] Dimitrios Bagkavos and Prakash N. Patil. Improving the Wilcoxon signed rank test by a kernel smooth probability integral transformation. Statistics & Probability Letters, 171(??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220303291.

Barabesi:2024:ITV

[BP24] Lucio Barabesi and Luca Pratelli. An inequality for the total variation distance between high-dimensional centered Gaussian laws. Statistics & Probability Letters, 211(??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001172.

Bondi:2024:SIP

[BPB24] Laura Bondi, Marcello Pagano, and Marco Bonetti. The sparsity index in Poisson size-biased sampling: Algorithms for the optimal unbiased estimation from small samples. Statistics & Probability Letters, 214(??):??, November 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400186X.

Boente:2020:TSB

[BPF20] Graciela Boente and Juan Carlos Pardo-Fernández. Testing for superiority between two variance functions. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302585.

Barabesi:2023:CJC

[BPR23] Lucio Barabesi, Luca Pratelli, and Pietro Rigo. On the Chvátal– Janson conjecture. Statistics & Probability Letters, 194(??): ??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002577.

Bravo:2023:LPE

[Bra23] Francesco Bravo. Local polynomial estimation of nonparametric general estimating equations. Statistics & Probability Letters, 197 (??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000299.

Borges:2014:CPT

[BRBB14] Patrick Borges, Josemar Rodrigues, Narayanaswamy Balakrishnan, and Jorge Bazán. A COM–Poisson type generalization of the binomial distribution and its properties and applications. Statistics & Probability Letters, 87(??):158–166, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0167715214000303. See erratum [BRBB23].

Borges:2023:ECP

[BRBB23] P. Borges, J. Rodrigues, N. Balakrishnan, and J. Bazán. Erratum to “A COM-Poisson type generalization of the binomial distribution and its properties and applications” [Statist. Probab. Lett. 87 (2014) 158–166]. Statistics & Probability Letters, 201(??): ??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001074. See [BRBB14].

Browne:2022:RME

[Bro22] Ryan P. Browne. Revitalizing the multivariate elliptical leptokurtic-normal distribution and its application in model-based clustering. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001687.

Beknazaryan:2022:NRM

[BS22a] Aleksandr Beknazaryan and Hailin Sang. Nonparametric regression with modified ReLU networks. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001560.

Bhabak:2022:RSS

[BS22b] Arnab Bhabak and Subhamay Saha. Risk-sensitive semi-Markov decision problems with discounted cost and general utilities. Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000281.

Bouzebda:2022:NRM

[BS22c] Salim Bouzebda and Yousri Slaoui. Nonparametric recursive method for moment generating function kernel-type estimators. Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000359.

Bayramoglu:2024:APT

[BS24] Ismihan Bayramoglu and Alexei Stepanov. Asymptotic properties of m-th spacings based on records. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522400021X.

Braun:2020:DSF

[BSB20] W. John Braun, James Stafford, and Patrick Brown. Data sharpening via Firth’s adjusted score function. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301346.

Bandyopadhyay:2022:SCI

[BSB22] Uttam Bandyopadhyay, Suman Sarkar, and Atanu Biswas. Sequential confidence interval for comparing two Bernoulli distributions in a non-conventional set-up. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771522100225X.

Beran:2020:EPT

[BSST20] Jan Beran, Farzad Sabzikar, Donatas Surgailis, and Klaus Telkmann. On the empirical process of tempered moving averages. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302054.

Bhaskaran:2025:NIM

[BT25] Rajeev Bhaskaran and Stefan Tappe. A note on invariant manifolds for stochastic partial differential equations in the framework of the variational approach. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224002517.

Bilodeau:2023:TLA

[BTS23] Blair Bilodeau, Yanbo Tang, and Alex Stringer. On the tightness of the Laplace approximation for statistical inference. Statistics & Probability Letters, 198(??):??, July 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000639.

Bao:2021:EMS

[BTSL21] Zhenyu Bao, Jingwen Tang, Yan Shen, and Wei Liu. Equivalence of p-th moment stability between stochastic differential delay equations and their numerical methods. Statistics & Probability Letters, 168(??):??, January 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302558.

Barman:2022:BLP

[BU22] Kalyan Barman and Neelesh S. Upadhye. On Brascamp–Lieb and Poincaré type inequalities for generalized tempered stable distribution. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001444.

Barman:2024:SFL

[BU24] Kalyan Barman and Neelesh S. Upadhye. On Stein factors for Laplace approximation and their application to random sums. Statistics & Probability Letters, 206(??):??, March 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002201.

Budny:2022:IPI

[Bud22a] Katarzyna Budny. Improved probability inequalities for Mardia’s coefficient of kurtosis. Statistics & Probability Letters, 191(??): ??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001808.

Budny:2022:VPS

[Bud22b] Katarzyna Budny. The variance of the power of a shifted random variable with applications. Statistics & Probability Letters, 182 (??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002753.

Buonaguidi:2022:DPD

[Buo22] B. Buonaguidi. The disorder problem for diffusion processes with the ϵ-linear and expected total miss criteria. Statistics & Probability Letters, 189(??):??, October 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001092.

Burch:2021:FUC

[Bur21] Brent D. Burch. A family of unbiased confidence intervals for a ratio of variance components. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302686.

Boguslavskaya:2020:RIF

[BV20] Elena Boguslavskaya and Lioudmila Vostrikova. Revisiting integral functionals of geometric Brownian motion. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301371.

Bystrov:2023:EIN

[BV23] Alexander A. Bystrov and Nadezhda V. Volodko. Exponential inequalities for the number of subgraphs in the Erdős–Rényi random graph. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002760.

Besson:2020:SAC

[BVG20] Olivier Besson, François Vincent, and Xavier Gendre. A Stein’s approach to covariance matrix estimation using regularization of Cholesky factor and log-Cholesky metric. Statistics & Probability Letters, 167(??):??, December 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301966.

Bhaskaran:2023:DPE

[BW23] Aishwarya Bhaskaran and Matt P. Wand. Dispersion parameter extension of precise generalized linear mixed model asymptotics. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002048.

Cheng:2020:ODP

[CA20] Jing Cheng and Mingyao Ai. Optimal designs for panel data linear regressions. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300729.

Cui:2020:APN

[CB20] Jing Cui and Nana Bi. Averaging principle for neutral stochastic functional differential equations with impulses and non-Lipschitz coefficients. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030078X.

Chakraborty:2022:EAM

[CBK22] Saptarshi Chakraborty, Suman K. Bhattacharya, and Kshitij Khare. Estimating accuracy of the MCMC variance estimator: Asymptotic normality for batch means estimators. Statistics & Probability Letters, 183(??):??, April 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002868.

Can:2020:CLO

[CÇ20] Bugra Can and Mine Çağlar. Conditional law and occupation times of two-sided sticky Brownian motion. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301590.

Chen:2024:PLD

[CC24] Zhangting Chen and Dongya Cheng. Precise large deviations for non-centralized sums of partial sums and random sums of heavytailed END random variables. Statistics & Probability Letters, 211 (??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001032.

Chakrabarty:2020:RVF

[CCH20] Arijit Chakrabarty, Sukrit Chakraborty, and Rajat Subhra Hazra. Regular variation and free regular infinitely divisible laws. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302536.

Chen:2020:DCR

[CCY20] Tianyao Chen, Xue Cheng, and Jingping Yang. Decomposing correlated random walks on common and counter movements. Statistics & Probability Letters, 158(??):??, March 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302627.

Chen:2020:SLLa

[CD20a] Xiongzhi Chen and R. W. Doerge. A strong law of large numbers related to multiple testing normal means. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303396.

Chenavier:2020:ETR

[CD20b] Nicolas Chenavier and Ahmad Darwiche. Extremes for transient random walks in random sceneries under weak independence conditions. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219303037.

Chakraborty:2021:UCB

[CD21] Saptarshi Chakraborty and Swagatam Das. On uniform concentration bounds for bi-clustering by using the Vapnik–Chervonenkis theory. Statistics & Probability Letters, 175(??):??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522100064X.

Chakrabarty:2022:SCK

[CD22] Anish Chakrabarty and Swagatam Das. On strong consistency of kernel k-means: a Rademacher complexity approach. Statistics & Probability Letters, 182(??):??, March 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002534.

Cheng:2025:EMR

[CDLZ25] Jiang Cheng, Chao Deng, Shuyan Liu, and Xudong Zeng. An equilibrium model of reinsurance pricing. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224002499.

Capaldo:2024:GGM

[CDP24] Marco Capaldo, Antonio Di Crescenzo, and Franco Pellerey. Generalized Gini’s mean difference through distortions and copulas, and related minimizing problems. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223002055.

Chadjiconstantinidis:2024:DSM

[CE24] Stathis Chadjiconstantinidis and Serkan Eryilmaz. On δ-shock model with a change point in intershock time distribution. Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000154.

Christensen:2023:UFP

[CFH23] Sören Christensen, Simon Fischer, and Oskar Hallmann. Uniqueness of first passage time distributions via Fredholm integral equations. Statistics & Probability Letters, 203(??):??, December 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001360.

Chi:2023:RAB

[CH23a] Jui-Lin Chi and Jyy-I Hong. The range of asymmetric branching random walk. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002188.

Choi:2023:GMC

[CH23b] Michael C. H. Choi and Zhipeng Huang. Generalized Markov chain tree theorem and Kemeny’s constant for a class of non-Markovian matrices. Statistics & Probability Letters, 193(??): ??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002528.

Chauvet:2020:LSP

[Cha20] Guillaume Chauvet. Large sample properties of the Midzuno sampling scheme with probabilities proportional to size. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303268.

Chaari:2021:LDP

[Cha21] Sonia Chaari. Large deviations principle for white noise distributions with growth condition. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302881.

Chae:2024:WUB

[Cha24] Minwoo Chae. Wasserstein upper bounds of L^{p} -norms for multivariate densities in Besov spaces. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001007.

Chen:2020:SLLb

[Che20a] Xiongzhi Chen. A strong law of large numbers for simultaneously testing parameters of Lancaster bivariate distributions. Statistics & Probability Letters, 167(??):??, December 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302145.

Chen:2020:KTB

[Che20b] Yiqing Chen. A Kesten-type bound for sums of randomly weighted subexponential random variables. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219303074.

Chen:2021:UDS

[Che21] Li-Pang Chen. Ultrahigh-dimensional sufficient dimension reduction with measurement error in covariates. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302340.

Chen:2024:LDM

[Che24a] Zengcai Chen. Lower deviation for the maximum of two-speed branching Brownian motion. Statistics & Probability Letters, 206 (??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002262.

Cheng:2024:SMG

[Che24b] Dan Cheng. Smooth Matérn Gaussian random fields: Euler characteristic, expected number and height distribution of critical points. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000853.

Cheng:2024:IAE

[Che24c] Fuxia Cheng. The integrated absolute error of the kernel error distribution estimator in the first-order autoregression model.

Statistics & Probability Letters, 214(??):??, November 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001846.

Cheng:2024:OTA

[Che24d] Ziling Cheng. Occupation times for age-structured branching processes. Statistics & Probability Letters, 211(??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001147.

Chen:2025:HTB

[Che25] Yiqing Chen. The heavy-tail behavior of the difference of two dependent random variables. Statistics & Probability Letters, 218 (??):??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002761.

Chen:2025:BSB

[CHHS25] Yu Chen, Zongqing Hu, Jie Hu, and Lei Shu. Block structurebased covariance tensor decomposition for group identification in matrix variables. Statistics & Probability Letters, 216(??): ??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002207.

Chin:2022:DCL

[Chi22] Calvin Wooyoung Chin. Deriving the central limit theorem from the de Moivre–Laplace theorem. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002558.

Chen:2020:CTR

[CHW20] Ting-Li Chen, Su-Yun Huang, and Weichung Wang. A consistency theorem for randomized singular value decomposition. Statistics & Probability Letters, 161(??):??, June 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300468.

Cinque:2022:NCP

[Cin22] Fabrizio Cinque. A note on the conditional probabilities of the telegraph process. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000384.

Crucinio:2023:PMS

[CJ23] Francesca R. Crucinio and Adam M. Johansen. Properties of marginal sequential Monte Carlo methods. Statistics & Probability Letters, 203(??):??, December 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001384.

Chin:2020:NIC

[CJM20] Wooyoung Chin, Paul Jung, and Greg Markowsky. A note on invariance of the Cauchy and related distributions. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303141.

Choi:2023:AVU

[CK23] Woohyun Choi and Ilmun Kim. Averaging p-values under exchangeability. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002619.

Catania:2020:REL

[CL20] Leopoldo Catania and Alessandra Luati. Robust estimation of a location parameter with the integrated Hogg function. Statistics & Probability Letters, 164(??):??, September 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301152.

Cheng:2022:IHI

[CL22] Xueqin Cheng and Yanpeng Li. An improved Hoeffding’s inequality for sum of independent random variables. Statistics & Probability Letters, 183(??):??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002923.

Cui:2023:CTM

[CL23] Jiazhen Cui and Qiaojing Liu. Cramér-type moderate deviations for the log-likelihood ratio of inhomogeneous Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 192(??):??, January 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002036.

Chen:2023:CAP

[CLLP23] Baixi Chen, Zhiming Li, Zhi Li, and Can Peng. The calculation of alias pattern in three-level regular designs. Statistics & Probability Letters, 202(??):??, November 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001372.

Cheng:2022:GAH

[CLP22] Guanghui Cheng, Zhi Liu, and Liuhua Peng. Gaussian approximations for high-dimensional non-degenerate U-statistics via exchangeable pairs. Statistics & Probability Letters, 182(??): ??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002571.

Chen:2023:IVT

[CLP23] Cathy W. S. Chen, Feng-Chi Liu, and Aljo Clair Pingal. Integervalued transfer function models for counts that show zero inflation. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002140.

Chernova:2020:ADK

[CLR20] O. Chernova, F. Lavancier, and P. Rochet. Averaging of density kernel estimators. Statistics & Probability Letters, 158(??): ??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302913.

Cui:2021:NCI

[CLZZ21] Zhenyu Cui, Chihoon Lee, Lingjiong Zhu, and Yunfan Zhu. Nonconvex isotonic regression via the Myersonian approach. Statistics & Probability Letters, 179(??):??, December 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001723.

Cheng:2021:PCR

[CM21a] Hui-Hui Cheng and Yong-Hua Mao. Polynomial convergence for reversible jump processes. Statistics & Probability Letters, 173 (??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000432.

Cheng:2021:VFC

[CM21b] Zhi-Wen Cheng and Yong-Hua Mao. Variational formula of capacity for asymmetric Markov chains. Statistics & Probability Letters, 179(??):??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001929.

Chang:2023:GWL

[CM23] Mengmeng Chang and Yu Miao. Generalized weak laws of large numbers in Hilbert spaces. Statistics & Probability Letters, 197 (??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000548.

Chen:2020:COF

[CML20] Jianbin Chen, Rahul Mukerjee, and Dennis K. J. Lin. Construction of optimal fractional order-of-addition designs via block designs. Statistics & Probability Letters, 161(??):??, June 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300316.

Criens:2024:CMN

[CN24] David Criens and Lars Niemann. A class of multidimensional nonlinear diffusions with the Feller property. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000269.

Colwell:2023:HTN

[Col23] David B. Colwell. Hitting times, number of jumps, and occupation times for continuous-time finite state Markov chains. Statistics & Probability Letters, 195(??):??, April 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522300010X.

Caponera:2022:RCA

[CP22] Alessia Caponera and Victor M. Panaretos. On the rate of convergence for the autocorrelation operator in functional autoregression. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001274.

Chakraborty:2020:HCO

[CPD20] Saptarshi Chakraborty, Debolina Paul, and Swagatam Das. Hierarchical clustering with optimal transport. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300845.

Castellares:2020:GGL

[CPL20] Fredy Castellares, Silvio C. Patrício, and Artur J. Lemonte. On gamma-Gompertz life expectancy. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220301358.

Chen:2025:BEE

[CQ25] Wenxuan Chen and Zhi Qu. Berry–Esseen expansion and Cramértype large deviation for run and tumble particles on one dimension. Statistics & Probability Letters, 218(??):??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002773.

Christen:2025:HOH

[CR25] J. A. Christen and F. J. Rubio. On harmonic oscillator hazard functions. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002736.

Cruz:2020:SSD

[Cru20] Juan Alberto Rojas Cruz. Sensitivity of the stationary distributions of denumerable Markov chains. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301693.

Cheng:2020:CPG

[CS20] Dan Cheng and Armin Schwartzman. On critical points of Gaussian random fields under diffeomorphic transformations. Statistics & Probability Letters, 158(??):??, March 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303189.

Chen:2021:JSL

[CS21] Pingyan Chen and Soo Hak Sung. On the Jajte strong law of large numbers. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001000.

Chen:2024:ODS

[CS24] Qikun Chen and Michael Stewart. Optimal detection of sparse gamma scale admixture with twice the null mean. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400052X.

Cui:2021:NSI

[CT21] Yifan Cui and Eric Tchetgen Tchetgen. On a necessary and sufficient identification condition of optimal treatment regimes with an instrumental variable. Statistics & Probability Letters, 178(??): ??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001425.

Chae:2020:WUB

[CW20] Minwoo Chae and Stephen G. Walker. Wasserstein upper bounds of the total variation for smooth densities. Statistics & Probability Letters, 163(??):??, August 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300742.

Chen:2020:IIS

[CWW20] Zhenlong Chen, Jun Wang, and Dongsheng Wu. On intersections of independent space–time anisotropic Gaussian fields. Statistics & Probability Letters, 166(??):??, November 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301772.

Chen:2023:EHT

[CX23] Zhongyuan Chen and Jun Xie. Estimating heterogeneous treatment effects versus building individualized treatment rules: Connection and disconnection. Statistics & Probability Letters, 199 (??):??, August 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000780.

Chen:2024:HOD

[CX24] Zhenlong Chen and Peng Xu. Higher-order derivative of local times for space–time anisotropic Gaussian random fields. Statistics & Probability Letters, 214(??):??, November 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001664.

Chen:2021:PRH

[CY21] Xin Chen and Wenjie Ye. A probabilistic representation for heat flow of harmonic map on manifolds with time-dependent Riemannian metric. Statistics & Probability Letters, 177(??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001279.

Chen:2025:IBG

[CY25] Xiaohong Chen and Yanping Yi. Information bounds for Gaussian copula parameter in stationary semiparametric Markov models. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002232.

Cui:2023:WPM

[CZ23] Fengfeng Cui and Weidong Zhao. Well-posedness of mean reflected BSDEs with non-Lipschitz coefficients. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002310.

Chen:2024:EER

[CZ24] Ping Chen and Tusheng Zhang. Exponential ergodicity for reflected SDEs with interaction in a multidimensional general domain. Statistics & Probability Letters, 213(??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001378.

Chen:2024:PGW

[CZG24] Yong Chen, Shuolin Zhang, and Hongjun Gao. Probabilistic global well-posedness to the nonlocal Degasperis–Procesi equation. Statistics & Probability Letters, 206(??):??, March 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002249.

Chen:2020:MFF

[CZLC20] Xiaolin Chen, Yahui Zhang, Yi Liu, and Xiaojing Chen. Modelfree feature screening for ultra-high dimensional competing risks data. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301188.

Dembinska:2021:NOR

[DAA21] Anna Dembińska, Masoumeh Akbari, and Jafar Ahmadi. On numbers of observations in random regions determined by records for tail-less populations. Statistics & Probability Letters, 179(??): ??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001759.

Daly:2024:SMS

[Dal24] Fraser Daly. On Stein’s method for stochastically monotone singlebirth chains. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002171.

Dasgupta:2024:AHB

[Das24] Amites Dasgupta. Azuma–Hoeffding bounds for a class of urn models. Statistics & Probability Letters, 204(??):??, January 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001645.

Daw:2021:URD

[Daw21] Lara Daw. A uniform result for the dimension of fractional Brownian motion level sets. Statistics & Probability Letters, 169(??): ??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522030287X.

Das:2021:BFC

[DB21] Nabaneet Das and Subir Kumar Bhandari. Bound on FWER for correlated normal. Statistics & Probability Letters, 168(??): ??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302467.

Dhifaoui:2022:LCD

[DB22] Zouhaier Dhifaoui and Jean-Marc Bardet. Local correlation dimension of multidimensional stochastic process. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002248.

Dey:2023:FGZ

[DB23] Monitirtha Dey and Subir Kumar Bhandari. FWER goes to zero for correlated normal. Statistics & Probability Letters, 193(??): ??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002139.

Dornemann:2023:FDE

[DD23] Nina Dörnemann and Holger Dette. Fluctuations of the diagonal entries of a large sample precision matrix. Statistics & Probability Letters, 198(??):??, July 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000627.

Davezies:2022:MZL

[DDG22] Laurent Davezies, Xavier D’Haultfœuille, and Yannick Guyonvarch. The Marcinkiewicz–Zygmund law of large numbers for exchangeable arrays. Statistics & Probability Letters, 188(??):??, September 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001031.

DiCiccio:2020:ETM

[DDR20] Cyrus J. DiCiccio, Thomas J. DiCiccio, and Joseph P. Romano. Exact tests via multiple data splitting. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301681.

Demers:2021:EDN

[Dem21] Simon Demers. Expected duration of the no-information minimum rank problem. Statistics & Probability Letters, 168(??): ??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302534.

Caterina:2024:MSP

[DF24] Claudia Di Caterina and Davide Ferrari. Model selection by pathwise marginal likelihood thresholding. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001834.

Deng:2023:AEL

[DFG23a] Guantie Deng, Xiequan Fan, and Zhi-Qiang Gao. Asymptotic expansions in the local limit theorem for a branching Wiener process. Statistics & Probability Letters, 199(??):??, August 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000809.

Doukhan:2023:CMD

[DFG23b] Paul Doukhan, Xiequan Fan, and Zhi-Qiang Gao. Cramér moderate deviations for a supercritical Galton–Watson process. Statistics & Probability Letters, 192(??):??, January 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002243.

Dai:2020:NIC

[DH20] Shuang Dai and Zhensheng Huang. Nonparametric inference for covariate-adjusted model. Statistics & Probability Letters, 162 (??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300699.

Dumitrescu:2022:NEM

[DH22] Laura Dumitrescu and Darcy Harcourt. Nonparametric estimation of marginal distributions for unordered pairs. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002285.

Ding:2020:SST

[Din20] Xue Ding. Some sphericity tests for high dimensional data based on ratio of the traces of sample covariance matrices. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302597.

Ding:2021:FWL

[Din21] Peng Ding. The Frisch–Waugh–Lovell theorem for standard errors. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302480.

Demirci:2023:CTP

[DIY23] Yunus Emre Demirci, Ümiṫ Işlak, and Mehmet Akif Yıldız. On a cover time problem on a dynamic graph with steps at random times. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000457.

Diop:2022:ECP

[DK22] Mamadou Lamine Diop and William Kengne. Epidemic change-point detection in general causal time series. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000323.

Dai:2021:TCI

[DL21] Yin Dai and Ruinan Li. Transportation cost inequality for backward stochastic differential equations with mean reflection. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001292.

Dong:2024:NMC

[DL24] Yuexiao Dong and Zeda Li. A note on marginal coordinate test in sufficient dimension reduction. Statistics & Probability Letters, 204 (??):??, January 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001712.

Dai:2025:QBE

[DLLY25] Lichun Dai, Pengfei Liu, Yiming Liu, and Guangren Yang. The quantile-based empirical likelihood for the difference of quantiles. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002219.

Dau:2020:EAL

[DLS20] Hai Dang Dau, Thomas Laloë, and Rémi Servien. Exact asymptotic limit for kernel estimation of regression level sets. Statistics & Probability Letters, 161(??):??, June 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300249.

Dong:2022:RFS

[DLZZ22] Yan Dong, Daoji Li, Zemin Zheng, and Jia Zhou. Reproducible feature selection in high-dimensional accelerated failure time models. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002376.

Dolera:2020:UCP

[DM20] Emanuele Dolera and Edoardo Mainini. On uniform continuity of posterior distributions. Statistics & Probability Letters, 157(??): ??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302731.

Dedecker:2021:ASI

[DM21] Jérôme Dedecker and Florence Merlevède. Almost sure invariance principle for the Kantorovich distance between the empirical and the marginal distributions of strong mixing sequences. Statistics & Probability Letters, 171(??):??, April 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302947.

Dong:2024:CKM

[DM24] Aqi Dong and Volodymyr Melnykov. Contaminated Kent mixture model for clustering non-spherical directional data with heavy tails or scatter. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000270.

Fernandez-de-Marcos:2024:STS

[dMGP24] Alberto Fernández de Marcos and Eduardo García-Portugués. A stereographic test of spherical uniformity. Statistics & Probability Letters, 215(??):??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001871.

Noia:2024:BSS

[DML24] Antonio Di Noia, Gianluca Mastrantonio, and Giovanna Jona Lasinio. Bayesian size-and-shape regression modelling. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001529.

Djonguet:2024:TIF

[DMN24] Terence Kevin Manfoumbi Djonguet, Alban Mbina Mbina, and Guy Martial Nkiet. Testing independence of functional variables by an Hilbert–Schmidt independence criterion estimator. Statistics & Probability Letters, 207(??):??, April 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002390.

Noia:2023:CHT

[DMPP23] Antonio Di Noia, Marzia Marcheselli, Caterina Pisani, and Luca Pratelli. Censoring heavy-tail count distributions for parameter estimation with an application to stable distributions. Statistics & Probability Letters, 202(??):??, November 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522300127X.

Dette:2020:ODE

[DMS20a] Holger Dette, Viatcheslav B. Melas, and Petr Shpilev. Optimal designs for estimating individual coefficients in polynomial regression with no intercept. Statistics & Probability Letters, 158(??): ??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302822.

DiNardo:2020:NLD

[DMS20b] E. Di Nardo, M. Marena, and P. Semeraro. On non-linear dependence of multivariate subordinated Lévy processes. Statistics & Probability Letters, 166(??):??, November 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301735.

Dette:2021:NOD

[DMS21] Holger Dette, Viatcheslav B. Melas, and Petr Shpilev. A note on optimal designs for estimating the slope of a polynomial regression. Statistics & Probability Letters, 170(??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302959.

Do:2024:GBD

[DNH24] Quan Huu Do, Binh T. Nguyen, and Lam Si Tung Ho. A generalization bound of deep neural networks for dependent data. Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000294.

Dabo-Niang:2022:AES

[DNTV22] Sophie Dabo-Niang, Baba Thiam, and Thomas Verdebout. Asymptotic efficiency of some nonparametric tests for location on hyperspheres. Statistics & Probability Letters, 188(??):??, September 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000979.

Dobler:2022:SPL

[Döb22] Christian Döbler. A short proof of Lévy’s continuity theorem without using tightness. Statistics & Probability Letters, 185 (??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000438.

Dobler:2024:BEB

[Döb24] Christian Döbler. The Berry–Esseen bound in de Jong’s CLT. Statistics & Probability Letters, 215(??):??, December 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400213X.

Doumas:2022:MIC

[Dou22] Aristides V. Doumas. On the minimum of independent collecting processes via the Stirling numbers of the second kind. Statistics & Probability Letters, 185(??):??, June 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000372.

Dermitzakis:2022:MPS

[DP22] Vaios Dermitzakis and Konstadinos Politis. Monotonicity properties for solutions of renewal equations. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221001887.

Durre:2023:CIU

[DP23] Alexander Dürre and Davy Paindaveine. On the consistency of incomplete U-statistics under infinite second-order moments. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002279.

Diaz–Pachon:2020:HTA

[DPSR20] Daniel Andrés Díaz-Pachón, Juan Pablo Sáenz, and J. Sunil Rao. Hypothesis testing with active information. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300456.

Dixit:2021:PIS

[DR21] Anand Dixit and Vivekananda Roy. Posterior impropriety of some sparse Bayesian learning models. Statistics & Probability Letters, 171(??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221000018.

Davydov:2024:DAW

[DR24] Youri Davydov and Vladimir Rotar. The distribution of argmaximum or a winner problem. Statistics & Probability Letters, 211 (??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001214.

Drmota:2020:POR

[DS20a] Michael Drmota and Benedikt Stufler. Pattern occurrences in random planar maps. Statistics & Probability Letters, 158(??): ??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303128.

Dung:2020:RCC

[DS20b] Le Van Dung and Ta Cong Son. On the rate of convergence in the central limit theorem for arrays of random vectors. Statistics & Probability Letters, 158(??):??, March 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303177.

Drekic:2021:NTN

[DS21] Steve Drekic and Michael Z. Spivey. On the number of trials needed to obtain k consecutive successes. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000948.

Dai:2022:SOF

[DS22a] Guozheng Dai and Zhonggen Su. On the second order fluctuations for minimal difference partitions. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001225.

Duffield:2022:EKI

[DS22b] Samuel Duffield and Sumeetpal S. Singh. Ensemble Kalman inversion for general likelihoods. Statistics & Probability Letters, 187 (??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000967.

Djaloud:2024:NKE

[DS24] Toihir Soulaimana Djaloud and Cheikh Tidiane Seck. Nonparametric kernel estimation of conditional copula density. Statistics & Probability Letters, 212(??):??, September 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001238.

Dai:2025:CTS

[DS25] Shengtao Dai and Xiaojun Song. Consistent tests for semiparametric conditional independence. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224002220.

Derumigny:2024:LBT

[DSH24] Alexis Derumigny and Johannes Schmidt-Hieber. Lower bounds for the trade-off between bias and mean absolute deviation. Statistics & Probability Letters, 213(??):??, October 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001512.

Debaly:2021:NSM

[DT21a] Zinsou Max Debaly and Lionel Truquet. A note on the stability of multivariate non-linear time series with an application to time series of counts. Statistics & Probability Letters, 179(??): ??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001589.

DelSole:2021:CCA

[DT21b] Timothy DelSole and Michael K. Tippett. Correcting the corrected AIC. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000262.

Diez:2023:BET

[DT23] Charles-Philippe Diez and Ciprian A. Tudor. Berry–Essén theorem for random determinants. Statistics & Probability Letters, 197 (??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000287.

Dumbgen:2020:DRP

[DW20] Lutz Dümbgen and Jon A. Wellner. The density ratio of Poisson binomial versus Poisson distributions. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301656.

Ambroggio:2024:DEG

[DY24] Umberto De Ambroggio and Hiu Ching Yip. Degree evolution in a general growing network. Statistics & Probability Letters, 211 (??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001202.

Dmitriev:2021:MRF

[DZ21] Daniil Dmitriev and Maksim Zhukovskii. On monotonicity of Ramanujan function for binomial random variables. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001097.

Dong:2021:CFD

[DZZ21] Ruipeng Dong, Jia Zhou, and Zemin Zheng. Controlling the false discovery rate for latent factors via unit-rank deflation. Statistics & Probability Letters, 178(??):??, November 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001401.

Estala-Arias:2020:POM

[EA20] Samuel Estala-Arias. Pseudodifferential operators and Markov processes on certain totally disconnected groups. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301140.

El-Aroui:2021:RTP

[EA21] Mhamed-Ali El-Aroui. Real-time prequential goodness-of-fit testing of life distributions in renewal processes. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001085.

Efromovich:2024:NDE

[Efr24] Sam Efromovich. Nonparametric density estimation over its unknown support for right censored data. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224000531.

Eghbal:2021:SIA

[Egh21] Negar Eghbal. Some inequalities for absolute moments of feasible acceptable random variables. Statistics & Probability Letters, 172 (??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000195.

Egidi:2022:ESS

[Egi22] Leonardo Egidi. Effective sample size for a mixture prior. Statistics & Probability Letters, 183(??):??, April 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002856.

Estrada:2022:MEF

[EH22] Luisa F. Estrada and Michael A. Högele. Moment estimates in the first Borel–Cantelli lemma with applications to mean deviation frequencies. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001663.

Elster:2020:QIL

[EK20] Clemens Elster and Katy Klauenberg. A quantile inequality for location-scale distributions. Statistics & Probability Letters, 165 (??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301541.

Eisenberg:2022:IFS

[EK22] Julia Eisenberg and Paul Krühner. On Itô’s formula for semimartingales with jumps and non-C^{2} functions. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000049.

Edelmann:2021:RBP

[EMS21] Dominic Edelmann, Tamás F. Móri, and Gábor J. Székely. On relationships between the Pearson and the distance correlation coefficients. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302637.

Endo:2024:WAS

[EN24] Koya Endo and Yumiharu Nakano. Weak approximation of Schrödinger–Föllmer diffusion. Statistics & Probability Letters, 213(??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001408.

Edelmann:2023:PIM

[ERR23] Dominic Edelmann, Donald Richards, and Thomas Royen. Product inequalities for multivariate Gaussian, gamma, and positively upper orthant dependent distributions. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000445.

Fakhfakh:2024:CSK

[FA24] Raouf Fakhfakh and Omar Alzeley. Cauchy–Stieltjes Kernel families: Some properties. Statistics & Probability Letters, 209(??): ??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000646.

Fakhfakh:2020:VFB

[Fak20] Raouf Fakhfakh. Variance function of Boolean additive convolution. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300808.

Fakhfakh:2022:CMP

[Fak22] Raouf Fakhfakh. A characterization of the Marchenko–Pastur probability measure. Statistics & Probability Letters, 191(??): ??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522200178X.

Falconer:2022:IDI

[Fal22]  Kenneth J. Falconer.  Intermediate dimension of images

of sequences under fractional Brownian motion. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002625.

Fan:2024:SML

[Fan24] Xiequan Fan. Sharp moderate and large deviations for sample quantiles. Statistics & Probability Letters, 205(??):??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522300175X.

Finkelstein:2023:IAC

[FC23] Maxim Finkelstein and Ji Hwan Cha. Improving the age composition of dynamic populations of manufactured items. Statistics & Probability Letters, 201(??):??, October 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001037.

Farcomeni:2021:CMC

[FD21] Alessio Farcomeni and Francesco Dotto. A correction to make Chao estimator conservative when the number of sampling occasions is finite. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001164.

Forcina:2024:MTR

[FD24] Antonio Forcina and Valentino Dardanoni. Methods for testing the random utility model. Statistics & Probability Letters, 215(??): ??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001998.

Feriel:2021:NLL

[FE21] Bouhadjera Feriel and Ould Saïd Elias. Nonparametric local linear estimation of the relative error regression function for twice censored data. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001474.

Feng:2023:BKT

[Fen23] Fengxiang Feng. Baum–Katz-type complete and complete moment convergence theorems for the maximum of partial sums under sub-linear expectations. Statistics & Probability Letters, 197 (??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000421.

Ferrando:2021:OCB

[FEVC21] L. Ferrando, I. Epifanio, and N. Ventura-Campos. Ordinal classification of 3D brain structures by functional data analysis. Statistics & Probability Letters, 179(??):??, December 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001899.

Fu:2025:GMF

[FF25] Zongkui Fu and Dandan Fei. General mean-field reflected backward stochastic differential equations with locally monotone coefficients. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002426.

Forde:2022:RLF

[FFGS22] Martin Forde, Masaaki Fukasawa, Stefan Gerhold, and Benjamin Smith. The Riemann–Liouville field and its GMC as H → 0, and skew flattening for the rough Bergomi model. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002273.

Fazekas:2024:CRL

[FFS24] István Fazekas, Borbála Fazekas, and Michael Ochieng Suja. Convergence rate for the longest T -contaminated runs of heads. Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000282.

Fourdrinier:2021:CME

[FHM21] Dominique Fourdrinier, Anis M. Haddouche, and Fatiha Mezoued. Covariance matrix estimation under data-based loss. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522100122X.

Franke:2021:NAV

[FHSP21] B. Franke, C. R. Hwang, A. Ouled Said, and H. M. Pai. A note on the asymptotic variance of drift accelerated diffusions. Statistics & Probability Letters, 175(??):??, August 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000900.

Fernley:2024:URT

[FJ24] John Fernley and Emmanuel Jacob. A universal right tail upper bound for supercritical Galton–Watson processes with bounded offspring. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000518.

Formanov:2021:JPD

[FK21] Sh. K. Formanov and I. M. Khamdamov. On joint probability distribution of the number of vertices and area of the convex hulls generated by a Poisson point process. Statistics & Probability Letters, 169(??):??, February 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302698.

Feng:2023:ADD

[FLH23] Qunqiang Feng, Xing Li, and Zhishui Hu. Asymptotic degree distribution in a homogeneous evolving network model. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200253X.

Fan:2024:ESL

[FLLW24] Wujing Fan, Shihu Li, Yu Li, and Changxuan Wang. The exponential stability for locally monotone stochastic partial differential equations. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001992.

Fu:2022:PLD

[FLW22] Ke-Ang Fu, Yang Liu, and Jiangfeng Wang. Precise large deviations in a bidimensional risk model with arbitrary dependence between claim-size vectors and waiting times. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000025.

Feng:2023:LFT

[FLW23] Xingdong Feng, Qiaochu Liu, and Caixing Wang. A lackof-fit test for quantile regression process models. Statistics & Probability Letters, 192(??):??, January 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001936.

Foucart:2020:EIF

[FLZ20] Clément Foucart, Pei-Sen Li, and Xiaowen Zhou. On the entrance at infinity of Feller processes with no negative jumps. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301620.

Fan:2020:WDM

[FM20] Xiequan Fan and Xiaohui Ma. On the Wasserstein distance for a martingale central limit theorem. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301954.

Feng:2021:DPE

[FM21] Yarong Feng and Hosam M. Mahmoud. Dynamic Pólya– Eggenberger urns. Statistics & Probability Letters, 174(??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000511.

Fermin:2022:PCM

[FMR22] Lisandro Fermín, José Marcano, and Luis-Angel Rodríguez. A proof of consistency of the MLE for nonlinear Markov-switching AR processes. Statistics & Probability Letters, 183(??):??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002911.

Forcina:2023:MLL

[For23] Antonio Forcina. Marginal log-linear models and mediation analysis. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002449.

Forrester:2024:GDD

[For24] Peter J. Forrester. On the gamma difference distribution. Statistics & Probability Letters, 211(??):??, August 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001056.

Foss:2022:MFD

[FP22] Sergey Foss and Timofei Prasolov. Moments of the first descending epoch for a random walk with negative drift. Statistics & Probability Letters, 189(??):??, October 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001080.

Faouzi:2020:ZOP

[FPBK20] Tarik Faouzi, Emilio Porcu, Moreno Bevilacqua, and Igor Kondrashuk. Zastavnyi operators and positive definite radial functions. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302664.

Fulman:2020:SMN

[FR20] Jason Fulman and Adrian Röllin. Stein’s method and Narayana numbers. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301383.

Fried:2021:RUS

[Fri21] Sela Fried. On the restrictiveness of the usual stochastic order and the likelihood ratio order. Statistics & Probability Letters, 170 (??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220303151.

Fried:2022:RHR

[Fri22] Sela Fried. The restrictiveness of the hazard rate order and the moments of the maximal coordinate of a random vector uniformly distributed on the probability n-simplex. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000311.

Frolov:2021:ULB

[Fro21] Andrei N. Frolov. On upper and lower bounds for probabilities of combinations of events. Statistics & Probability Letters, 173 (??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000353.

Ferreira:2020:OBA

[FRR^{+}20] Paulo H. Ferreira, Eduardo Ramos, Pedro L. Ramos, Jhon F. B. Gonzales, Vera L. D. Tomazella, Ricardo S. Ehlers, Eveliny B. Silva, and Francisco Louzada. Objective Bayesian analysis for the Lomax distribution. Statistics & Probability Letters, 159(??): ??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303232.

Forde:2020:CLB

[FS20a] Martin Forde and Benjamin Smith. The conditional law of the Bacry–Muzy and Riemann–Liouville log correlated Gaussian fields and their GMC, via Gaussian Hilbert and fractional Sobolev spaces. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300353.

Forsstrom:2020:FSI

[FS20b] Malin Palö Forsström and Jeffrey E. Steif. A few surprising integrals. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302810.

Fuchs:2021:OSK

[FS21a] Sebastian Fuchs and Klaus D. Schmidt. On order statistics and Kendall’s tau. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302753.

Fung:2021:TAB

[FS21b] Thomas Fung and Eugene Seneta. Tail asymptotics for the bivariate equi-skew generalized hyperbolic distribution and its Variance–Gamma special case. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221001449.

Fung:2024:TAS

[FS24] Thomas Fung and Eugene Seneta. A theorem on the asymptotics of skew-normal type integrals. Statistics & Probability Letters, 209 (??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522400049X.

Fulman:2024:BAT

[Ful24] Jason Fulman. Beta approximation for the two alleles Moran model by Stein’s method. Statistics & Probability Letters, 208 (??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000208.

Ferrer:2025:ESS

[FV25] Clemente Ferrer and Ronny Vallejos. Is the effective sample size always less than n ? A spatial regression approach. Statistics & Probability Letters, 218(??):??, March 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002785.

Feng:2020:LMD

[FWY20] Tingting Feng, Shaochen Wang, and Guangyu Yang. Large and moderate deviation principles for the Erdős–Kac theorem in function fields. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300833.

Feng:2024:MIA

[FY24] Decheng Feng and Gaihua Yang. Maximal inequalities and applications for reverse demimartingales based on concave Young functions. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000105.

Fujita:2024:SMP

[FYY24] Takahiko Fujita, Shotaro Yagishita, and Naohiro Yoshida. Some martingale properties of the simple random walk and its maximum process. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000452.

Feng:2023:MIS

[FZ23] Decheng Feng and Xiaomei Zhang. Maximal inequalities and the strong law of large numbers for strong demimartingales. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002218.

Ganesan:2020:EPI

[Gan20] Ghurumuruhan Ganesan. Extremal paths in inhomogeneous random graphs. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302391.

Ganesan:2021:DEE

[Gan21a] Ghurumuruhan Ganesan. Deviation estimates for Eulerian edit numbers of random graphs. Statistics & Probability Letters, 171 (??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522030328X.

Ganesan:2021:NCI

[Gan21b] Ghurumuruhan Ganesan. Nominal correlation of inhomogeneous random sequences. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302595.

Gao:2021:ECR

[Gao21] Zhi-Qiang Gao. Exact convergence rate in the central limit theorem for a branching process in a random environment. Statistics & Probability Letters, 178(??):??, November 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001565.

Gao:2024:LDM

[Gao24] Yunshi Gao. Large deviations for mean field model in Erdős– Rényi graph. Statistics & Probability Letters, 205(??):??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001773.

Gapeev:2020:OSP

[Gap20]  Pavel V. Gapeev.  Optimal stopping problems for run-

ning minima with positive discounting rates. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302029.

Gatto:2024:SJP

[Gat24] R. Gatto. Stationary jump processes for planar directions obtained by wrapping. Statistics & Probability Letters, 205(??): ??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001797.

Gaunt:2023:MVG

[Gau23] Robert E. Gaunt. On the moments of the variance-gamma distribution. Statistics & Probability Letters, 201(??):??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001086.

Gajda:2021:PLP

[GB21] Janusz Gajda and Luisa Beghin. Prabhakar Lévy processes. Statistics & Probability Letters, 178(??):??, November 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001243.

Guo:2024:CPB

[GCW24] Bing Guo, Xueping Chen, and Xiaodi Wang. Component projection balanced designs for order of addition experiments. Statistics & Probability Letters, 211(??):??, August 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001159.

Guo:2021:QSE

[GGX21] Junyu Guo, Xiaotian Guo, and Longjie Xie. Quantitative stability estimates for multiscale stochastic dynamical systems. Statistics & Probability Letters, 178(??):??, November 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001553.

Gagnon:2023:TPB

[GH23] Philippe Gagnon and Yoshiko Hayashi. Theoretical properties of Bayesian Student-t linear regression. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002061.

Guerbyenne:2024:SVM

[GHH24] Hafida Guerbyenne, Fayçal Hamdi, and Malika Hamrat. The log GARCH stochastic volatility model. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001548.

Ghosh:2023:OGU

[Gho23a] Abhik Ghosh. Optimal guessing under nonextensive framework and associated moment bounds. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000366.

Ghosh:2023:ICC

[Gho23b] Indranil Ghosh. On the issue of convergence of certain divergence measures related to finding most nearly compatible probability distribution under the discrete set-up. Statistics & Probability Letters, 203(??):??, December 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001396.

Galganov:2024:SCR

[GI24] Oleksii Galganov and Andrii Ilienko. Short cycles of random permutations with cycle weights: Point processes approach. Statistics & Probability Letters, 213(??):??, October 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400138X.

Girard:2020:ANE

[Gir20] Didier A. Girard. Asymptotic near-efficiency of the “Gibbsenergy (GE) and empirical-variance” estimating functions for fitting Matérn models — II: Accounting for measurement errors via “conditional GE mean”. Statistics & Probability Letters, 162 (??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300298.

Giraudo:2024:SNE

[Gir24] Davide Giraudo. Some notes on ergodic theorem for U-statistics of order m for stationary and not necessarily ergodic sequences. Statistics & Probability Letters, 210(??):??, July 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000865.

Gajek:2020:ASP

[GK20] Lesław Gajek and Elżbieta Krajewska. Approximating sums of products of dependent random variables. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301061.

Gupta:2022:ITS

[GK22] Neha Gupta and Arun Kumar. Inverse tempered stable subordinators and related processes with Mellin transform. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200058X.

Guo:2021:EEC

[GL21a] Xianping Guo and Zhong-Wei Liao. Estimate the exponential convergence rate of f-ergodicity via spectral gap. Statistics & Probability Letters, 168(??):??, January 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302273.

Guo:2021:LLN

[GL21b] Xiaofan Guo and Xinpeng Li. On the laws of large numbers for pseudo-independent random variables under sublinear expectation. Statistics & Probability Letters, 172(??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000043.

Gao:2023:LDA

[GLL23a] Qingwu Gao, Jiańan Lin, and Xijun Liu. Large deviations of aggregate amount of claims in compound risk model with arbitrary dependence between claim sizes and waiting times. Statistics & Probability Letters, 197(??):??, June 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000330.

Guo:2023:NPI

[GLL23b] Xiaofan Guo, Shan Li, and Xinpeng Li. Notes on Peng’s independence in sublinear expectation theory. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222002322.

Gomtsyan:2024:SCE

[GLLOS24] M. Gomtsyan, C. Lévy-Leduc, S. Ouadah, and L. Sansonnet. Sign consistent estimation in a sparse Poisson model. Statistics & Probability Letters, 209(??):??, June 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000762.

Gaunt:2025:SCD

[GLS25] Robert E. Gaunt, Siqi Li, and Heather L. Sutcliffe. A Stein characterisation of the distribution of the product of correlated normal random variables. Statistics & Probability Letters, 216(??): ??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002384.

Gregoratti:2021:LSV

[GM21] M. Gregoratti and D. Maran. Least singular value and condition number of a square random matrix with i.i.d. rows. Statistics & Probability Letters, 173(??):??, June 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000328.

Gangopadhyay:2022:ASC

[GM22] Ujan Gangopadhyay and Krishanu Maulik. Almost sure convergence of randomized urn models with application to elephant random walk. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001699.

Glavas:2024:EIS

[GM24] Lenka Glavaš and Pavle Mladenović. Extremes in incomplete samples from moving averages of random variables from the domain of attraction of the Gumbel distribution. Statistics & Probability Letters, 207(??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002353.

Ghosh:2023:CPO

[GN23] Sugata Ghosh and Asok K. Nanda. Conditional precedence orders for stochastic comparison of random variables. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002152.

Gokhale:2025:WPS

[Gok25] Soham Gokhale. Well-posedness for the stochastic Landau– Lifshitz–Gilbert equation with helicity driven by jump noise. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002542.

Ghodrati:2023:MRO

[GP23] Laya Ghodrati and Victor M. Panaretos. Minimax rate for optimal transport regression between distributions. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002711.

Gutierrez-Pavon:2020:DLT

[GPP20] Jonathan Gutierrez-Pavón and Carlos G. Pacheco. A density for the local time of the Brox diffusion. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300948.

Gutierrez-Pavon:2022:QDM

[GPP22] Jonathan Gutierrez-Pavón and Carlos G. Pacheco. Quenched distributions for the maximum, minimum and local time of the Brox diffusion. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002005.

Gava:2021:SLT

[GR21] Renato J. Gava and Bruna L. F. Rezende. Some limit theorems for dependent Bernoulli random variables. Statistics & Probability Letters, 170(??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303138.

Grabchak:2021:EMS

[Gra21] Michael Grabchak. An exact method for simulating rapidly decreasing tempered stable distributions in the finite variation case. Statistics & Probability Letters, 170(??):??, March 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303187.

Grahovac:2022:IST

[Gra22] Danijel Grahovac. Intermittency in the small-time behavior of Lévy processes. Statistics & Probability Letters, 187(??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000852.

Galimberti:2020:NCM

[GS20] Giuliano Galimberti and Gabriele Soffritti. A note on the consistency of the maximum likelihood estimator under multivariate linear cluster-weighted models. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219302767.

Gut:2021:NZE

[GS21] Allan Gut and Ulrich Stadtmüller. The number of zeros in elephant random walks with delays. Statistics & Probability Letters, 174(??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000742.

Gairat:2022:SBM

[GS22a] Alexander Gairat and Vadim Shcherbakov. Skew Brownian motion with dry friction: Joint density approach. Statistics & Probability Letters, 187(??):??, August 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000876.

Gut:2022:ERW

[GS22b] Allan Gut and Ulrich Stadtmüller. The elephant random walk with gradually increasing memory. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001432.

Guerrero:2020:GMD

[GSL20] Victor M. Guerrero and Claudia Solis-Lemus. A generalized measure of dispersion. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301097.

Grygierek:2020:GFE

[GT20] Jens Grygierek and Christoph Thäle. Gaussian fluctuations for edge counts in high-dimensional random geometric graphs. Statistics & Probability Letters, 158(??):??, March 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303207.

Guo:2024:SNB

[GTH24] Zheng-Yan Guo, Ze-Yu Tao, and Ze-Chun Hu. A study on the negative binomial distribution motivated by Chvátal’s theorem. Statistics & Probability Letters, 207(??):??, April 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000063.

Grazian:2020:LBP

[GVL20] Clara Grazian, Cristiano Villa, and Brunero Liseo. On a lossbased prior for the number of components in mixture models. Statistics & Probability Letters, 158(??):??, March 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303025.

Gaunt:2020:SMS

[GW20] Robert E. Gaunt and Neil Walton. Stein’s method for the single server queue in heavy traffic. Statistics & Probability Letters, 156 (??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302123.

Ghaffari:2021:PIO

[GW21] N. Ghaffari and S. G. Walker. Parseval’s identity and optimal transport maps. Statistics & Probability Letters, 170(??): ??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302923.

Ghaffari:2023:BRR

[GW23] N. Ghaffari and S. G. Walker. W_{2} barycenters for radially related distributions. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000123.

Gao:2022:TEM

[GWWY22] Xiangyu Gao, Jianqiao Wang, Yanxia Wang, and Hongfu Yang. The truncated Euler–Maruyama method for CIR model driven by fractional Brownian motion. Statistics & Probability Letters, 189 (??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001262.

Geng:2022:WCM

[GX22] Xi Geng and Aihua Xia. When is the Conway–Maxwell–Poisson distribution infinitely divisible? Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002261.

Gao:2022:MDR

[GYZ22] Yanping Gao, Siyu Yi, and Yongdao Zhou. Maximin L_{1}- distance range-fixed level-augmented designs. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222000621.

Geng:2024:DEL

[GZ24] Shuli Geng and Lixin Zhang. Decorrelated empirical likelihood for generalized linear models with high-dimensional longitudinal data. Statistics & Probability Letters, 211(??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001044.

Guo:2020:ICM

[GZH20] Junhao Guo, Jie Zhou, and Sanfeng Hu. Intrinsic covariance matrix estimation for multivariate elliptical distributions. Statistics & Probability Letters, 162(??):??, July 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300778.

Gzyl:2021:FHO

[Gzy21a] Henryk Gzyl. Forced harmonic oscillators, waves on a forced string and changes of measure. Statistics & Probability Letters, 179(??): ??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001942.

Gzyl:2021:HOW

[Gzy21b] Henryk Gzyl. Harmonic oscillators, waves and Gaussian processes. Statistics & Probability Letters, 172(??):??, May 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000055.

Husain:2024:WSC

[HA24] Bushra Husain and Fariha Aslam. Weighted Simplex Centroid Mixture Experiments for third order Becker’s models: the Roptimal approach. Statistics & Probability Letters, 213(??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001585.

Huong:2025:ABS

[HA25] P. T. Huong and P. T. Anh. On the asymptotic behavior of solutions to bilinear Caputo stochastic fractional differential equations. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002414.

Hafouta:2023:ASI

[Haf23] Yeor Hafouta. An almost sure invariance principle for some classes of non-stationary mixing sequences. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222002413.

Hagele:2020:PAR

[Häg20] Miriam Hägele. Precise asymptotics of ruin probabilities for a class of multivariate heavy-tailed distributions. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301747.

Hahn:2020:ERM

[Hah20] Georg Hahn. On the expected runtime of multiple testing algorithms with bounded error. Statistics & Probability Letters, 165 (??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301474.

Hahn:2022:OMC

[Hah22] Georg Hahn. Online multivariate changepoint detection with type I error control and constant time/memory updates per series. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002200.

Hashimoto:2021:PPM

[Has21] Shintaro Hashimoto. Predictive probability matching priors for a certain non-regular model. Statistics & Probability Letters, 174 (??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000584.

Hsieh:2020:SFE

[HC20] Jin-Jian Hsieh and Yung-Yu Chen. Survival function estimation of current status data with dependent censoring. Statistics & Probability Letters, 157(??):??, February 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302676.

Hua:2022:MAR

[HC22] Chen-Wei Hua and Ting-Li Chen. On multiple acceleration of reversible Markov chain. Statistics & Probability Letters, 189(??): ??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001195.

Hinojosa-Calleja:2023:EUM

[HC23] Adrián Hinojosa-Calleja. Exact uniform modulus of continuity for q-isotropic Gaussian random fields. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223000378.

Hsu:2022:DBO

[HCW22] Yu-Sheng Hsu, Pei-Chun Chen, and Cheng-Hsun Wu. Doublebarrier option pricing equations under extended geometric Brownian motion with bankruptcy risk. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000116.

Ho:2022:SMO

[HD22] Lam Si Tung Ho and Vu Dinh. Searching for minimal optimal neural networks. Statistics & Probability Letters, 183(??): ??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002947.

Henzi:2023:SNI

[HD23a] Alexander Henzi and Lutz Dümbgen. Some new inequalities for beta distributions. Statistics & Probability Letters, 195(??): ??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522300007X.

Hima:2023:LDP

[HD23b] Abdoulaye Soumana Hima and Ibrahim Dakaou. Large deviation principle for Reflected Stochastic Differential Equations driven by G-Brownian motion in non-convex domains. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002206.

Hussain:2024:RRM

[HDC24] Abid Hussain, Steve Drekic, and Salman A. Cheema. A relative-rank measure for the rank transformation. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001566.

He:2024:MDM

[He24] Wei He. Multi-dimensional mean-reflected BSDEs driven by G-Brownian motion with time-varying non-Lipschitz coefficients. Statistics & Probability Letters, 206(??):??, March 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002018.

Hazra:2022:HFR

[HFC22] Nil Kamal Hazra, Maxim Finkelstein, and Ji Hwan Cha. On a hazard (failure) rate process with delays after shocks. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002388.

He:2020:ODI

[HH20] Lei He and Daojiang He. R-optimal designs for individual prediction in random coefficient regression models. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771521930330X.

Hildebrandt:2020:GFD

[Hil20] Florian Hildebrandt. On generating fully discrete samples of the stochastic heat equation on an interval. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300535.

Hirao:2021:FFF

[Hir21] Masatake Hirao. Finite frames, frame potentials and determinantal point processes on the sphere. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000912.

Hamura:2024:PRM

[HIS24] Yasuyuki Hamura, Kaoru Irie, and Shonosuke Sugasawa. Posterior robustness with milder conditions: Contamination models revisited. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000993.

Huang:2021:FSL

[HJKL21] Jian Huang, Yuling Jiao, Lican Kang, and Yanyan Liu. Fitting sparse linear models under the sufficient and necessary condition for model identification. Statistics & Probability Letters, 168(??): ??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302285.

Hilbert:2020:RSD

[HJO20] Astrid Hilbert, Imane Jarni, and Youssef Ouknine. On reflected stochastic differential equations driven by regulated semimartingales. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302157.

Hashorva:2021:MMS

[HK21] Enkelejd Hashorva and Alfred Kume. Multivariate max-stable processes and homogeneous functionals. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221000286.

Hazelton:2024:WLB

[HK24] Martin L. Hazelton and Masoud Karimi. When lattice bases are Markov bases. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000750.

Herve:2021:APM

[HL21a] Loïc Hervé and James Ledoux. Asymptotic of products of Markov kernels. application to deterministic and random forward/ backward products. Statistics & Probability Letters, 179(??):??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001668.

Huang:2021:LBM

[HL21b] Chunfeng Huang and Ao Li. On Lévy’s Brownian motion and white noise space on the circle. Statistics & Probability Letters, 171 (??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000031.

He:2020:ECS

[HLL20a] Lidan He, Qiang Liu, and Zhi Liu. Edgeworth corrections for spot volatility estimator. Statistics & Probability Letters, 164(??): ??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301127.

Hong:2020:ALH

[HLL20b] Wei Hong, Shihu Li, and Wei Liu. Asymptotic log-Harnack inequality and applications for SPDE with degenerate multiplicative noise. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301139.

Hu:2024:MLS

[HLS24] Bo Hu, Dennis K. J. Lin, and Fasheng Sun. Mixed-level screening designs based on skew-symmetric conference matrices. Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000488.

Huang:2022:BEB

[HLX22] Xulan Huang, Yingqiu Li, and Kainan Xiang. Berry–Esseen bound for a supercritical branching processes with immigration in a random environment. Statistics & Probability Letters, 190(??): ??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001535.

Hansen:2021:QCU

[HMRT21] Joel Hansen, Itzamar Marquez, Mrinal K. Roychowdhury, and Eduardo Torres. Quantization coefficients for uniform distributions on the boundaries of regular polygons. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000225.

Hyodo:2020:EBH

[HNP20] Masashi Hyodo, Takahiro Nishiyama, and Tatjana Pavlenko. On error bounds for high-dimensional asymptotic distribution of L_{2}-type test statistic for equality of means. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302834.

Hirsch:2023:APO

[HNS23] Christian Hirsch, Matthias Neumann, and Volker Schmidt. Asymptotic properties of one-layer artificial neural networks with sparse connectivity. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002115.

Huang:2025:MPG

[HP25] Ying Huang and Jun Peng. Minimizing the penalized goalreaching probability with multiple dependent risks. Statistics & Probability Letters, 217(??):??, February 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002566.

Hu:2024:SNA

[HPH24] Pingbo Hu, Xiuyuan Peng, and Xinglin Hu. Some new asymptotic results for series estimation under clustered dependence. Statistics & Probability Letters, 213(??):??, October 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001251.

Hu:2022:TDF

[HPN22] Shuang Hu, Zuoxiang Peng, and Saralees Nadarajah. Tail dependence functions of the bivariate Hüsler–Reiss model. Statistics & Probability Letters, 180(??):??, January 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001978.

Hashorva:2020:ADS

[HR20] Enkelejd Hashorva and Didier Rullière. Asymptotic domination of sample maxima. Statistics & Probability Letters, 160(??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300067.

Hassairi:2022:ERP

[HR22] Abdelhamid Hassairi and Amel Roula. Exponential and related probability distributions on symmetric matrices. Statistics & Probability Letters, 187(??):??, August 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000815.

Hossjer:2022:SLU

[HS22] Ola Hössjer and Arvid Sjölander. Sharp lower and upper bounds for the covariance of bounded random variables. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002790.

He:2022:ODD

[HSS22] Ping He, Yuncong Shen, and Wenjie Sun. One-dimensional diffusion and stochastic differential equation. Statistics & Probability Letters, 183(??):??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002844.

Hille:2021:EEC

[HSWZ21] Sander C. Hille, Tomasz Szarek, Daniel T. H. Worm, and Maria A. Ziemlańska. Equivalence of equicontinuity concepts for Markov operators derived from a Schur-like property for spaces of measures. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302674.

Han:2021:PAV

[HSZ21] Xiyue Han, Alexander Schied, and Zhenyuan Zhang. A probabilistic approach to the Φ-variation of classical fractal functions with critical roughness. Statistics & Probability Letters, 168(??): ??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302236.

Hooshangifar:2021:BOD

[HT21] M. Hooshangifar and H. Talebi. Bayesian optimal design for non-linear model under non-regularity condition. Statistics & Probability Letters, 169(??):??, February 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302789.

Horvath:2023:FCP

[HT23a] Lajos Horváth and Lorenzo Trapani. L_{p}-functionals for change point detection in random coefficient autoregressive models. Statistics & Probability Letters, 201(??):??, October 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000536.

Huong:2023:WPR

[HT23b] P. T. Huong and N. T. The. Well-posedness and regularity for solutions of Caputo stochastic fractional delay differential equations. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002814.

Huang:2023:NEG

[Hua23] Chien-Hao Huang. Nonsymmetric examples for Gaussian correlation inequalities. Statistics & Probability Letters, 201(??): ??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001098.

Huang:2024:LSM

[Hua24a] Dongzhou Huang. The “logarithmic scale” Minkowski dimension of the most visited sites of two-dimensional Brownian motion. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001761.

Huang:2024:NCA

[Hua24b] Xiangyu Huang. On negative correlation of Arboreal Gas for specific parameters. Statistics & Probability Letters, 213(??): ??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001433.

Hughes:2021:OED

[Hug21] John Hughes. On the occasional exactness of the distributional transform approximation for direct Gaussian copula models with discrete margins. Statistics & Probability Letters, 177(??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001218.

Hughes:2024:BTB

[Hug24] John Hughes. A Bayesian tour of binomial inference. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001980.

Hu:2020:ASB

[HW20] Shulan Hu and Ran Wang. Asymptotics of stochastic Burgers equation with jumps. Statistics & Probability Letters, 162 (??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300730.

Huang:2023:SCR

[HW23a] Chuying Huang and Xu Wang. Strong convergence rate of the Euler scheme for SDEs driven by additive rough fractional noises. Statistics & Probability Letters, 194(??):??, March 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002553.

Huang:2023:DEE

[HW23b] Lu-Jing Huang and Tao Wang. Dirichlet eigenvalues and exit time moments for symmetric Markov processes. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002176.

Han:2024:HCS

[HW24] Yuecai Han and Guanyu Wu. Hölder continuity of stochastic heat equation with rough Gaussian noise. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224000889.

Hwang:2020:RTS

[Hwa20] Leng-Cheng Hwang. A robust two-stage procedure for the Poisson process under the linear exponential loss function. Statistics & Probability Letters, 163(??):??, August 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300766.

Hwang:2021:WLS

[Hwa21] Eunju Hwang. Weighted least squares estimation in a binary random coefficient panel model with infinite variance. Statistics & Probability Letters, 168(??):??, January 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302352.

Han:2023:GOF

[HWW23] Yong Han, Yuefei Wang, and Zipeng Wang. On the GFF with one free boundary condition. Statistics & Probability Letters, 193(??): ??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002127.

Hao:2022:SCM

[HWX22] Tao Hao, Jiaqiang Wen, and Jie Xiong. Solvability of a class of mean-field BSDEs with quadratic growth. Statistics & Probability Letters, 191(??):??, December 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001742.

Hong:2021:DLT

[HX21a] Minhao Hong and Fangjun Xu. Derivatives of local times for some Gaussian fields II. Statistics & Probability Letters, 172 (??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000250.

Hu:2021:EAP

[HX21b] Yaozhong Hu and Yuejuan Xi. Estimation of all parameters in the reflected Ornstein–Uhlenbeck process from discrete observations. Statistics & Probability Letters, 174(??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000614.

He:2023:MDH

[HX23] Yuheng He and Xiaofeng Xue. Moderate deviations of hitting times of a family of density-dependent Markov chains. Statistics & Probability Letters, 195(??):??, April 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000044.

Han:2023:NAC

[HY23] Yuecai Han and Zhe Yin. A note on asymptotics of classical likelihood ratio tests for high-dimensional normal distributions. Statistics & Probability Letters, 199(??):??, August 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000834.

He:2021:EMP

[HZY21] Guoman He, Hanjun Zhang, and Gang Yang. Exponential mixing property for absorbing Markov processes. Statistics & Probability Letters, 179(??):??, December 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001693.

Hao:2021:ROD

[HZZZ21] Honghua Hao, Xiaoyuan Zhu, Xinfeng Zhang, and Chongqi Zhang. R-optimal design of the second-order Scheffé mixture model. Statistics & Probability Letters, 173(??):??, June 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000316.

Idais:2020:NLO

[Ida20] Osama Idais. A note on locally optimal designs for generalized linear models with restricted support. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219303384.

Iksanov:2021:FLT

[IK21] Alexander Iksanov and Oleh Kondratenko. Functional limit theorems for discounted exponential functional of random walk and discounted convergent perpetuity. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001103.

Il-Kwang:2025:PFL

[IKCSKH25] Paek Il-Kwang, Kang Chol-Su, and Kim Kyong-Hui. Pricing formula of lookback option in stochastic delay differential equation model. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002529.

Iafrate:2020:ARL

[IM20] Francesco Iafrate and Claudio Macci. Asymptotic results for the last zero crossing time of a Brownian motion with nonnull drift. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030184X.

Iuliano:2023:ARA

[IM23] Antonella Iuliano and Claudio Macci. Asymptotic results for the absorption time of telegraph processes with a non-standard barrier at the origin. Statistics & Probability Letters, 196(??):??, May 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522300024X.

Imamura:2024:LLA

[IMT24] Takumi Imamura, Hiroki Masuda, and Hayato Tajima. On local likelihood asymptotics for Gaussian mixed-effects model with system noise. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000439.

Iafrate:2021:STG

[IO21] F. Iafrate and E. Orsingher. On the sojourn time of a generalized Brownian meander. Statistics & Probability Letters, 168(??): ??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302303.

Itkin:2024:GRD

[Itk24] David Itkin. Generalized Rank Dirichlet distributions. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001748.

Inverardi:2024:IHM

[ITM24] Pier Luigi Novi Inverardi, Aldo Tagliani, and Mariyan Milev. Indeterminate Hamburger moment problem: Entropy convergence. Statistics & Probability Letters, 212(??):??, September 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522400124X.

Ip:2020:NDM

[IW20] Ryan H. L. Ip and K. Y. K. Wu. A note on discrete multivariate Markov random field models. Statistics & Probability Letters, 156 (??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302342.

Izmirlian:2020:SCA

[Izm20] Grant Izmirlian. Strong consistency and asymptotic normality for quantities related to the Benjamini–Hochberg false discovery rate procedure. Statistics & Probability Letters, 160(??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030016X.

Jones:2021:RPP

[JA21] Ben Jones and Andreas Artemiou. Revisiting the predictive power of kernel principal components. Statistics & Probability Letters, 171(??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220303229.

Jakubowski:2021:CCI

[Jak21] Adam Jakubowski. A complement to the Chebyshev integral inequality. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302376.

Janson:2021:PBV

[Jan21] Svante Janson. On the probability that a binomial variable is at most its expectation. Statistics & Probability Letters, 171 (??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220303230.

Janson:2024:CLT

[Jan24] Svante Janson. On a central limit theorem in renewal theory. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001724.

Jasinski:2020:SCM

[Jas20] Krzysztof Jasiński. Some concept of Markov property of discrete order statistics arising from independent and non-identically distributed variables. Statistics & Probability Letters, 160(??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300213.

Jammalamadaka:2020:MCM

[JB20] S. Rao Jammalamadaka and Sudeep R. Bapat. Middle censoring in the multinomial distribution with applications. Statistics & Probability Letters, 167(??):??, December 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302194.

Jones:2021:SFI

[JB21] M. C. Jones and N. Balakrishnan. Simple functions of independent beta random variables that follow beta distributions. Statistics & Probability Letters, 170(??):??, March 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522030314X.

Jones:2025:AMB

[JB25] M. C. Jones and N. Balakrishnan. On absolute moment-based upper bounds for L-moments. Statistics & Probability Letters, 216 (??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002189.

Jammalamadaka:2024:TUC

[JGA24] S. Rao Jammalamadaka, Kaushik Ghosh, and Sridhar Akiri. Testing uniformity on the circle using spacings when data are rounded. Statistics & Probability Letters, 210(??):??, July 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400083X.

Jin:2022:CRO

[JH22] Zhumengmeng Jin and James P. Hobert. On the convergence rate of the “out-of-order” block Gibbs sampler. Statistics & Probability Letters, 188(??):??, September 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001043.

Jacobovic:2022:CNC

[JK22] Royi Jacobovic and Offer Kella. A characterization of normality via convex likelihood ratios. Statistics & Probability Letters, 186 (??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000542.

Ji:2025:SCM

[JL25] Qing Ji and Jicheng Liu. Strong convergence of multi-scale stochastic differential equations with a full dependence. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002402.

Jevtic:2020:FSM

[JLL20] Petar Jevtić, Nicolas Lanchier, and Axel La Salle. First and second moments of the size distribution of bond percolation clusters on rings, paths and stars. Statistics & Probability Letters, 161 (??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300171.

Jie:2024:ODM

[JLZ24] Lijuan Jie, Liangqing Luo, and Hua Zhang. One-dimensional McKean–Vlasov stochastic Volterra equations with Hölder diffusion coefficients. Statistics & Probability Letters, 205(??):??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001943.

Jones:2022:DMR

[Jon22] M. C. Jones. Duals of multiplicative relationships involving beta and gamma random variables. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001821.

Joutard:2022:SAP

[Jou22] Cyrille Joutard. Saddlepoint approximations for the probability mass functions of some nonparametric test statistics. Statistics & Probability Letters, 188(??):??, September 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001079.

Jaworski:2020:NCV

[JP20] Piotr Jaworski and Marcin Pitera. A note on conditional variance and characterization of probability distributions. Statistics & Probability Letters, 163(??):??, August 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301036.

Jiang:2020:BLU

[JQS20] Hong Jiang, Jianwei Qian, and Yuqin Sun. Best linear unbiased predictors and estimators under a pair of constrained seemingly unrelated regression models. Statistics & Probability Letters, 158 (??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303153.

Jia:2023:CPT

[JQWD23] Kai Jia, Renjun Qiu, Zhengming Wang, and Xiaojun Duan. On the compression phenomenon of typical discrepancies. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000469.

Justice:2020:TBA

[JS20] Sam Justice and N. D. Shyamalkumar. On a tail bound for analyzing random trees. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300493.

Janssen:2020:NBK

[JSV20] Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. A note on the behaviour of a kernel-smoothed kernel density estimator.

Statistics & Probability Letters, 158(??):??, March 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303098.

Ji:2022:CSE

[JSWZ22] Ronglin Ji, Xuejun Shi, Shijie Wang, and Jinming Zhou. Convexity and sublinearity of g-expectations. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001249.

Jurek:2021:LTN

[Jur21] Zbigniew J. Jurek. On a limit theorem for a non-linear scaling. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001255.

Jurek:2023:WUC

[Jur23] Zbigniew J. Jurek. Which Urbanik class L_{k}, do the hyperbolic and the generalized logistic characteristic functions belong to? Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522300038X.

Jiang:2024:BME

[JW24] Jie Jiang and Lichun Wang. Bayes minimax estimator of the mean vector in an elliptically contoured distribution. Statistics & Probability Letters, 213(??):??, October 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400155X.

Jiang:2021:SOA

[JWW21] Bochuan Jiang, Zuzheng Wang, and Yaping Wang. Strong orthogonal arrays of strength two-plus based on the Addelman– Kempthorne method. Statistics & Probability Letters, 175(??): ??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000766.

Ji:2022:TBJ

[JX22] Huijie Ji and Fubao Xi. The tail behavior of jump-diffusion Cox–Ingersoll–Ross processes with regime-switching. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002339.

Jacobovic:2023:PTP

[JZ23] Royi Jacobovic and Or Zuk. A phase transition for the probability of being a maximum among random vectors with general iid coordinates. Statistics & Probability Letters, 199(??): ??, August 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000718.

Kabaila:2024:YHC

[Kab24] Paul Kabaila. On Yu and Hoff’s confidence intervals for treatment means. Statistics & Probability Letters, 213(??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001391.

Kapelko:2022:MAD

[Kap22] Rafał Kapelko. On the moment absolute deviation of order statistics from uniform distribution. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002406.

Kargol:2024:LCS

[Kar24] Alina Kargol. On the local compactness of spaces of positive measures. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000609.

Kawai:2021:GAS

[Kaw21] Reiichiro Kawai. A general approach to sample path generation of infinitely divisible processes via shot noise representation. Statistics & Probability Letters, 174(??):??, July 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000535.

Kawakami:2023:DCN

[Kaw23] Hajime Kawakami. Doob’s consistency of a non-Bayesian updating process. Statistics & Probability Letters, 203(??):??, December 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001451.

Kayid:2024:FRI

[Kay24] Mohamed Kayid. Further results involving residual and past extropy with their applications. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001706.

Kharazmi:2021:JIG

[KB21] Omid Kharazmi and Narayanaswamy Balakrishnan. Jenseninformation generating function and its connections to some wellknown information measures. Statistics & Probability Letters, 170 (??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302984.

Khan:2021:SPM

[KBM21] Ruhul Ali Khan, Dhrubasish Bhattacharyya, and Murari Mitra. On some properties of the mean inactivity time function. Statistics & Probability Letters, 170(??):??, March 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302960.

Koeneman:2022:IAT

[KC22] Scott H. Koeneman and Joseph E. Cavanaugh. An improved asymptotic test for the Jaccard similarity index for binary data. Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000074.

Kiesse:2024:DSO

[KD24] Tristan Senga Kiessé and Gilles Durrieu. On a discrete symmetric optimal associated kernel for estimating count data distributions. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000476.

Kattumannil:2021:NPE

[KDN21] Sudheesh K. Kattumannil, Isha Dewan, and Sreelaksmi N. Nonparametric estimation of Gini index with right censored observations. Statistics & Probability Letters, 175(??):??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000754.

Kondratiev:2023:CPP

[KdS23] Yuri Kondratiev and José L. da Silva. Compound Poisson processes: Potentials, green measures and random times. Statistics & Probability Letters, 197(??):??, June 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000391.

Kella:2024:ITC

[Kel24] Offer Kella. On independence of time and cause. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001682.

Kengne:2021:SCM

[Ken21] William Kengne. Strongly consistent model selection for general causal time series. Statistics & Probability Letters, 171(??): ??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220303035.

Kevei:2024:BPI

[Kev24] Péter Kevei. Branching processes with immigration in a random environment — the Grincevičius–Grey setup. Statistics & Probability Letters, 214(??):??, November 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001688.

Khartov:2022:CQI

[Kha22] A. A. Khartov. A criterion of quasi-infinite divisibility for discrete laws. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000426.

Kyong-Il:2024:EUS

[KIMG24] Ri Kyong-Il and Sin Myong-Guk. Existence and uniqueness of solution for fully coupled fractional forward–backward stochastic differential equations with delay and anticipated term. Statistics & Probability Letters, 206(??):??, March 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001785.

Kim:2024:IRE

[KjL24] Bora Kim and Myoung jae Lee. Instrument-residual estimator for multi-valued instruments under full monotonicity. Statistics & Probability Letters, 213(??):??, October 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001561.

Kerger:2020:PEO

[KK20] Phillip Kerger and Kei Kobayashi. Parameter estimation for one-sided heavy-tailed distributions. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220301115.

Kataria:2022:EEE

[KK22a] K. K. Kataria and M. Khandakar. Extended eigenvalue– eigenvector method. Statistics & Probability Letters, 183(??): ??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221003011.

Kim:2022:IPD

[KK22b] Bara Kim and Jeongsim Kim. Identification of parameters from the distribution of the maximum or minimum of Poisson random variables. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002054.

Kumar:2024:NBA

[KK24] Amit N. Kumar and Poleen Kumar. A negative binomial approximation to the distribution of the sum of maxima of indicator random variables. Statistics & Probability Letters, 208 (??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000099.

Kim:2025:ESF

[KK25] Bara Kim and Jeongsim Kim. Extension of a strong form of the three-dimensional Gaussian product inequality. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002451.

Kasinathan:2024:APH

[KKC^{+}24] Ramkumar Kasinathan, Ravikumar Kasinathan, Dimplekumar Chalishajar, Dumitru Baleanu, and Varshini Sandrasekaran. The averaging principle of Hilfer fractional stochastic pantograph equations with non-Lipschitz conditions. Statistics & Probability Letters, 215(??):??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001901.

Konzou:2020:RCG

[KKG20] Essomanda Konzou, Efoévi Koudou, and Kossi E. Gneyou. Rate of convergence of generalized inverse Gaussian and Kummer distributions to the gamma distribution via Stein’s method. Statistics & Probability Letters, 159(??):??, April 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303293.

Kim:2024:RSQ

[KKH24] Kon-Gun Kim, Mun-Chol Kim, and Ho-Jin Hwang. Representation of solutions to quadratic 2BSDEs with unbounded terminal values. Statistics & Probability Letters, 213(??):??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001603.

Kang:2021:NSB

[KKI21] Taegyu Kang, Young Min Kim, and Jongho Im. A note on stationary bootstrap variance estimator under long-range dependence. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302741.

Kim:2022:LEU

[KKO22] Kon-Gun Kim, Mun-Chol Kim, and Hun O. Local existence and uniqueness of solutions to quadratic BSDEs with weak monotonicity and general growth generators. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000785.

Kim:2024:PFB

[KKS24] Kyong-Hui Kim, Jong-Kuk Kim, and Myong Guk Sin. Pricing formula for a barrier call option based on stochastic delay differential equation. Statistics & Probability Letters, 205(??): ??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001670.

Kreer:2024:WDW

[KKT24] Markus Kreer, Ayse Kizilersu, and Anthony W. Thomas. When is the discrete Weibull distribution infinitely divisible? Statistics & Probability Letters, 215(??):??, December 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002074.

Kloodt:2021:IFN

[Klo21] Nick Kloodt. Identification in a fully nonparametric transformation model with heteroscedasticity. Statistics & Probability Letters, 170(??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220303217.

Kim:2020:SCA

[KM20] Daehong Kim and Masakuni Matsuura. Semi-classical asymptotics for scattering length of symmetric stable processes. Statistics & Probability Letters, 167(??):??, December 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302248.

Kolev:2022:NCB

[KM22] Nikolai Kolev and Sabrina Mulinacci. New characterizations of bivariate discrete Schur-constant models. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001954.

Kozubowski:2025:WTR

[KMP25] Tomasz J. Kozubowski, Dorota Młynarczyk, and Anna K. Panorska. Waiting time representation of discrete distributions. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400244X.

Kondratiev:2021:ABF

[KMS21] Yuri Kondratiev, Yuliya Mishura, and René L. Schilling. Asymptotic behaviour and functional limit theorems for a time changed Wiener process. Statistics & Probability Letters, 170(??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030300X.

Kleyntssens:2022:BMM

[KN22] T. Kleyntssens and S. Nicolay. From the Brownian motion to a multifractal process using the Lévy–Ciesielski construction. Statistics & Probability Letters, 186(??):??, July 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000529.

Kaminski:2021:SLN

[KO21a] Łukasz Kamiński and Adam Osȩkowski. Sharp Lorentz-norm estimates for BMO martingales. Statistics & Probability Letters, 173 (??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000304.

Kim:2021:GCT

[KO21b] Mun-Chol Kim and Hun O. A general comparison theorem for reflected BSDEs. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000201.

Koperberg:2024:CMC

[Kop24] Twan Koperberg. Couplings and matchings: Combinatorial notes on Strassen’s theorem. Statistics & Probability Letters, 209(??): ??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000580.

Ki:2021:IHC

[KP21] Dohyeong Ki and Byeong U. Park. Intrinsic Hölder classes of density functions on Riemannian manifolds and lower bounds to convergence rates. Statistics & Probability Letters, 169(??): ??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302625.

Kattumannil:2022:NPE

[KP22a] Sudheesh K. Kattumannil and Sreedevi E. P. Non-parametric estimation of cumulative (residual) extropy. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222000414.

Kim:2022:NAW

[KP22b] Yoon Tae Kim and Hyun Suk Park. Normal approximation when a chaos grade is greater than two. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000141.

Kundu:2023:BEC

[KP23] Anupam Kundu and Mohsen Pourahmadi. Bayesian estimation of constrained mean-covariance of normal distributions.

Statistics & Probability Letters, 194(??):??, March 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002589.

Kock:2024:RMD

[KP24] Anders Bredahl Kock and David Preinerstorfer. A remark on moment-dependent phase transitions in high-dimensional Gaussian approximations. Statistics & Probability Letters, 211(??): ??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001184.

Kamps:2023:RVP

[KR23] Udo Kamps and Diana Rauwolf. A record-values property of a renewal process with random inspection time. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000093.

Klein:2024:EAR

[KR24] Thierry Klein and Paul Rochet. Efficiency of the averaged rankbased estimator for first order Sobol index inference. Statistics & Probability Letters, 207(??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002389.

Kreer:2022:EPD

[Kre22] Markus Kreer. An elementary proof for dynamical scaling for certain fractional non-homogeneous Poisson processes. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002583.

Kriukov:2022:UBR

[Kri22] Nikolai Kriukov. Uniform bounds for ruin probability in multidimensional risk model. Statistics & Probability Letters, 190(??): ??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001559.

Kongjiw:2023:LLT

[KRN23] Hattacha Kongjiw, Petcharat Rattanawong, and Kritsana Neammanee. Local limit theorems for collective risk models. Statistics & Probability Letters, 201(??):??, October 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000913.

Kroll:2022:UCH

[Kro22] Martin Kroll. On the universal consistency of histograms anonymised by a randomised response technique. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000347.

Krystecki:2022:PRP

[Kry22] Konrad Krystecki. Parisian ruin probability for two-dimensional Brownian risk model. Statistics & Probability Letters, 182(??): ??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002819.

Kovac:2020:CEM

[KS20a] Vjekoslav Kovač and Mario Stipčić. Convergence of ergodicmartingale paraproducts. Statistics & Probability Letters, 164(??): ??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301292.

Kraus:2020:RRP

[KS20b] David Kraus and Marco Stefanucci. Ridge reconstruction of partially observed functional data is asymptotically optimal. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301164.

Kremer:2020:AAR

[KS20c] D. Kremer and H.-P. Scheffler. About atomless random measures on δ-rings. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301085.

Kojevnikov:2022:BEB

[KS22] Denis Kojevnikov and Kyungchul Song. A Berry–Esseen bound for vector-valued martingales. Statistics & Probability Letters, 186 (??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000517.

Karimi:2024:SOS

[KSAA24] Nader Karimi, Erfan Salavati, Hirbod Assa, and Hojatollah Adibi. A stochastic optimal stopping model for storable commodity prices. Statistics & Probability Letters, 204(??):??, January 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001657.

Kaiser:2024:RRW

[KSH24] Robin Kaiser and Ecaterina Sava-Huss. Random rotor walks and i.i.d. sandpiles on Sierpiński graphs. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224000592.

Karamzadeh:2020:CSR

[KSMF20] M. Karamzadeh, A. R. Soltani, and H. A. Mardani-Fard. On a class of spatial renewal processes: Renewal processes synchronization probabilities. Statistics & Probability Letters, 158(??): ??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303049.

Kadankova:2020:SNM

[KSW20] Tetyana Kadankova, Thomas Simon, and Min Wang. On some new moments of Gamma type. Statistics & Probability Letters, 165 (??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301577.

Kokoszka:2022:LTB

[KSW22] Piotr Kokoszka, Deepak Singh, and Haonan Wang. Long term behavior of incomplete and time varying product ratings.

Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200013X.

Kume:2025:RCA

[KSW25] Alfred Kume, Tomonari Sei, and Andrew T. A. Wood. On the representation and computational aspects of the distribution of a linear combination of independent noncentral chi-squared random variables. Statistics & Probability Letters, 218(??):??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002608.

Kolliopoulos:2024:LPA

[KSX24] Nikolaos Kolliopoulos, David Sanchez, and Amy Xiao. Largepopulation asymptotics for the maximum of diffusive particles with mean-field interaction in the noises. Statistics & Probability Letters, 212(??):??, September 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001196.

Kubilius:2020:CQV

[Kub20] K. Kubilius. CLT for quadratic variation of Gaussian processes and its application to the estimation of the Orey index. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301486.

Kumar:2022:SFH

[Kum22] Vivek Kumar. Stochastic fractional heat equation perturbed by general Gaussian and non-Gaussian noise. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222000104.

Kutlu:2021:DRQ

[Kut21] Merve Kutlu. On a denseness result for quasi-infinitely divisible distributions. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001012.

Kume:2024:OSR

[KVW24] Alfred Kume, Cristiano Villa, and Stephen G. Walker. Optimal square-root pooling from expert opinions. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223002377.

Kevei:2021:MSD

[KW21] Péter Kevei and Péter Wiandt. Moments of the stationary distribution of subcritical multitype Galton–Watson processes with immigration. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000298.

Katselis:2021:CIV

[KXBS21] Dimitrios Katselis, Xiaotian Xie, Carolyn L. Beck, and R. Srikant. On concentration inequalities for vector-valued Lipschitz functions. Statistics & Probability Letters, 173(??):??, June 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522100033X.

Khardani:2022:NRR

[KY22] Salah Khardani and Anne Françoise Yao. Nonparametric recursive regression estimation on Riemannian Manifolds. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002364.

Korolev:2021:BCR

[KZ21] Victor Korolev and Alexander Zeifman. Bounds for convergence rate in laws of large numbers for mixed Poisson random sums. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302212.

Khaledi:2023:SSG

[KZB23] Majid Jafari Khaledi, Hamid Zareifard, and Hossein Boojari. A spatial skew-Gaussian process with a specified covariance function. Statistics & Probability Letters, 192(??):??, January 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001948.

Lando:2021:TIL

[Lan21] Tommaso Lando. A test for the increasing log-odds rate family. Statistics & Probability Letters, 170(??):??, March 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303205.

Lando:2023:TDI

[Lan23] Tommaso Lando. Testing departures from the increasing hazard rate property. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002498.

Lando:2024:NIN

[Lan24] Tommaso Lando. Nonparametric inference for NBUE distributions based on the TTT transform. Statistics & Probability Letters, 212(??):??, September 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001263.

Lardjane:2022:SUC

[Lar22] Salim Lardjane. Strong uniform consistency of the frequency polygon density estimator for stable non-anticipative stochastic processes. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522200150X.

Lassance:2023:ASE

[Las23] Nathan Lassance. An analytical shrinkage estimator for linear regression. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002735.

Lauritzen:2024:TVC

[Lau24] Steffen Lauritzen. Total variation convergence preserves conditional independence. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522400169X.

Lando:2020:SOS

[LBB20] Tommaso Lando and Lucio Bertoli-Barsotti. Second-order stochastic dominance for decomposable multiparametric families with applications to order statistics. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219303372.

Liu:2022:RFC

[LBL^{+}22] Yan Liu, Zhidong Bai, Hua Li, Jiang Hu, Zhihui Lv, and Shurong Zheng. RDS free CLT for spiked eigenvalues of high-dimensional covariance matrices. Statistics & Probability Letters, 187(??): ??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000827.

Li:2020:LDS

[LBZ20] Rong Li, Xiuchun Bi, and Shuguang Zhang. Large deviations for sums of claims in a general renewal risk model with the regression dependent structure. Statistics & Probability Letters, 165(??): ??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301607.

Luo:2020:MBT

[LBZH20] Wenya Luo, Zhidong Bai, Shurong Zheng, and Yongchang Hui. A modified BDS test. Statistics & Probability Letters, 164(??): ??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300973.

Li:2024:UAF

[LCCZ24] Mingjun Li, Zhangting Chen, Dongya Cheng, and Junyi Zhou. Uniform asymptotics for finite-time ruin probabilities of a bidimensional compound risk model with stochastic returns. Statistics & Probability Letters, 207(??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002365.

Liu:2021:ATD

[LCF21] Yang Liu, Zhenlong Chen, and Ke-Ang Fu. Asymptotics for a time-dependent renewal risk model with subexponential main claims and delayed claims. Statistics & Probability Letters, 177 (??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522100136X.

Lau:2020:VIM

[LCH20] John W. Lau, Edward Cripps, and Wendy Hui. Variational inference for multiplicative intensity models. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300237.

Liu:2020:LLN

[LD20] Qun Liu and Zhishan Dong. Limit laws for the number of triangles in the generalized random graphs with random node weights. Statistics & Probability Letters, 161(??):??, June 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300365.

Lemyre:2021:NRE

[LD21a] Felix Camirand Lemyre and Geoffrey Decrouez. Nonparametric recursive estimation of the copula. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302327.

Liang:2021:ELB

[LD21b] Wei Liang and Hongsheng Dai. Empirical likelihood based on synthetic right censored data. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302650.

Lu:2022:SCE

[LDM22] Tianshi Lu, Juan Du, and Chunsheng Ma. Stochastic comparison for elliptically contoured random fields. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001407.

Le:2022:ECS

[Le22] V. Le. On the extinction of continuous state branching processes with competition. Statistics & Probability Letters, 185 (??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000293.

Lee:2020:WEO

[Lee20] Jeonghwa Lee. Wavelet estimation in OFBM: Choosing scale parameter in different sampling methods and different parameter values. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301802.

Lee:2021:HEO

[Lee21a] Jeonghwa Lee. Hurst estimation for operator scaling random fields. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001504.

Leeb:2021:NIC

[Lee21b] William Leeb. A note on identifiability conditions in confirmatory factor analysis. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001528.

Leeb:2022:OSV

[Lee22] William Leeb. Optimal singular value shrinkage for operator norm loss: Extending to non-square matrices. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000633.

Lee:2023:ESF

[Lee23] Taeho Lee. Exact simulation for the first hitting time of Brownian motion and Brownian bridge. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222001754.

Lefebvre:2024:ESF

[Lef24] Mario Lefebvre. Exact solution to a first-passage problem for an Ornstein–Uhlenbeck process with jumps and its integral. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001803.

Leuenberger:2022:BSA

[LEJT22] Michael Leuenberger, Esther Eustache, Raphaël Jauslin, and Yves Tillé. Balancing a sample almost perfectly. Statistics & Probability Letters, 180(??):??, January 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001917.

Levy:2021:DSI

[Lev21] Edmond Levy. On the density for sums of independent Mittag-Leffler variates with common order. Statistics & Probability Letters, 179(??):??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001735.

Liu:2024:OMA

[LG24] Shanqi Liu and Hongjun Gao. The Onsager–Machlup action functional for degenerate stochastic differential equations in a class of norms. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522300233X.

Li:2023:MVD

[LGHG23] Yan Li, Zhigen Gao, Wei Huang, and Jianhua Guo. Matrixvariate data analysis by two-way factor model with replicated observations. Statistics & Probability Letters, 202(??):??, November 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001281.

Long:2022:NDN

[LH22] Yuhang Long and Tao Huang. A note on a dynamic network model with homogeneous structure. Statistics & Probability Letters, 184 (??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000013.

Lin:2025:PCL

[LH25] Yong Lin and Mian Huang. Penalized composite likelihood estimation for hidden Markov models with unknown number of states. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002165.

Li:2024:VCF

[LHW24] Min Li, Chengming Huang, and Nan Wang. A variation of constant formula for Caputo–Hadamard fractional stochastic differential equations ^{⋆}. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001858.

Li:2023:SHE

[Li23] Xiang Li. Stochastic heat equation with Ornstein–Uhlenbeck operator. Statistics & Probability Letters, 201(??):??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001116.

Li:2024:IMB

[Li24a] Hsin-Lun Li. An imitation model based on the majority. Statistics & Probability Letters, 206(??):??, March 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002316.

Li:2024:NME

[Li24b] Xiao Li. Nearly minimax empirical Bayesian prediction of independent Poisson observables. Statistics & Probability Letters, 208 (??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000440.

Liao:2020:PRB

[Lia20] Chen-Tuo Liao. Partially replicated block designs for two-level factorial experiments. Statistics & Probability Letters, 158(??): ??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302998.

Light:2023:IHI

[Lig23] Bar Light. Improving Hoeffding’s inequality using higher moments information. Statistics & Probability Letters, 201(??):??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001062.

Lin:2024:NOS

[Lin24] Yi-Shen Lin. A note on one-sided solutions for optimal stopping problems driven by Lévy processes. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223002134.

Liu:2020:SRR

[Liu20a] Kai Liu. Some regularity results on stochastic convolutions in point delay differential equations perturbed by noise. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302706.

Liu:2020:GTA

[Liu20b] Yang Liu. A general treatment of alternative expectation formulae. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301668.

Liu:2024:SLC

[Liu24] Jiawei Liu. A scaling limit of controlled branching processes. Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000506.

Li:2020:CSL

[LKA20] Zhiming Li, Qingxun Kong, and Mingyao Ai. Construction of some s-level regular designs with general minimum lower-order confounding. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302005. See correction [ZCL22].

Li:2020:PWS

[LL20a] Chen Li and Xiaohu Li. Preservation of weak SAI’s under increasing transformations with applications. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301310.

Li:2020:TQT

[LL20b] Ruinan Li and Yumeng Li. Talagrand’s quadratic transportation cost inequalities for reflected SPDEs driven by space–time white noise. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300377.

Laib:2021:AOT

[LL21a] Naâmane Laïb and Tewfik Lounis. Asymptotically optimal tests for non-linear autoregressive model with β-ARCH errors. Statistics & Probability Letters, 178(??):??, November 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001462.

Li:2021:SDR

[LL21b] Chen Li and Xiaohu Li. On stochastic dependence in residual lifetime and inactivity time with some applications. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000821.

Li:2023:SHC

[LL23a] Chen Li and Xiaohu Li. On k-out-of-n systems with homogeneous components and one independent cold standby redundancy. Statistics & Probability Letters, 203(??):??, December 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001426.

Lu:2023:GWP

[LL23b] Dawei Lu and Huangding Lv. A Galton–Watson process with a threshold at 1 and an immigration at 0. Statistics & Probability Letters, 201(??):??, October 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001050.

Lavigne:2024:QUP

[LL24a] Aurore Lavigne and Silvia Liverani. Quantifying the uncertainty of partitions for infinite mixture models. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001542.

Li:2024:FAS

[LL24b] Ge Li and Jicheng Liu. Fluctuation analysis of synchronized system. Statistics & Probability Letters, 206(??):??, March 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522300216X.

Lifshts:2024:IZR

[LL24c] M. A. Lifshts and I. M. Lialinov. Intersections of Zipf random sets: Maximal weighted relevance. Statistics & Probability Letters, 208 (??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000191.

Liu:2024:ASP

[LL24d] Shuning Liu and Guangying Lv. Almost sure polynomial stability and stabilization of stochastic differential systems with impulsive effects. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002043.

Li:2025:CWE

[LLL^{+}25] Junpeng Li, Guanghui Li, Wei Leng, Chongqi Zhang, and Hongyu Su. Construction of weighted efficiency optimal designs for experiments with mixtures. Statistics & Probability Letters, 216(??): ??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002244.

Lu:2020:SRI

[LLM20] Tianshi Lu, Nikolai Leonenko, and Chunsheng Ma. Series representations of isotropic vector random fields on balls. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302299.

Levin:2024:LCM

[LLM24] Bruce Levin and Erik Learned-Miller. Log-concavity of multinomial likelihood functions under interval censoring constraints on frequencies or their partial sums. Statistics & Probability Letters, 207(??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224000087.

Litimein:2023:LLE

[LLMB23] Ouahiba Litimein, Ali Laksaci, Boubaker Mechab, and Salim Bouzebda. Local linear estimate of the functional expectile regression. Statistics & Probability Letters, 192(??):??, January 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522200195X.

Li:2020:PCE

[LLWZ20] Min Li, Min-Qian Liu, Xiao-Lei Wang, and Yong-Dao Zhou. Prediction for computer experiments with both quantitative and qualitative factors. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301619.

Lee:2021:FSF

[LM21] Sharon X. Lee and Geoffrey J. McLachlan. On formulations of skew factor models: Skew factors and/or skew errors. Statistics & Probability Letters, 168(??):??, January 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302388.

Liu:2023:ILD

[LM23] Zhenxia Liu and Mainza Mbokoma. An improvement on the large deviations for longest runs in Markov chains. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002504.

Lanchier:2024:LBK

[LM24a] Nicolas Lanchier and Max Mercer. Limiting behavior of a kindness model. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001743.

Liang:2024:NAU

[LM24b] Wanfeng Liang and Xiaoyan Ma. A new approach for ultrahigh-dimensional covariance matrix estimation. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001530.

Lyons:2024:CCF

[LMN24] Terry J. Lyons, Vlad Margarint, and Sina Nejad. Convergence to closed-form distribution for the backward SLE_{κ} at some random times and the phase transition at κ = 8. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001827.

Li:2022:GLD

[LMS22] Deli Li, Yu Miao, and George Stoica. A general large deviation result for partial sums of i.i.d. super-heavy tailed random variables. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000050.

Li:2022:OSC

[LMSY22] Butong Li, Yongna Meng, Xiaobin Sun, and Ting Yang. Optimal strong convergence rate for a class of McKean– Vlasov SDEs with fast oscillating perturbation. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001791.

Li:2021:SGB

[LMZ21] Biao Li, Yutao Ma, and Zhengliang Zhang. On the spectral gap of Boltzmann measures on the unit sphere. Statistics & Probability Letters, 169(??):??, February 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302662.

Luati:2020:HCR

[LN20] Alessandra Luati and Marco Novelli. The Hammersley–Chapman– Robbins inequality for repeatedly monitored quantum system. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301553.

Longla:2022:DMP

[LNN22] Martial Longla, Mathias Muia Nthiani, and Fidel Djongreba Ndikwa. Dependence and mixing for perturbations of copulabased Markov chains. Statistics & Probability Letters, 180(??): ??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002017.

Liu:2022:SDE

[LNYZ22] Yujie Liu, Minwen Niu, Dacheng Yao, and Hanqin Zhang. Stationary distributions and ergodicity of reflection-type Markov chains. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001420.

Lou:2023:TDF

[Lou23] Shuwen Lou. On transition density functions of skew Brownian motions with two-valued drift. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222002255.

Lukusa:2020:NWT

[LP20] Martin T. Lukusa and Frederick Kin Hing Phoa. A note on the weighting-type estimations of the zero-inflated Poisson regression model with missing data in covariates. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303001.

Leipus:2021:THO

[LPŠ21] Remigijus Leipus, Saulius Paukštys, and Jonas Šiaulys. Tails of higher-order moments of sums with heavy-tailed increments and application to the Haezendonck–Goovaerts risk measure. Statistics & Probability Letters, 170(??):??, March 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303011.

Lautier:2024:MLE

[LPY24] Jackson P. Lautier, Vladimir Pozdnyakov, and Jun Yan. On the maximum likelihood estimation of a discrete, finite support distribution under left-truncation and competing risks. Statistics & Probability Letters, 207(??):??, April 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001979.

Liu:2022:OEL

[LQ22] Youming Liu and Xinyu Qi. Optimal estimation for lower bound of the packing number. Statistics & Probability Letters, 186 (??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000748.

Lv:2024:PLI

[LQZ24] Yang Lv, Guoyou Qin, and Zhongyi Zhu. Population-level information for improving quantile regression efficiency. Statistics & Probability Letters, 215(??):??, December 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001962.

Li:2022:PBW

[LR22] W. Li and F. J. Rubio. On a prior based on the Wasserstein information matrix. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001705.

Levine:2020:IPT

[LRS20] Michael Levine, Donald Richards, and Jianxi Su. Independence properties of the truncated multivariate elliptical distributions. Statistics & Probability Letters, 161(??):??, June 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300328.

Liu:2024:CTC

[LRW24] Rong-Li Liu, Yan-Xia Ren, and Yingrui Wang. Coalescence times for critical Galton–Watson processes with immigration. Statistics & Probability Letters, 210(??):??, July 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000907.

Lei:2022:RRC

[LS22] Xiaoyu Lei and Qi-Man Shao. A refined randomized concentration inequality. Statistics & Probability Letters, 187(??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000888.

Li:2024:CEE

[LSFZ24] Li Li, Pengfei Shi, Qingliang Fan, and Wei Zhong. Causal effect estimation with censored outcome and covariate selection. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001578.

Lhaut:2022:UCB

[LSS22] Stéphane Lhaut, Anne Sabourin, and Johan Segers. Uniform concentration bounds for frequencies of rare events. Statistics & Probability Letters, 189(??):??, October 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001493.

Li:2022:NAS

[LT22a] Yanpeng Li and Boping Tian. Non-asymptotic sub-Gaussian error bounds for hypothesis testing. Statistics & Probability Letters, 189 (??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001377.

Lifshits:2022:MRA

[LT22b] M. A. Lifshits and A. A. Tadevosian. On the maximum of random assignment process. Statistics & Probability Letters, 187(??): ??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001006.

Liu:2022:PPE

[LT22c] Huiyan Liu and Zhongquan Tan. Point processes of exceedances by Gaussian random fields with applications to asymptotic locations of extreme order statistics. Statistics & Probability Letters, 189 (??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001419.

Li:2023:ONA

[LT23] Yanpeng Li and Boping Tian. Optimal non-asymptotic concentration of centered empirical relative entropy in the highdimensional regime. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000275.

Li:2024:SGM

[LT24] Yanpeng Li and Boping Tian. On the sub-Gaussianity of the missing mass. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002158.

Li:2022:RPE

[LTZ22] Kangqiang Li, Songqiao Tang, and Lixin Zhang. Robust parameter estimation of regression models under weakened moment assumptions. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001912.

Lu:2022:AMC

[Lu22] Yingdong Lu. Approximating the magnetization in the Curie– Weiss model. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001547.

Lu:2024:IRT

[Lu24] Tianshi Lu. Isotropic random tangential vector fields on spheres. Statistics & Probability Letters, 213(??):??, October 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400141X.

Luczynska:2021:UEF

[Łuc21] Gabriela Łuczyńska. Unique ergodicity for function systems on the circle. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000468.

Luo:2020:RBT

[Luo20] Peng Luo. Reflected BSDEs with time-delayed generators and nonlinear resistance. Statistics & Probability Letters, 163(??): ??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300687.

Lv:2022:SAS

[Lv22] Xiang Lv. Stability analysis of semilinear stochastic differential equations. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002194.

Levental:2023:FDO

[LV23] S. Levental and P. Vellaisamy. Formulas for the divergence operator in isonormal Gaussian space. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222002565.

Lv:2020:BSS

[LW20] Guangying Lv and Jinlong Wei. Blowup solutions for stochastic parabolic equations. Statistics & Probability Letters, 166(??): ??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301796.

Li:2024:SBC

[LW24a] Huimin Li and Jinru Wang. Sparse basis covariance matrix estimation for high dimensional compositional data via hard thresholding. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000579.

Li:2024:DPA

[LW24b] Libo Li and Zhuoshu Wu. Defaultable perpetual American put option in a last passage time model. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002419.

Li:2024:OPU

[LW24c] Qian Li and Li Wang. Option pricing under jump diffusion model. Statistics & Probability Letters, 211(??):??, August 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001068.

Li:2024:TTI

[LW24d] Ruinan Li and Xinyu Wang. Talagrand’s transportation inequality for SPDEs with locally monotone drifts. Statistics & Probability Letters, 204(??):??, January 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001694.

Liang:2022:RSP

[LWM22] Wanfeng Liang, Yue Wu, and Xiaoyan Ma. Robust sparse precision matrix estimation for high-dimensional compositional data. Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000098.

Li:2021:OTS

[LWP21] Yingqiu Li, Yushao Wei, and Zhaohui Peng. Occupation times for spectrally negative Lévy processes on the last exit time. Statistics & Probability Letters, 175(??):??, August 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000730.

Luo:2021:IBR

[LWX21] Guowang Luo, Mixia Wu, and Liwen Xu. IPW-based robust estimation of the SAR model with missing data. Statistics & Probability Letters, 172(??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000274.

Lai:2025:SSE

[LWZ25] Peng Lai, Zhou Wang, and Yurong Zhang. Semi-supervised estimation of a single-index varying-coefficient model. Statistics & Probability Letters, 218(??):??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002815.

Li:2024:SEM

[LX24a] Shuang Li and Jie Xiong. SIR epidemic model with non-Lipschitz stochastic perturbations. Statistics & Probability Letters, 210 (??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000786.

Lo:2024:RNA

[LX24b] Tiffany Y. Y. Lo and Aihua Xia. On the rate of normal approximation for Poisson continuum percolation. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000798.

Lu:2020:GAJ

[LXC20] Fei Lu, Liugen Xue, and Xiong Cai. GEE analysis in joint mean-covariance model for longitudinal data. Statistics & Probability Letters, 160(??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300080.

Li:2023:SPG

[LXH23] Fu-Bo Li, Kun Xu, and Ze-Chun Hu. A study on the Poisson, geometric and Pascal distributions motivated by Chvátal’s conjecture. Statistics & Probability Letters, 200(??):??, September 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000950.

Lan:2020:MAS

[LXZ20] Guangqiang Lan, Fang Xia, and Mei Zhao. p-th moment (p ∈ (0, 1)) and almost sure exponential stability of the exact solutions and modified truncated EM method for stochastic differential equations. Statistics & Probability Letters, 160(??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300043.

Liu:2020:GCO

[LYC20] Xin Liu, Rong-Xian Yue, and Kashinath Chatterjee. Geometric characterization of D-optimal designs for random coefficient regression models. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219303426.

Li:2022:CSF

[LYL22] Hui Li, Liuqing Yang, and Min-Qian Liu. Construction of space-filling orthogonal Latin hypercube designs. Statistics & Probability Letters, 180(??):??, January 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002078.

Lyu:2020:PHM

[Lyu20] Yangyang Lyu. Precise high moment asymptotics for parabolic Anderson model with log-correlated Gaussian field. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303086.

Liu:2021:ODC

[LYY21] Xin Liu, Min Ye, and Rong-Xian Yue. Optimal designs for comparing population curves in hierarchical models. Statistics & Probability Letters, 178(??):??, November 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001541.

Li:2020:GLK

[LZ20] Siran Li and Xunjie Zheng. A generalization of Lemma 1 in Kotlarski (1967). Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301176.

Li:2021:ASC

[LZ21] Jingyu Li and Yong Zhang. An almost sure central limit theorem for the stochastic heat equation. Statistics & Probability Letters, 177(??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221001115.

Li:2022:SPS

[LZ22a] Zenghu Li and Yaping Zhu. Survival probability for super-Brownian motion with absorption. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000566.

Lu:2022:FET

[LZ22b] Dawei Lu and Yinbing Zhou. The first exit time of fractional Brownian motion from the minimum and maximum parabolic domains. Statistics & Probability Letters, 186(??):??, July 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000578.

Liu:2023:SES

[LZ23a] Yu Liu and Xiaoyang Zhuang. Shrinkage estimation of semiparametric spatial autoregressive panel data model with fixed effects. Statistics & Probability Letters, 194(??):??, March 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002590.

Liu:2023:UBI

[LZ23b] Zhiguang Liu and Quanxin Zhu. Ultimate boundedness of impulsive stochastic delay differential equations with delayed impulses. Statistics & Probability Letters, 199(??):??, August 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000810.

Li:2024:TIR

[LZ24a] Ruinan Li and Beibei Zhang. A transportation inequality for reflected SPDEs on infinite spatial domain. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001839.

Li:2024:EMS

[LZ24b] Yanfang Li and Guohuan Zhao. Euler–Maruyama scheme for SDE driven by Lévy process with Hölder drift. Statistics & Probability Letters, 215(??):??, December 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001895.

Lu:2025:OWM

[LZ25] Minyuan Lu and Bu Zhou. A one-way MANOVA test for high-dimensional data using clustering subspaces. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002621.

Li:2023:ELI

[LZH23] Yuling Li, Huanchao Zhou, and Jiang Hu. The eigenvector LSD of information plus noise matrices and its application to linear regression model. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000354.

Li:2025:ECR

[LZLX25] Yingqiu Li, Xin Zhang, Zhan Lu, and Sheng Xiao. Exact convergence rate of the central limit theorem and polynomial convergence rate for branching processes in a random environment. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002372.

Liu:2020:TRC

[LZMZ20] Yan Liu, Sanguo Zhang, Shuangge Ma, and Qingzhao Zhang. Tests for regression coefficients in high dimensional partially linear models. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300754.

Li:2024:BAL

[LZZ24a] Fanqun Li, Mingtao Zhao, and Kongsheng Zhang. Bayesian adaptive lasso estimation of large graphical model based on modified Cholesky decomposition. Statistics & Probability Letters, 206(??): ??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002286.

Liang:2024:BSA

[LZZ24b] Fei Liang, Yidi Zhang, and Shuangshuang Zhao. Blow-up solution to an abstract non-local stochastic heat equation with Lévy noise. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002006.

Ma:2021:HCR

[Ma21] Chunsheng Ma. Hyperbolic cosine ratio and hyperbolic sine ratio random fields. Statistics & Probability Letters, 179(??):??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001747.

Matula:2022:NEL

[MA22] Przemysław Matuła and André Adler. A note on exact laws of large numbers for the range of a sample from Pareto-type distributions. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002595.

Ma:2023:MEP

[Ma23] Chunsheng Ma. Multivariate exponential power Lévy processes and random fields. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000305.

Maheshwari:2023:TSF

[Mah23] Aditya Maheshwari. Tempered space fractional negative binomial process. Statistics & Probability Letters, 196(??):??, May 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000238.

Mao:2022:IOD

[Mao22] Lu Mao. Identification of the outcome distribution and sensitivity analysis under weak confounder–instrument interaction. Statistics & Probability Letters, 189(??):??, October 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001390.

Marie:2020:NET

[Mar20a] Nicolas Marie. Nonparametric estimation of the trend in reflected fractional SDE. Statistics & Probability Letters, 158(??): ??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303050.

Marzougue:2020:NOS

[Mar20b] Mohamed Marzougue. A note on optional Snell envelopes and reflected backward SDEs. Statistics & Probability Letters, 165(??): ??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522030136X.

Martsynyuk:2021:NDC

[Mar21a] Yuliya V. Martsynyuk. Nonparametric detection of change in the slope and intercept in linear structural errors-in-variables models. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302601.

Marzougue:2021:MLT

[Mar21b] Mohamed Marzougue. Monotonic limit theorem for BSDEs with regulated trajectories. Statistics & Probability Letters, 176(??): ??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001139.

Marie:2022:PES

[Mar22] Nicolas Marie. Projection estimators of the stationary density of a differential equation driven by the fractional Brownian motion. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002066.

Marzougue:2023:NLD

[Mar23] Mohamed Marzougue. Non-linear Dynkin games over split stopping times. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002346.

Maurya:2024:LSD

[Mau24] Shambhu Nath Maurya. Limiting spectral distribution of Toeplitz and Hankel matrices with dependent entries. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000610.

Mayer:2023:TSE

[May23] Alexander Mayer. Two-step estimation in linear regressions with adaptive learning. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002747.

Maatouk:2022:NSH

[MBR22] Hassan Maatouk, Xavier Bay, and Didier Rullière. A note on simulating hyperplane-truncated multivariate normal distributions. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001730.

Maillot:2021:CDE

[MC21] Bertrand Maillot and Christophe Chesneau. On the conditional density estimation for continuous time processes with values in functional spaces. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001413.

Manna:2024:OCD

[MD24] Soumen Manna and Ashish Das. Optimal choice designs for the main effects plus specified two-factor interaction effects model for binary attributes. Statistics & Probability Letters, 209(??): ??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000555.

McFadden:2025:OCR

[MDS25] Ryan McFadden, Fraser Daly, and Seva Shneer. Optimal curing rate allocation in the SIS epidemic model. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002530.

Malekzadeh:2021:EMT

[MEA21] Ahad Malekzadeh and Asghar Esmaeli-Ayan. An exact method for testing equality of several groups in panel data models. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001188.

Merkle:2023:CPW

[Mer23] Ana Merkle. Causal predictability and weak solutions of the stochastic differential equations with driving semimartingales. Statistics & Probability Letters, 197(??):??, June 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000408.

Martins:2022:NRF

[MFF22] Ana Paula Martins, Helena Ferreira, and Marta Ferreira. A new random field on lattices. Statistics & Probability Letters, 186 (??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000669.

Marrelec:2020:CMD

[MG20] Guillaume Marrelec and Alain Giron. Cumulants of multiinformation density in the case of a multivariate normal distribution. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302330.

Mousavinasr:2020:CFD

[MGD20] S. Mousavinasr, C. R. Gonçalves, and C. C. Y. Dorea. Convergence to Fréchet distribution via Mallows distance. Statistics & Probability Letters, 163(??):??, August 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300791.

Marrelec:2021:EDP

[MGM21] Guillaume Marrelec, Alain Giron, and Laura Messio. Exponential decay of pairwise correlation in Gaussian graphical models with an equicorrelational one-dimensional connection pattern. Statistics & Probability Letters, 171(??):??, April 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303199.

Martin-Gonzalez:2024:NSR

[MGMSP24] Ehyter M. Martín-González, Antonio Murillo-Salas, and Henry Pantí. A note on series representation for the q-scale function of a class of spectrally negative Lévy processes. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000841.

Martino:2020:HMM

[MGP20] Andrea Martino, Giuseppina Guatteri, and Anna Maria Paganoni. Hidden Markov models for multivariate functional data. Statistics & Probability Letters, 167(??):??, December 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302200.

Manta:2022:ESE

[MHHJC22] Alexandra Manta, Anson T. Y. Ho, Kim P. Huynh, and David T. Jacho-Chávez. Estimating social effects in a multilayered linear-in-means model with network data. Statistics & Probability Letters, 183(??):??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002832.

Mitov:2021:CBP

[Mit21] Kosto V. Mitov. A critical branching process with immigration in varying environments. Statistics & Probability Letters, 168(??): ??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302315.

Mojirsheibani:2024:KTR

[MK24] Majid Mojirsheibani and Arin Khudaverdyan. A kernel-type regression estimator for NMAR response variables with applications to classification. Statistics & Probability Letters, 215(??): ??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002153.

Marengo:2020:UBC

[MKF20] James E. Marengo, Quinn T. Kolt, and David L. Farnsworth. An upper bound for a cyclic sum of probabilities. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301644.

Miao:2024:GLA

[ML24] Yu Miao and Deli Li. A general logarithmic asymptotic behavior for partial sums of i.i.d. random variables. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000129.

Mou:2025:PEH

[MLC25] Keyi Mou, Zhiming Li, and Jinlong Cheng. Parameter estimation and hypothesis tests in logistic model for complex correlated data. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002633.

Ma:2023:SAC

[MLZ23] Li Ma, Yujing Li, and Quanxin Zhu. Stability analysis for a class of stochastic delay nonlinear systems driven by G-Lévy Process. Statistics & Probability Letters, 195(??):??, April 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000019.

Malinovsky:2022:NDI

[MM22] Yaakov Malinovsky and John W. Moon. On negative dependence inequalities and maximal scores in round-robin tournaments. Statistics & Probability Letters, 185(??):??, June 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000402.

Mijatovic:2020:NES

[MMB20] Aleksandar Mijatović, Veno Mramor, and Gerónimo Uribe Bravo. A note on the exact simulation of spherical Brownian motion. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301395.

Makigusa:2020:ANC

[MN20] Natsumi Makigusa and Kanta Naito. Asymptotic normality of a consistent estimator of maximum mean discrepancy in Hilbert space. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302421.

Milheiro-Oliveira:2022:ASM

[MO22] Paula Milheiro-Oliveira. An alternative sequential method for the state estimation of a partially observed SETAR(1) process. Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000128.

Marchione:2023:PRH

[MO23a] Manfred Marvin Marchione and Enzo Orsingher. Pseudoprocesses related to higher-order equations of vibrations of rods. Statistics & Probability Letters, 199(??):??, August 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000767.

Matsumoto:2023:LBM

[MO23b] Hiroyuki Matsumoto and Jun Otani. Laplacian and Brownian motion on positive definite matrices, revisited. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002097.

Modarres:2023:ADM

[Mod23] Reza Modarres. Analysis of distance matrices. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002334.

Mojirsheibani:2021:NPB

[Moj21] Majid Mojirsheibani. A note on the performance of bootstrap kernel density estimation with small re-sample sizes. Statistics & Probability Letters, 178(??):??, November 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001516.

Merlevede:2020:FCN

[MP20a] Florence Merlevède and Magda Peligrad. Functional CLT for nonstationary strongly mixing processes. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219302275.

Mokkadem:2020:OEI

[MP20b] Abdelkader Mokkadem and Mariane Pelletier. Online estimation of integrated squared density derivatives. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301838.

Mokkadem:2020:REI

[MP20c] Abdelkader Mokkadem and Mariane Pelletier. Recursive estimators of integrated squared density derivatives. Statistics & Probability Letters, 157(??):??, February 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771521930255X.

Mokkadem:2021:CLI

[MP21] Abdelkader Mokkadem and Mariane Pelletier. A compact law of the iterated logarithm for online estimator of hazard rate under random censoring. Statistics & Probability Letters, 178(??): ??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001486.

Mnatsakanov:2024:RRD

[MP24] Robert M. Mnatsakanov and Denys Pommeret. On recovering the relative distribution, Part 1: the moment-recovered approach. Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400004X.

Marsiglietti:2025:NSD

[MP25] Arnaud Marsiglietti and Puja Pandey. A note on statistical distances for discrete log-concave measures. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002268.

Mandal:2020:ISP

[MPD20] B. N. Mandal, Rajender Parsad, and Sukanta Dash. Incomplete split-plot designs: Construction and analysis. Statistics & Probability Letters, 166(??):??, November 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301723.

Mathew:2022:JEL

[MPK22] Litty Mathew, Anisha P., and Sudheesh K. Kattumannil. A jackknife empirical likelihood ratio test for strong mean inactivity time order. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001511.

Moresco:2022:LBM

[MR22] Marlon Ruoso Moresco and Marcelo Brutti Righi. On the link between monetary and star-shaped risk measures. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522100290X.

Masmoudi:2023:IDM

[MR23] Afif Masmoudi and Hajer Rejeb. Infinitely divisible matrix gamma distribution: Asymptotic behaviour and parameters estimation. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522200270X.

Mishura:2024:FDB

[MR24] Yuliya Mishura and Kostiantyn Ralchenko. Fractional diffusion Bessel processes with Hurst index H ∈ (0, 1/2). Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002328.

Mishura:2022:PEC

[MRD22] Yuliya Mishura, Kostiantyn Ralchenko, and Olena Dehtiar. Parameter estimation in CKLS model by continuous observations. Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000153.

Mercadier:2022:LHS

[MRG22] Cécile Mercadier, Olivier Roustant, and Christian Genest. Linking the Hoeffding–Sobol and Möbius formulas through a decomposition of Kuo, Sloan, wasilkowski, and Woźniakowski. Statistics & Probability Letters, 185(??):??, June 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000335.

Mishura:2020:MWS

[MRZ20] Yuliya Mishura, Kostiantyn Ralchenko, and Mounir Zili. On mild and weak solutions for stochastic heat equations with piecewiseconstant conductivity. Statistics & Probability Letters, 159(??): ??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303281.

Marchand:2020:NSO

[MS20a] Éric Marchand and William E. Strawderman. On the nonstochastic ordering of some quadratic forms. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220301024.

Maurya:2020:PCF

[MS20b] Shambhu Nath Maurya and Koushik Saha. Process convergence of fluctuations of linear eigenvalue statistics of band Toeplitz matrices. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301784.

Mordant:2021:MNM

[MS21] Gilles Mordant and Johan Segers. Maxima and near-maxima of a Gaussian random assignment field. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221000493.

Maejima:2023:RCF

[MS23a] Makoto Maejima and Noriyoshi Sakuma. Rates of convergence in the free central limit theorem. Statistics & Probability Letters, 197 (??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000263.

Miller:2023:ASR

[MS23b] Joshua C. Miller and Natalia A. Stepanova. Adaptive signal recovery with Subbotin noise. Statistics & Probability Letters, 196 (??):??, May 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000159.

Montes:2020:MVS

[MSM20] Ignacio Montes, Juan Jesús Salamanca, and Susana Montes. A modified version of stochastic dominance involving dependence. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301516.

Merkle:2020:BAP

[MST20] Milan Merkle, Yuri F. Saporito, and Rodrigo S. Targino. Bayesian approach for parameter estimation of continuous-time stochastic volatility models using Fourier transform methods. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302469.

Ma:2021:GER

[MT21a] Hanmin Ma and Dejian Tian. Generalized entropic risk measures and related BSDEs. Statistics & Probability Letters, 174 (??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000729.

Mania:2021:MTM

[MT21b] M. Mania and R. Tevzadze. On martingale transformations of multidimensional Brownian motion. Statistics & Probability Letters, 175(??):??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771522100081X.

Miao:2021:LDI

[MT21c] Yu Miao and Yanyan Tang. Large deviation inequalities of LS estimator in nonlinear regression models. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302339.

Mallein:2022:LSP

[MT22] Bastien Mallein and Pavel Tesemnikov. On the length of the shortest path in a sparse Barak–Erdős graph. Statistics & Probability Letters, 190(??):??, November 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001651.

Maeda:2023:EWT

[MU23] Takashi Maeda and Yuya Unten. On expected waiting time until given words appear in random sequence. Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000329.

Muirhead:2020:SMB

[Mui20] Stephen Muirhead. A second moment bound for critical points of planar Gaussian fields in shrinking height windows. Statistics & Probability Letters, 160(??):??, May 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300018.

Mustafayev:2022:NKI

[Mus22] Heybetkulu Mustafayev. A note on the Kawada–Itô theorem. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002236.

Milbradt:2020:HPB

[MW20] Cassandra Milbradt and Martin Wahl. High-probability bounds for the reconstruction error of PCA. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220300444.

Meng:2022:SPT

[MX22] Weiwei Meng and Chengxun Xi. Some properties of 2-type Markov branching processes with immigration and instantaneous resurrection. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001250.

Miao:2024:CMD

[MY24] Yu Miao and Qing Yin. Cramér’s moderate deviations for the LS estimator of the autoregressive processes in the neighborhood of the unit root. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000622.

Myronyuk:2021:CTI

[Myr21]  Margaryta Myronyuk.  Characterization theorems for Q-

independent random variables with values in a Banach space. Statistics & Probability Letters, 168(??):??, January 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302492.

Ma:2022:AMD

[MYX22] Xiaocui Ma, Haitao Yue, and Fubao Xi. The averaging method for doubly perturbed distribution dependent SDEs. Statistics & Probability Letters, 189(??):??, October 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001389.

Mahmoud:2020:DCD

[MZ20] Hosam M. Mahmoud and Panpan Zhang. Distributions in the constant-differentials Pólya process. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521930238X.

Malyshkin:2021:MLR

[MZ21] Y. A. Malyshkin and M. E. Zhukovskii. MSO 0–1 law for recursive random trees. Statistics & Probability Letters, 173(??): ??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000237.

Naaman:2021:TCM

[Naa21] Michael Naaman. On the tight constant in the multivariate Dvoretzky–Kiefer–Wolfowitz inequality. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771522100050X.

Nakata:2022:LDS

[Nak22] Toshio Nakata. Large deviations for super-heavy tailed random walks. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002029.

Nalecz:2021:MHF

[Nał21] Marek Nałȩcz. Matrix hypergeometric function and its application to computation of characteristic function of spherically symmetric distributions with phase-type amplitude. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000249.

Nasri:2020:NCS

[Nas20] Bouchra R. Nasri. On non-central squared copulas. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300079.

Novo:2021:NPS

[NAV21] Silvia Novo, Germán Aneiros, and Philippe Vieu. A k NN procedure in semiparametric functional data analysis. Statistics & Probability Letters, 171(??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522030331X.

Hu:2021:RCS

[nHL21] Zhang nan Hu and Bing Li. Random covering sets in metric space with exponentially mixing property. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522030225X.

Nielsen:2022:RCP

[Nie22] Michael Nielsen. Regular conditional probabilities and strictly proper loss functions. Statistics & Probability Letters, 185(??): ??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522200030X.

Nishimori:2024:BHP

[Nis24] Yasuhito Nishimori. A bidirectional hitting probability for the symmetric Hunt processes. Statistics & Probability Letters, 204 (??):??, January 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001669.

Nguyen:2024:RGU

[NLS24] Hoang Nam Nguyen, Abdel Lisser, and Vikas Vikram Singh. Random games under normal mean–variance mixture distributed independent linear joint chance constraints. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000051.

Nieman:2025:IED

[NMR25] Dennis Nieman, Michel Mandjes, and Liron Ravner. Input estimation from discrete workload observations in a Lévy-driven storage system. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002190.

Nazarov:2021:GPC

[NN21] A. I. Nazarov and Ya. Yu. Nikitin. Gaussian processes centered at their online average, and applications. Statistics & Probability Letters, 170(??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303163.

Nane:2020:ABS

[NNO20] Erkan Nane, Eze R. Nwaeze, and McSylvester Ejighikeme Omaba. Asymptotic behaviour of solution and non-existence of global solution to a class of conformable time-fractional stochastic equation. Statistics & Probability Letters, 163(??):??, August 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522030095X.

Nogales:2021:NCE

[Nog21] A. G. Nogales. A note on conditional expectation for Markov kernels. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001590.

Nasari:2021:CES

[NOH21] Masoud M. Nasari and Mohamedou Ould-Haye. A consistent estimator for skewness of partial sums of dependent data. Statistics & Probability Letters, 171(??):??, April 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303242.

Novak:2022:T

[Nov22] S. Y. Novak. On the T -test. Statistics & Probability Letters, 189 (??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001201.

Nguyen:2020:ESS

[NP20] Giang T. Nguyen and Oscar Peralta. An explicit solution to the Skorokhod embedding problem for double exponential increments. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030170X.

Nourdin:2022:GFG

[NP22] Ivan Nourdin and Fei Pu. Gaussian fluctuation for Gaussian Wishart matrices of overall correlation. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002315.

Nazarov:2024:SBA

[NP24] Alexander Nazarov and Yulia Petrova. L_{2}-small ball asymptotics for some demeaned Gaussian processes. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002146.

Neufeld:2021:SMO

[NS21] Ariel Neufeld and Julian Sester. On the stability of the martingale optimal transport problem: a set-valued map approach. Statistics & Probability Letters, 176(??):??, September 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000936.

Nakajima:2022:ANL

[NS22] Shohei Nakajima and Yasutaka Shimizu. Asymptotic normality of least squares type estimators to stochastic differential equations driven by fractional Brownian motions. Statistics & Probability Letters, 187(??):??, August 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000657.

Nair:2023:SUR

[NS23a] R. Dhanya Nair and E. I. Abdul Sathar. Some useful results related to various measures of extropy and their interrelationship. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002425.

Nakamura:2023:IDD

[NS23b] Takashi Nakamura and Masatoshi Suzuki. On infinitely divisible distributions related to the Riemann hypothesis. Statistics & Probability Letters, 201(??):??, October 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522300113X.

Nevzorov:2024:UDS

[NS24] Valery B. Nevzorov and Alexei Stepanov. On uppermost discrete spacing. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000300.

Nummi:2024:NSC

[NV24] Patrik Nummi and Lauri Viitasaari. Necessary and sufficient conditions for continuity of hypercontractive processes and fields.

Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400018X.

Noguchi:2024:AOS

[NW24] Kimihiro Noguchi and Mayla C. Ward. Asymptotic optimality of the square-root transformation on the gamma distribution using the Kullback–Leibler information number criterion. Statistics & Probability Letters, 210(??):??, July 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000877.

Najnudel:2020:MBF

[NWY20] Joseph Najnudel, Ching-Tang Wu, and Ju-Yi Yen. On moments of Brownian functionals and their interpretation in terms of random walks. Statistics & Probability Letters, 161(??):??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300274.

Oakes:2025:IWR

[Oak25] David Oakes. On the intransitivity of the win ratio. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002360.

Ogasawara:2023:WDT

[Oga23] Haruhiko Ogasawara. The Wishart distribution with two different degrees of freedom. Statistics & Probability Letters, 200(??): ??, September 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000901.

Osorio:2024:GTS

[OGR24] Felipe Osorio, Ángelo Gárate, and Cibele M. Russo. The gradient test statistic for outlier detection in generalized estimating equations. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000567.

Orihara:2021:DON

[OH21] Shunichiro Orihara and Etsuo Hamada. Determination of the optimal number of strata for propensity score subclassification. Statistics & Probability Letters, 168(??):??, January 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302546.

Okudo:2021:SPS

[OK21] Michiko Okudo and Fumiyasu Komaki. Shrinkage priors for singlespiked covariance models. Statistics & Probability Letters, 176(??): ??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000894.

Okamura:2020:ECI

[Oka20] Kazuki Okamura. An equivalence criterion for infinite products of Cauchy measures. Statistics & Probability Letters, 163(??): ??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301000.

Olsen:2024:NLI

[Ols24] Niels Lundtorp Olsen. A new likelihood inequality for models with latent variables. Statistics & Probability Letters, 206(??): ??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002225.

Osekowski:2021:EPI

[Osȩ21] Adam Osȩkowski. An extension of Pratelli’s inequality. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001371.

Osekowski:2024:SET

[Osȩ24] Adam Osȩkowski. A sharp L^{p} estimate for the total variation process. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002213.

Ostrovski:2022:TEB

[Ost22a] Vladimir Ostrovski. Testing equivalence to binary generalized linear models with application to logistic regression. Statistics & Probability Letters, 191(??):??, December 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001778.

Ostrovski:2022:TEP

[Ost22b] Vladimir Ostrovski. Testing equivalence to power law distributions. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002492.

Oudghiri:2021:GET

[Oud21] Yannis Oudghiri. Generalizations of Efron’s theorem. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001206.

Okuno:2023:DVC

[OY23] Akifumi Okuno and Keisuke Yano. Dependence of variance on covariate design in nonparametric link regression. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002292.

Pacini:2020:PSA

[Pac20] David Pacini. Proximal statistic: Asymptotic normality. Statistics & Probability Letters, 167(??):??, December 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301991.

Pakshirajan:2020:PTL

[Pak20] R. P. Pakshirajan. Power type law of the iterated logarithm for Gaussian variables. Statistics & Probability Letters, 157(??): ??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302688.

Pashley:2022:NDM

[Pas22] Nicole E. Pashley. Note on the delta method for finite population inference with applications to causal inference. Statistics & Probability Letters, 188(??):??, September 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001055.

Pirenne:2024:EPS

[PC24] Sarah Pirenne and Gerda Claeskens. Exact post-selection inference for adjusted R squared selection. Statistics & Probability Letters, 211(??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001020.

Perez-Cendejas:2024:SOH

[PCPS24] Ulises Pérez-Cendejas and Gerardo Pérez-Suárez. Stochastic ordering for hitting times of fractional Brownian motions. Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000221.

Park:2020:NMR

[PD20] Seunghwan Park and Michael J. Daniels. A note on monotonicity in repeated attempt selection models. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302317.

Power:2021:BMA

[PD21] Michael Declan Power and Yuexiao Dong. Bayesian model averaging sliced inverse regression. Statistics & Probability Letters, 174 (??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000651.

Pekalp:2022:SNB

[Pek22] Mustafa Hilmi Pekalp. Some new bounds for the mean value function of the residual lifetime process. Statistics & Probability Letters, 187(??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000803.

Peng:2023:AEE

[Pen23] Jingfu Peng. Adaptive and efficient estimation in the Gaussian sequence model. Statistics & Probability Letters, 195(??): ??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002759.

Perreault:2024:SCK

[Per24] Samuel Perreault. Simultaneous computation of Kendall’s tau and its jackknife variance. Statistics & Probability Letters, 213(??): ??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001500.

Perez-Fernandez:2023:URE

[PF23] Raúl Pérez-Fernández. On the use of robust estimators of multivariate location for heterogeneous data. Statistics & Probability Letters, 203(??):??, December 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522300144X.

Pauwelyn:2020:PND

[PG20] P. J. Pauwelyn and M. A. Guerry. On perturbations of non-diagonalizable stochastic matrices of order 3. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302792.

Pandey:2021:DDG

[PG21] Kaushlendra Pandey and Abhishek K. Gupta. k-th distance distributions for generalized Gauss–Poisson process in R^{n}. Statistics & Probability Letters, 172(??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000109.

Pinelis:2020:MPP

[Pin20] Iosif Pinelis. Monotonicity properties of the Poisson approximation to the binomial distribution. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302042.

Pinelis:2021:BLB

[Pin21a] Iosif Pinelis. Best lower bound on the probability of a binomial exceeding its expectation. Statistics & Probability Letters, 179(??): ??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001863.

Pinelis:2021:MPP

[Pin21b] Iosif Pinelis. Monotonicity preservation properties of kernel regression estimators. Statistics & Probability Letters, 177(??): ??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522100119X.

Pinelis:2021:MPG

[Pin21c] Iosif Pinelis. Monotonicity properties of the gamma family of distributions. Statistics & Probability Letters, 171(??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220303308.

Pinelis:2022:ICB

[Pin22] Iosif Pinelis. Improved concentration bounds for sums of independent sub-exponential random variables. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200181X.

Pires:2024:LFI

[Pir24] Benito Pires. Lyapunov function for interacting reinforced stochastic processes via Hopfield’s energy function. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001815.

Patrone:2024:MUP

[PK24] Paul N. Patrone and Anthony J. Kearsley. Minimizing uncertainty in prevalence estimates. Statistics & Probability Letters, 205(??):

??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001700.

Peligrad:2023:CSC

[PP23] M. Peligrad and C. Peligrad. Convergence of series of conditional expectations. Statistics & Probability Letters, 200(??):??, September 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000937.

Pratelli:2021:ASC

[PR21] Luca Pratelli and Pietro Rigo. On the almost sure convergence of sums. Statistics & Probability Letters, 172(??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000079.

Privault:2021:RCH

[Pri21] Nicolas Privault. Recursive computation of the Hawkes cumulants. Statistics & Probability Letters, 177(??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001231.

Proia:2021:CPS

[PS21] Frédéric Proïa and Marius Soltane. Comments on the presence of serial correlation in the random coefficients of an autoregressive process. Statistics & Probability Letters, 170(??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302911.

Panda:2025:UCK

[PSP^{+}25] Sambit Panda, Cencheng Shen, Ronan Perry, Jelle Zorn, Antoine Lutz, Carey E. Priebe, and Joshua T. Vogelstein. Universally consistent k-sample tests via dependence measures. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002475.

Przybylowicz:2021:EUS

[PSX21] Paweł Przybyłowicz, Michaela Szölgyenyi, and Fanhui Xu. Existence and uniqueness of solutions of SDEs with discontinuous drift and finite activity jumps. Statistics & Probability Letters, 174 (??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000341.

Peligrad:2024:LLT

[PSZ24] Magda Peligrad, Hailin Sang, and Na Zhang. On the local limit theorems for linear sequences of lower psi-mixing Markov chains. Statistics & Probability Letters, 210(??):??, July 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000774.

Perrone:2023:KTE

[PvdHZ23] Elisa Perrone, Edwin R. van den Heuvel, and Zhuozhao Zhan. Kendall’s tau estimator for bivariate zero-inflated count data. Statistics & Probability Letters, 199(??):??, August 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000822.

Pycke:2021:NFO

[Pyc21] Jean-Renaud Pycke. A new family of omega-square-type statistics with Bahadur local optimality for the location family of generalized logistic distributions. Statistics & Probability Letters, 170 (??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220303023.

Peng:2020:HTP

[PZ20]  Shige Peng and Quan Zhou.  A hypothesis-testing per-

spective on the G-normal distribution theory. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771521930269X.

Pang:2024:SNF

[PZL24] Shanqi Pang, Yan Zhu, and Xiao Lin. Some new families of asymmetric nested orthogonal arrays with any strength. Statistics & Probability Letters, 214(??):??, November 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001883.

Quang:2025:CRL

[QHK25] Nguyen Van Quang, Nguyen Van Huan, and Phan Tri Kien. Convergence rates in the limit theorems for random sums of morthogonal random variables. Statistics & Probability Letters, 216 (??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002177.

Qiu:2022:DPM

[QK22] Qinjing Qiu and Reiichiro Kawai. A decoupling principle for Markov-modulated chains. Statistics & Probability Letters, 182 (??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002637.

Qiang:2023:RSE

[QP23] Beidi Qiang and Edsel A. Peña. Robust simultaneous estimation of location parameters. Statistics & Probability Letters, 193(??): ??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002437.

Qian:2020:HTI

[QTLT20] Manling Qian, Li Tao, Erqian Li, and Maozai Tian. Hypothesis testing for the identity of high-dimensional covariance matrices. Statistics & Probability Letters, 161(??):??, June 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522030002X.

Qiao:2021:SDS

[QW21] Huijie Qiao and Jiang-Lun Wu. Supports for degenerate stochastic differential equations with jumps and applications. Statistics & Probability Letters, 177(??):??, October 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001383.

Qian:2023:PBR

[QX23] Zhongmin Qian and Xingcheng Xu. Probability bounds for reflecting diffusion processes. Statistics & Probability Letters, 199 (??):??, August 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000792.

Rigon:2023:CPB

[RA23] Tommaso Rigon and Emanuele Aliverti. Conjugate priors and bias reduction for logistic regression models. Statistics & Probability Letters, 202(??):??, November 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001256.

Radchenko:2023:APW

[Rad23] Vadym Radchenko. Averaging principle for the wave equation driven by a stochastic measure. Statistics & Probability Letters, 201(??):??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223001128.

Ratanov:2021:TPT

[Rat21] Nikita Ratanov. On telegraph processes, their first passage times and running extrema. Statistics & Probability Letters, 174(??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000638.

Ruzankin:2020:RPA

[RB20] Pavel S. Ruzankin and Igor S. Borisov. On the rate of Poisson approximation to Bernoulli partial sum processes. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300572.

Roy:2024:ABO

[RB24] Rahul Roy and Subir Kumar Bhandari. Asymptotic Bayes’ optimality under sparsity for exchangeable dependent multivariate normal test statistics. Statistics & Probability Letters, 207(??): ??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002535.

Renaud:2024:NOD

[Ren24] Jean-François Renaud. A note on the optimal dividends problem with transaction costs in a spectrally negative Lévy model with Parisian ruin. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522300202X.

Roy:2020:STI

[RG20] Angshuman Roy and Anil K. Ghosh. Some tests of independence based on maximum mean discrepancy and ranks of nearest neighbors. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300961.

Rajesh:2022:KEE

[RGS22] Richu Rajesh, Rajesh G., and S. M. Sunoj. Kernel estimation of extropy function under length-biased sampling. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002522.

Ribas:2020:ELS

[Rib20] J. M. Grau Ribas. An extension of the Last-Success-Problem. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302378.

Riekert:2022:CRE

[Rie22] Adrian Riekert. Convergence rates for empirical measures of Markov chains in dual and Wasserstein distances. Statistics & Probability Letters, 189(??):??, October 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001468.

Rivest:2021:LPE

[Riv21] Louis-Paul Rivest. Limiting properties of an equiprobable sampling scheme for 0–1 matrices. Statistics & Probability Letters, 172 (??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000092.

Ramtirthkar:2022:NLA

[RK22] Mukund Ramtirthkar and Mohan Kale. A note on the local asymptotic mixed normality of a controlled branching process with a random control function. Statistics & Probability Letters, 181(??): ??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002327.

Rajchert:2023:CFD

[RK23] Andrew Rajchert and Uri Keich. Controlling the false discovery rate via competition: Is the +1 needed? Statistics & Probability Letters, 197(??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000433.

Ruzankin:2020:FME

[RL20] Pavel S. Ruzankin and Artem V. Logachov. A fast mode estimator in multidimensional space. Statistics & Probability Letters, 158 (??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303165.

Ratibenyakool:2024:RCT

[RN24] Yuttana Ratibenyakool and Kritsana Neammanee. Rate of convergence of trinomial formula to Black–Scholes formula. Statistics & Probability Letters, 213(??):??, October 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001366.

Rozovsky:2021:OMC

[Roz21] L. V. Rozovsky. One more on the convergence rates in precise asymptotics. Statistics & Probability Letters, 171(??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220303266.

Ramos:2020:PPW

[RRL20] Eduardo Ramos, Pedro L. Ramos, and Francisco Louzada. Posterior properties of the Weibull distribution for censored data. Statistics & Probability Letters, 166(??):??, November 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301760.

Robertson:2021:QRR

[RRPB21] B. L. Robertson, M. Reale, C. J. Price, and J. A. Brown. Quasirandom ranked set sampling. Statistics & Probability Letters, 171 (??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220303321.

Russell:2022:OGP

[RS22] Oliver Russell and Wei Sun. An opposite Gaussian product inequality. Statistics & Probability Letters, 191(??):??, December 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001766.

Rahimov:2024:DAN

[RS24] I. Rahimov and S. O. Sharipov. On deterministic approximation for nearly critical branching processes with dependent immigration. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002547.

Rosalsky:2021:NSD

[RT21] Andrew Rosalsky and Lê Vǎn Thành. A note on the stochastic domination condition and uniform integrability with applications to the strong law of large numbers. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001437.

Rasonyi:2022:SSG

[RT22] Miklós Rásonyi and Kinga Tikosi. On the stability of the stochastic gradient Langevin algorithm with dependent data stream. Statistics & Probability Letters, 182(??):??, March 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002789.

Ren:2021:DIT

[RTT21] Quanwei Ren, Hongjiong Tian, and Tianhai Tian. Deterministic implicit two-step Milstein methods for stochastic differential equations. Statistics & Probability Letters, 179(??):??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522100170X.

Roozegar:2021:ADR

[Rz21] Rasool Roozegar and Hamid Reza Taherizadeh zarch. On the asymptotic distribution of randomly weighted averages of random vectors. Statistics & Probability Letters, 179(??):??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001930.

Rice:2022:CBS

[RZ22] Gregory Rice and Chi Zhang. Consistency of binary segmentation for multiple change-point estimation with functional data. Statistics & Probability Letters, 180(??):??, January 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001905.

Said:2020:SRA

[Sai20] A. Ouled Said. Some remark on the asymptotic variance in a drift accelerated diffusion. Statistics & Probability Letters, 162 (??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300511.

Saigo:2021:IAM

[Sai21] Hiroshi Saigo. The independence assumption in the mixed randomized response model. Statistics & Probability Letters, 170(??): ??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302893.

Saigo:2025:MSE

[Sai25] Hiroshi Saigo. Mean square error and variance estimation of the sample ratio under Lahiri’s design. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002463.

Sang:2024:TDS

[San24] Yongli Sang. Test for diagonal symmetry in high dimension. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001840.

Sarantsev:2020:CRE

[Sar20] Andrey Sarantsev. Convergence rate to equilibrium in Wasserstein distance for reflected jump–diffusions. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220301632.

Sarantsev:2021:OPP

[Sar21a] Andrey Sarantsev. Optimal portfolio with power utility of absolute and relative wealth. Statistics & Probability Letters, 179(??): ??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001875.

Sarantsev:2021:SER

[Sar21b] Andrey Sarantsev. Sub-exponential rate of convergence to equilibrium for processes on the half-line. Statistics & Probability Letters, 175(??):??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221000778.

Sahu:2021:NMA

[SB21] Nitesh Sahu and Prabhu Babu. A new monotonic algorithm for the E-optimal experiment design problem. Statistics & Probability Letters, 174(??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000596.

Shah:2020:CTG

[SBK20] Imtiyaz A. Shah, H. M. Barakat, and A. H. Khan. Characterizations through generalized and dual generalized order statistics, with an application to statistical prediction problem. Statistics & Probability Letters, 163(??):??, August 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300857.

Shi:2020:NRE

[SBS20] Jianhong Shi, Xiuqin Bai, and Weixing Song. Nonparametric regression estimate with Berkson Laplace measurement error. Statistics & Probability Letters, 166(??):??, November 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522030167X.

Shi:2022:TTF

[SBS22] Jianhong Shi, Xiuqin Bai, and Weixing Song. Tweedietype formulae for a multivariate Laplace distribution. Statistics & Probability Letters, 183(??):??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002820.

Sun:2020:TLR

[SCGW20] Hongwei Sun, Yuehua Cui, Qian Gao, and Tong Wang. Trimmed LASSO regression estimator for binary response data. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303256.

Song:2024:SWQ

[SCMZ24] Yuping Song, Chunchun Cai, Huijue Mao, and Min Zhu. Selfweighted quantile regression estimation for diffusion parameter in jump-diffusion models. Statistics & Probability Letters, 206 (??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002341.

Sang:2020:ELT

[SD20] Yongli Sang and Xin Dang. Empirical likelihood test for diagonal symmetry. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521930241X.

Solari:2022:MSC

[SD22a] Aldo Solari and Vera Djordjilović. Multi split conformal prediction. Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000177.

Stepanov:2022:LTU

[SD22b] Alexei Stepanov and Anna Dembińska. Limit theorems for the uppermost m-th spacing based on weak geometric records. Statistics & Probability Letters, 183(??):??, April 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002935.

Singh:2020:SOS

[SDH20] Rakhi Singh, Ashish Das, and Daniel Horsley. SUE(s^{2})-optimal supersaturated designs. Statistics & Probability Letters, 158(??): ??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303190.

Selk:2024:LDH

[Sel24] Zachary Selk. Large deviations for high minima of Gaussian processes with nonnegatively correlated increments. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002250.

Sepanski:2020:NDE

[Sep20] Jungsywan H. Sepanski. A note on distortion effects on the strength of bivariate copula tail dependence. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301978.

Sester:2024:MMC

[Ses24] Julian Sester. A multi-marginal c-convex duality theorem for martingale optimal transport. Statistics & Probability Letters, 210 (??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000816.

Shimizu:2022:APJ

[SH22] Koki Shimizu and Hiroki Hashiguchi. Algorithm for the product of Jack polynomials and its application to the sphericity test. Statistics & Probability Letters, 187(??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000840.

Shashkov:2025:RPA

[Sha25] Timofei Shashkov. Ruin probability approximation for bidimensional Brownian risk model with tax. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224002748.

Shin:2020:LDB

[Shi20] Jiyong Shin. Liouville distorted Brownian motion. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302366.

Shpilev:2024:ODM

[Shp24] P. V. Shpilev. D-optimal designs for a multidimensional second-degree polynomial model with no intercept. Statistics & Probability Letters, 215(??):??, December 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001974.

Singh:2023:SGF

[Sin23] Rahul Singh. Some goodness of fit tests based on centre outward spacings. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002516.

Sun:2022:LSE

[SJK22] Chuanfeng Sun, Shaolin Ji, and Chuiliu Kong. The least squares estimator of random variables under convex operators on L_{F}^{∞}(μ) space. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002303.

Shen:2021:GFT

[SJSL21] Xiaoxi Shen, Chang Jiang, Lyudmila Sakhanenko, and Qing Lu. A goodness-of-fit test based on neural network sieve estimators. Statistics & Probability Letters, 174(??):??, July 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000626.

Shu:2023:PEI

[SJZ23] Huisheng Shu, Ziwei Jiang, and Xuekang Zhang. Parameter estimation for integrated Ornstein–Uhlenbeck processes with small Lévy noises. Statistics & Probability Letters, 199(??):??, August 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000755.

Skorniakov:2019:ANC

[Sko19] V. Skorniakov. On asymptotic normality of certain linear rank statistics. Statistics & Probability Letters, 145(??):63–73, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302918. See erratum [Sko20].

Skorniakov:2020:EAN

[Sko20] V. Skorniakov. Erratum to “On asymptotic normality of certain linear rank statistics” [Statist. Probab. Lett. 145 (2019) 63–73]. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302548.

Song:2025:EBC

[SL25] Chaegeun Song and Bing Li. On exact Bayesian credible sets for discrete parameters. Statistics & Probability Letters, 218(??): ??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002645.

Singh:2021:EMP

[SLA21] Vikas Vikram Singh, Abdel Lisser, and Monika Arora. An equivalent mathematical program for games with random constraints. Statistics & Probability Letters, 174(??):??, July 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000547.

S:2022:ABL

[SM22] Kiran Kumar A. S. and Shambhu Nath Maurya. Asymptotic behaviour of linear eigenvalue statistics of Hankel matrices. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002352.

Spencer:2024:SUL

[SM24] Neil A. Spencer and Jeffrey W. Miller. Strong uniform laws of large numbers for bootstrap means and other randomly weighted sums. Statistics & Probability Letters, 211(??):??, August 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001135.

Santos:2024:NIC

[SMRH24] Samuel S. Santos, Marlon R. Moresco, Marcelo B. Righi, and Eduardo Horta. A note on the induction of comonotonic additive risk measures from acceptance sets. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000130.

Sherwood:2020:APC

[SMS20] Ben Sherwood, Aaron J. Molstad, and Sumanta Singha. Asymptotic properties of concave L_{1}-norm group penalties. Statistics & Probability Letters, 157(??):??, February 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302779.

Shen:2020:FMS

[SN20] Chengwu Shen and Anh Ninh. Finding the modes of some multivariate discrete probability distributions: Application of the resource allocation problem. Statistics & Probability Letters, 156 (??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302251.

Self:2024:BSS

[SN24] Stella Self and Melissa Nolan. A Bayesian spatial scan statistic for multinomial data. Statistics & Probability Letters, 206(??): ??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002298.

Soltani:2022:RIE

[Sol22] A. R. Soltani. Recursive integral equations for random weights averages: Exponential functions and Cauchy distribution. Statistics & Probability Letters, 190(??):??, November 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200147X.

Struleva:2022:ILL

[SP22] M. A. Struleva and E. I. Prokopenko. Integro-local limit theorems for supercritical branching process in a random environment. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001966.

Samanta:2024:NQS

[SP24a] Sujit Kumar Samanta and Aysha Parveen. A note on Geo^{X} /G/1 queueing system. Statistics & Probability Letters, 207(??): ??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000026.

Soni:2024:GFN

[SP24b] Ritik Soni and Ashok Kumar Pathak. Generalized fractional negative binomial process. Statistics & Probability Letters, 207 (??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002444.

Sjolander:2022:BRN

[SPG22] Arvid Sjölander, Jose M. Peña, and Erin E. Gabriel. Bias results for nondifferential mismeasurement of a binary confounder. Statistics & Probability Letters, 186(??):??, July 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000645.

Selvamuthu:2023:LTE

[SPT23] Dharmaraja Selvamuthu, Shamiksha Pandey, and Paola Tardelli. Limit Theorems for an extended inverse Hawkes process with general exciting functions. Statistics & Probability Letters, 197 (??):??, June 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522300041X.

Sun:2024:ETC

[SQT24] Chengfeng Sun, Zhaoyang Qiu, and Yanbin Tang. Ergodicity for two class stochastic partial differential equations with anisotropic viscosity. Statistics & Probability Letters, 207(??): ??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002456.

Sin:2022:EUS

[SRK22] Myong-Guk Sin, Kyong-Il Ri, and Kyong-Hui Kim. Existence and uniqueness of solution for coupled fractional meanfield forward–backward stochastic differential equations. Statistics & Probability Letters, 190(??):??, November 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001481.

Sandric:2023:HIN

[SŠ23a] Nikola Sandrić and Stjepan Šebek. Hoeffding’s inequality for non-irreducible Markov models. Statistics & Probability Letters, 200(??):??, September 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223000949.

Sen:2023:DAI

[SS23b] Ankita Sen and N. Selvaraju. Diffusion approximation of an infinite-server queue under Markovian environment with rapid switching. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000020.

Sarma:2024:SDT

[SS24] Utpal Jyoti Deba Sarma and Dharmaraja Selvamuthu. Study of discrete-time Hawkes process and its compensator. Statistics & Probability Letters, 214(??):??, November 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001615.

Sukeda:2025:FCM

[SS25] Issey Sukeda and Tomonari Sei. Frank copula is minimum information copula under fixed Kendall’s τ. Statistics & Probability Letters, 217(??):??, February 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400258X.

Sun:2022:LSD

[SSL22] Yan Sun, Qifan Song, and Faming Liang. Learning sparse deep neural networks with a spike-and-slab prior. Statistics & Probability Letters, 180(??):??, January 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522100208X.

Siri:2021:RCI

[ST21] Paola Siri and Barbara Trivellato. Robust concentration inequalities in maximal exponential models. Statistics & Probability Letters, 170(??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220303047.

Stewart:2023:TIN

[Ste23] Jonathan R. Stewart. On the time to identify the nodes in a random graph. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000032.

Stover:2020:SCR

[Sto20] Joseph P. Stover. A stochastic comparison result for the multitype contact process with unequal death rates. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300663.

Subramanian:2021:MRT

[Sub21] Sundarraman Subramanian. Median regression from twice censored data. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220302583.

Subramanian:2022:SCB

[Sub22] Sundarraman Subramanian. Simultaneous confidence bands for survival functions from twice censorship. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222000797.

Sun:2020:GAM

[Sun20] Qiang Sun. Gaussian approximations for maxima of random vectors under (2+ι)-th moments. Statistics & Probability Letters, 158 (??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301518.

Sun:2025:CTM

[Sun25] Mingyang Sun. Cramér type moderate deviation for random walks conditioned to stay positive. Statistics & Probability Letters, 216 (??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522400227X.

Suzuki:2018:MDI

[Suz18] Ryoichi Suzuki. Malliavin differentiability of indicator functions on canonical Lévy spaces. Statistics & Probability Letters, 137(??): 183–190, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic).  URL http://www.sciencedirect.
com/science/article/pii/S016771521830035X. See correction

[Suz20].

Suzuki:2020:CMD

[Suz20] Ryoichi Suzuki. Correction to “Malliavin differentiability of indicator functions on canonical Lévy spaces” [Statist. Probab. Lett. 137 (2018) 183–190]. Statistics & Probability Letters, 156(??): ??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302603.

Sottinen:2020:PLM

[SV20] Tommi Sottinen and Lauri Viitasaari. Prediction law of mixed Gaussian Volterra processes. Statistics & Probability Letters, 156 (??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302408.

Song:2020:GQL

[SW20] Xinyu Song and Yazhen Wang. GARCH quasi-likelihood ratios for SV model and the diffusion limit. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301206.

Su:2020:MDN

[SWZ20] Zheren Su, Yaping Wang, and Yingchun Zhou. On maximin distance and nearly orthogonal Latin hypercube designs. Statistics & Probability Letters, 166(??):??, November 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301814.

Song:2021:JIE

[SWZ21] Wenjie Song, Panyu Wu, and Guodong Zhang. Jensen’s inequality for g-expectations in general filtration spaces. Statistics & Probability Letters, 169(??):??, February 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302613.

Song:2024:NLV

[SX24] He Song and Kainan Xiang. Note on locality of volume growth rate of graphs. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002183.

Sun:2020:DFF

[SXX20] Xiaobin Sun, Longjie Xie, and Yingchao Xie. Derivative formula for the Feynman–Kac semigroup of SDEs driven by rotationally invariant α-stable process. Statistics & Probability Letters, 158 (??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219303104.

Shen:2022:LSE

[SXY22] Leyi Shen, Xiaoyu Xia, and Litan Yan. Least squares estimation for the linear self-repelling diffusion driven by α-stable motions. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002212.

Su:2024:EVT

[SY24] Wen Su and Yaodi Yong. Estimating a VaR-type ruin measure by Laguerre series expansion in classical compound Poisson risk model. Statistics & Probability Letters, 205(??):??, February 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223001864.

Song:2025:UST

[SY25] Xiaojun Song and Zixin Yang. Unified specification tests in partially linear quantile regression models. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002128.

Sylvester:2022:CTM

[Syl22] John Sylvester. The cover time of a (multiple) Markov chain with rational transition probabilities is rational. Statistics & Probability Letters, 187(??):??, August 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200102X.

Shi:2025:SLD

[SYM25] Jianan Shi, Zhenhong Yu, and Yu Miao. A supplement to the large deviations of infinite weighted sums of heavy tailed random variables. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522400275X.

Sattler:2024:CHM

[SZ24] Paavo Sattler and Georg Zimmermann. Choice of the hypothesis matrix for using the Wald-type-statistic. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000075.

Szewczak:2020:MZS

[Sze20] Zbigniew S. Szewczak. On the Marcinkiewicz–Zygmund strong laws for arbitrary dependent sequences. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771522030198X.

Szewczak:2022:MTR

[Sze22] Zbigniew S. Szewczak. A. de Moivre theorem revisited. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002224.

Takabatake:2024:QLA

[Tak24] Tetsuya Takabatake. Quasi-likelihood analysis of fractional Brownian motion with constant drift under high-frequency observations. Statistics & Probability Letters, 207(??):??, April 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002304.

Talarczyk:2020:BBM

[Tal20] Anna Talarczyk. Bifractional Brownian motion for H > 1 and 2HK ≤ 1. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302743.

Tanoue:2024:CIW

[Tan24] Yuta Tanoue. Concentration inequality and the weak law of large numbers for the sum of partly negatively dependent ϕ - subgaussian random variables. Statistics & Probability Letters, 206 (??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002031.

Tappe:2021:NWT

[Tap21] Stefan Tappe. A note on the von Weizsäcker theorem. Statistics & Probability Letters, 168(??):??, January 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302297.

Tempelman:2023:GGC

[Tem23] Arkady Tempelman. Generalization of the Glivenko–Cantelli theorem to finite-dimensional distributions of ergodic homogeneous random fields. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002802.

Turner:2023:EAV

[TG23] Rosanne J. Turner and Peter D. Grünwald. Exact anytimevalid confidence intervals for contingency tables and beyond. Statistics & Probability Letters, 198(??):??, July 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000597.

Tran:2022:ESN

[TL22] Ky Q. Tran and Bich T. N. Le. On exponential stability of non-autonomous stochastic differential equations with Markovian switching. Statistics & Probability Letters, 189(??):??, October 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001456.

Tan:2023:RPG

[TL23] Jiangrui Tan and Junping Li. Recurrence property for Galton– Watson processes in which individuals have variable lifetimes. Statistics & Probability Letters, 199(??):??, August 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000731.

Tian:2022:NMF

[TLW22] Bing Tian, Zili Liu, and Hong Wang. Non-marginal feature screening for varying coefficient competing risks model. Statistics & Probability Letters, 190(??):??, November 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001729.

Tang:2022:ADM

[TLX22] Ping Tang, Rongrong Lu, and Junshan Xie. Asymptotic distribution of the maximum interpoint distance for high-dimensional data. Statistics & Probability Letters, 190(??):??, November 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001237.

Torricelli:2024:CET

[Tor24] Lorenzo Torricelli. On the convolution equivalence of tempered stable distributions on the real line. Statistics & Probability Letters, 207(??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224000038.

Touhami:2021:SSB

[Tou21] Wajdi Touhami. On skew sticky Brownian motion. Statistics & Probability Letters, 173(??):??, June 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000481.

Thu:2022:DNC

[TP22] Pham-Gia Thu and Duong Thanh Phong. The distribution of the non-central Wilks statistic in the complex case. Statistics & Probability Letters, 184(??):??, May 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000062.

Tran:2021:ECS

[Tra21a] Ky Quan Tran. Exponential contraction of switching jump diffusions with a hidden Markov chain. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771522100153X.

Trapani:2021:TSS

[Tra21b] Lorenzo Trapani. A test for strict stationarity in a random coefficient autoregressive model of order 1. Statistics & Probability Letters, 177(??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001267.

Trede:2020:MLE

[Tre20] Mark Trede. Maximum likelihood estimation of high-dimensional Student-t copulas. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219303244.

Tabrizi:2020:NIL

[TSG20] Elham Tabrizi, Ehsan Bahrami Samani, and Mojtaba Ganjali. A note on the identifiability of latent variable models for mixed longitudinal data. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301851.

Takahashi:2023:DPB

[TT23] Hiroshi Takahashi and Yozo Tamura. Diffusion processes in Brownian environments on disconnected selfsimilar fractal sets in R. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002073.

Tuitman:2020:CMA

[TVY20] Jan Tuitman, Steven Vanduffel, and Jing Yao. Correlation matrices with average constraints. Statistics & Probability Letters, 165 (??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220301711.

Tian:2021:GPD

[TWW21] Rongrong Tian, Jinlong Wei, and Jiang-Lun Wu. On a generalized population dynamics equation with environmental noise. Statistics & Probability Letters, 168(??):??, January 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302479.

Tzioufas:2020:MEP

[Tzi20] Achillefs Tzioufas. Monotonicity of escape probabilities for branching random walks on Z^{d}. Statistics & Probability Letters, 167(??):??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302030.

Tang:2021:LTP

[TZT21] Linjun Tang, Shengchao Zheng, and Zhongquan Tan. Limit theorem on the pointwise maxima of minimum of vector-valued Gaussian processes. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000997.

Valentic:2023:SPD

[Val23] Ivana Valentić. Skew-product decomposition of Brownian motion on an ellipsoid. Statistics & Probability Letters, 195(??): ??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223000081.

Vila:2024:UBC

[VBS24] Roberto Vila, Narayanaswamy Balakrishnan, and Helton Saulo. An upper bound and a characterization for Gini’s mean difference based on correlated random variables. Statistics & Probability Letters, 207(??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000014.

Verreault:2022:PFP

[Ver22] William Verreault. On the probability of forming polygons from a broken stick. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001991.

Videla:2020:EMB

[Vid20] Leonardo A. Videla. On the expected maximum of a birth-and-death process. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219303116.

Vidmar:2021:NSC

[Vid21a] Matija Vidmar. A nonclassical solution to a classical SDE and a converse to Kolmogorov’s zero–one law. Statistics & Probability Letters, 175(??):??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000791.

Vidotto:2021:NRP

[Vid21b] Anna Vidotto. A note on the reduction principle for the nodal length of planar random waves. Statistics & Probability Letters, 174(??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000523.

Vidmar:2023:CMT

[Vid23] Matija Vidmar. Complete monotonicity of time-changed Lévy processes at first passage. Statistics & Probability Letters, 193(??): ??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002231.

VilelaMendes:2024:FLP

[Vil24] R. Vilela Mendes. On a family of Lévy processes without support in S′. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000464.

Virta:2021:TSW

[Vir21] Joni Virta. Testing for subsphericity when n and p are of different asymptotic order. Statistics & Probability Letters, 179(??): ??, December 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001711.

Voutilainen:2023:NAB

[VIVL23] Marko Voutilainen, Pauliina Ilmonen, Lauri Viitasaari, and Niko Lietzén. Note on asymptotic behavior of spatial sign autocovariance matrices. Statistics & Probability Letters, 192(??):??, January 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222001924.

Vishwakarma:2024:IBD

[VK24] P. Vishwakarma and K. K. Kataria. On integrals of birth–death processes at random time. Statistics & Probability Letters, 214(??): ??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001731.

Valjarevic:2021:SCM

[VM21]  Dragana Valjarević and Ana Merkle.  Statistical causal-

ity and measurable separability of σ-algebras. Statistics & Probability Letters, 177(??):??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001280.

vanNiekerk:2021:PDB

[vNBR21] J. van Niekerk, H. Bakka, and H. Rue. A principled distancebased prior for the shape of the Weibull model. Statistics & Probability Letters, 174(??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000602.

Vogt:2020:DBB

[Vog20] Michael Vogt. On the differences between L_{2} boosting and the lasso. Statistics & Probability Letters, 157(??):??, February 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302809.

Vogel:2021:NIT

[Vog21] Quirin Vogel. A note on the intersections of two random walks in two dimensions. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001577.

Valjarevic:2020:SCS

[VP20] Dragana Valjarević and Ljiljana Petrović. Statistical causality and separable processes. Statistics & Probability Letters, 167(??): ??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302182.

Valdora:2020:EBP

[VY20] Marina Valdora and Víctor Yohai. M estimators based on the probability integral transformation with applications to count data. Statistics & Probability Letters, 162(??):??, July 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300547.

Waite:2024:RRT

[Wai24] Timothy W. Waite. Replication in random translation designs. Statistics & Probability Letters, 215(??):??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001986.

Walker:2023:CWS

[Wal23] Stephen G. Walker. Comparing weak and strong convergence of density functions. Statistics & Probability Letters, 200(??):??, September 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001025.

Wang:2022:LTD

[Wan22a] Haixu Wang. Limit theorems for a discrete-time marked Hawkes process. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000037.

Wang:2022:NTP

[Wan22b] Li Wang. New testing procedures with k-FWER control for discrete data. Statistics & Probability Letters, 180(??):??, January 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522100198X.

Wang:2022:ECR

[Wan22c] Tao Wang. Ergodic convergence rates for time-changed symmetric Lévy processes in dimension one. Statistics & Probability Letters, 183(??):??, April 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715221002893.

Wang:2021:LRO

[WC21] Jiantian Wang and Bin Cheng. On likelihood ratio ordering of parallel systems with heterogeneous exponential components. Statistics & Probability Letters, 175(??):??, August 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000808.

Wilsher:2023:DNI

[WDG23] Michael Wilsher, Carl P. Dettmann, and Ayalvadi J. Ganesh. The distribution of the number of isolated nodes in the 1dimensional soft random geometric graph. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222002085.

Wei:2020:NFR

[WDP^{+}20] Bo Wei, Tian Dai, Limin Peng, Ying Guo, and Amita Manatunga. A new functional representation of broad sense agreement.

Statistics & Probability Letters, 158(??):??, March 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302652.

Weiss:2021:EMC

[Wei21] Christian H. Weiß. On Edgeworth models for count time series. Statistics & Probability Letters, 171(??):??, April 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302972.

Weissman:2023:SCD

[Wei23] Ishay Weissman. Some comments on “The density flatness phenomenon” by Alhakim and Molchanov and the Dickman distribution. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002541. See [AM19].

Weiss:2024:HOM

[Wei24] Christian H. Weiß. On higher-order moments of INGARCH processes. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001676.

Wang:2024:LDY

[WFGR24] Xiaochang Wang, Shui Feng, Yiping Guo, and Bruno N. Rémillard. Large deviations for the Yule–Walker estimator of near critical autoregressive processes. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001652.

Wang:2024:ACC

[WFJ24] Ya Wang, Linjiajie Fang, and Bingyi Jing. Adapted Chatterjee correlation coefficient. Statistics & Probability Letters, 215(??): ??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002104.

Wang:2025:SDS

[WFZ25] Kai Wang, Hongjie Fan, and Yanling Zhu. Stationary distribution of a stochastic generalized SIRI epidemic model with reinfection and relapse. Statistics & Probability Letters, 216(??): ??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002487.

Wang:2023:CDB

[WG23] Tao Wang and Zhong Guan. Choice of degree of Bernstein polynomial model. Statistics & Probability Letters, 200(??):??, September 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000925.

Wu:2024:SBG

[WH24] Qianyong Wu and Jiang Hu. A spectral based goodness-of-fit test for stochastic block models. Statistics & Probability Letters, 209 (??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000737.

Wu:2022:ICI

[WjWbK22] Fan Wu, Guan jun Wang, and Xin bing Kong. Inference on common intraday periodicity at high frequencies. Statistics & Probability Letters, 191(??):??, December 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001717.

Wei:2021:MAA

[WK21] Zheng Wei and Daeyoung Kim. Measure of asymmetric association for ordinal contingency tables via the bilinear extension copula. Statistics & Probability Letters, 178(??):??, November 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221001450.

Wang:2021:HOC

[WL21a] Weixin Wang and Taeyoung Lee. Higher-order central moments of matrix Fisher distribution on SO(3). Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302868.

Wu:2021:CCT

[WL21b] Hao Wu and Xuefeng Li. Converse comparison theorems for multidimensional anticipated backward stochastic differential equations. Statistics & Probability Letters, 168(??):??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522030256X.

Wang:2022:DSD

[WLL22] Yaping Wang, Sixu Liu, and Dennis K. J. Lin. On definitive screening designs using Paley’s conference matrices. Statistics & Probability Letters, 181(??):??, February 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002297.

Wang:2020:REE

[WLS20] Zhanfeng Wang, Kai Li, and Jian Qing Shi. A robust estimation for the extended t-process regression model. Statistics & Probability Letters, 157(??):??, February 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771521930272X.

Wen:2024:SER

[WLZ24] Xin Wen, Yang Li, and Zemin Zheng. Scalable efficient reproducible multi-task learning via data splitting. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000403.

Wei:2021:DDF

[WM21] Yixi Wei and Jiang-Hong Ma. Determinacy of a distribution with finitely many mass points by finitely many moments. Statistics & Probability Letters, 176(??):??, September 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000973.

Wang:2024:ULR

[WM24] Yining Wang and Michael P. McDermott. On the unbiasedness of the likelihood ratio test for the multivariate one-sided problem. Statistics & Probability Letters, 215(??):??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002001.

Wang:2022:CID

[WML22] Nanwei Wang, Hélène Massam, and Qiong Li. Confidence intervals for discrete log-linear models when MLE does not exist. Statistics & Probability Letters, 187(??):??, August 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222001018.

Wright:2020:GEO

[Wri20] Tommy Wright. A general exact optimal sample allocation algorithm: With bounded cost and bounded sample sizes. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301322.

Wright:2025:OTK

[Wri25] Tommy Wright. Optimal tightening of the KWW joint confidence region for a ranking. Statistics & Probability Letters, 217(??): ??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002578.

Wen:2020:SAB

[WS20] Jiaqiang Wen and Yufeng Shi. Solvability of anticipated backward stochastic Volterra integral equations. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302457.

Wang:2024:ABW

[WSM24] Jixia Wang, Lu Sun, and Yu Miao. Asymptotic behavior of weighted quadratic variation of tempered fractional Brownian motion. Statistics & Probability Letters, 207(??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002432.

Wang:2021:MCP

[WW21a] Qian Wang and Yanhui Wang. Moments of characteristic polynomials in certain random neural networks. Statistics & Probability Letters, 172(??):??, May 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221000067.

Wei:2021:CVB

[WW21b] Yuting Wei and Qihua Wang. Cross-validation-based model averaging in linear models with response missing at random. Statistics & Probability Letters, 171(??):??, April 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302935.

Wang:2023:SDS

[WW23] Wei Wang and Xiulian Wang. Stationary distributions for stochastic differential equations with memory driven by α-stable processes. Statistics & Probability Letters, 195(??):??, April 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002796.

Wu:2020:CTO

[WWL20] Jintang Wu, Mengshou Wang, and Xiaohu Li. Convex transform order of the maximum of independent Weibull random variables. Statistics & Probability Letters, 156(??):??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302433.

Wang:2023:ABS

[WWL23] Juan Wang, Xueke Wang, and Junping Li. Asymptotic behavior for supercritical branching processes. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223000068.

Wang:2020:LPP

[WWZ20] Feng Wang, Xian-Yuan Wu, and Rui Zhu. Last passage percolation on the complete graph. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220301012.

Wang:2021:EHD

[WX21] Moming Wang and Ningning Xia. Estimation of high-dimensional integrated covariance matrix based on noisy high-frequency data with multiple observations. Statistics & Probability Letters, 170 (??):??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302996.

Wang:2022:SAN

[WXY22] Ruifang Wang, Yong Xu, and Hongge Yue. Stochastic averaging for the non-autonomous mixed stochastic differential equations with locally Lipschitz coefficients. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522100256X.

Wu:2021:RCC

[WYWS21] Yi Wu, Wei Yu, Xuejun Wang, and Aiting Shen. The rate of complete consistency for recursive probability density estimator under strong mixing samples. Statistics & Probability Letters, 176(??): ??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000924.

Wang:2024:MPS

[WZ24a] Dongying Wang and Qi Zhou. A maximum projective stratification criterion for selecting space-filling designs. Statistics & Probability Letters, 209(??):??, June 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000543.

Wang:2024:MDP

[WZ24b] Shaochen Wang and Yangchun Zhang. Moderate deviation principle for guesswork. Statistics & Probability Letters, 207(??):

??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000117.

Wang:2024:TSS

[WZ24c] Xuanci Wang and Bin Zhang. Target selection in shrinkage estimation of covariance matrix: a structural similarity approach. Statistics & Probability Letters, 208(??):??, May 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224000178.

Wei:2021:TSA

[WZGY21] Yanfeng Wei, Liguo Zhang, Xiaoying Guo, and Ting Yang. A theoretical and simulation analysis on the power of the frequency domain causality test. Statistics & Probability Letters, 170(??): ??, March 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522030273X.

Wang:2022:QTR

[WZLL22] Lei Wang, Jing Zhang, Bo Li, and Xiaohui Liu. Quantile trace regression via nuclear norm regularization. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002613.

Wang:2025:GWE

[WZQ25] Kang Wang, Na Zou, and Hong Qin. Generalized wordlength enumerator for asymmetrical designs. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224002505.

Wang:2023:SNS

[WZZ23] Jian Wang, Jianliang Zhai, and Jiahui Zhu. The stochastic nonlinear Schrödinger equations driven by pure jump noise. Statistics & Probability Letters, 197(??):??, June 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000342.

Xu:2022:NRB

[XC22a] Jianjun Xu and Wenquan Cui. A new RKHS-based global testing for functional linear model. Statistics & Probability Letters, 182 (??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522100239X.

Xu:2022:HSI

[XC22b] Mingzhou Xu and Kun Cheng. How small are the increments of G-Brownian motion. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522200061X.

Xu:2020:MDN

[XCD20] Mingzhou Xu, Kun Cheng, and Yunzheng Ding. Moderate deviations for nonhomogeneous Markov chains. Statistics & Probability Letters, 157(??):??, February 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302780.

Xu:2022:CNC

[XCDS22] Hui Xu, Joel E. Cohen, Richard A. Davis, and Gennady Samorodnitsky. Cauchy, normal and correlations versus heavy tails. Statistics & Probability Letters, 186(??):??, July 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200075X.

Xu:2022:VCF

[XFJN22] Shuli Xu, Yuqiang Feng, Jun Jiang, and Na Nie. A variation of constant formula for Caputo fractional stochastic differential equations with jump–diffusion. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522200027X.

Xiang:2021:TRH

[XFZ^{+}21] Xuyan Xiang, Haiqin Fu, Jieming Zhou, Yingchun Deng, and Xiangqun Yang. Taboo rate and hitting time distribution of continuous-time reversible Markov chains. Statistics & Probability Letters, 169(??):??, February 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302728.

Xiong:2022:SVM

[XH22] Peihan Xiong and Taizhong Hu. On Samuel’s p-value model and the Simes test under dependence. Statistics & Probability Letters, 187(??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222000864.

Xie:2023:UCR

[Xie23] Haitian Xie. Uniform convergence results for the local linear regression estimation of the conditional distribution. Statistics & Probability Letters, 201(??):??, October 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001104.

Xun:2021:CHG

[XJG21] Li Xun, Renqiao Jiang, and Jianhua Guo. The conditional Haezendonck–Goovaerts risk measure. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220302716.

Xu:2024:CTM

[XK24] Mingzhou Xu and Xuhang Kong. Complete q-th moment convergence of moving average processes for m-widely acceptable random variables under sub-linear expectations. Statistics & Probability Letters, 214(??):??, November 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522400172X.

Xiong:2020:SPD

[XLNQ20] Zikang Xiong, Liwei Liu, Jianhui Ning, and Hong Qin. Sphere packing design for experiments with mixtures. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301103.

Xie:2024:ESS

[XNZL24] Lin Xie, Wenqing Ni, Shuicao Zheng, and Guowei Lei. The existence and smoothness of self-intersection local time for a class of Gaussian processes. Statistics & Probability Letters, 213(??): ??, October 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001597.

Xiong:2021:RNW

[XOC21] Xianzhu Xiong, Meijuan Ou, and Ailian Chen. Reweighted Nadaraya–Watson estimation of conditional density function in the right-censored model. Statistics & Probability Letters, 168(??): ??, January 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302364.

Xiong:2020:AEP

[XP20] Qian Xiong and Zuoxiang Peng. Asymptotic expansions of powered skew-normal extremes. Statistics & Probability Letters, 158 (??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521930313X.

Xu:2023:DII

[Xu23] Jie Xu. A deviation inequality for increment of a G-Brownian motion under G-expectation and applications. Statistics & Probability Letters, 198(??):??, July 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522300072X.

Xie:2020:CPN

[XX20a] Fang Xie and Zhijie Xiao. Consistency of ℓ_{1} penalized negative binomial regressions. Statistics & Probability Letters, 165(??): ??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522030119X.

Xiong:2020:ABS

[XX20b] Yafang Xiong and Xiaoming Xu. Anticipated backward stochastic differential equations with left-Lipschitz coefficient. Statistics & Probability Letters, 163(??):??, August 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300651.

Xiao:2023:HEN

[XXZ23] Xuan Xiao, Xingbai Xu, and Wei Zhong. Huber estimation for the network autoregressive model. Statistics & Probability Letters, 203(??):??, December 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001414.

Xu:2023:FSI

[XY23] Xiaoyan Xu and Xianye Yu. The fractional smoothness of integral functionals driven by Brownian motion. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002309.

Xiang:2020:ECI

[XZ20] Kainan Xiang and Lang Zou. Every countable infinite group admits a long range percolation with a phase transition. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301309.

Xiang:2022:SEB

[XZ22] Kainan Xiang and Lang Zou. Surviving ends in Bernoulli percolation on graphs roughly isometric to a tree. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000086.

Xie:2023:NSD

[XZ23] Jiansheng Xie and Ruisong Zhou. A note on an SLLN of Dvoretzky–Erdős type. Statistics & Probability Letters, 195(??): ??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222002826.

Xin:2020:DHT

[XZYC20] Cheng Xin, Minzhi Zhao, Qiang Yao, and Erjia Cui. On the distribution of the hitting time for the N-urn Ehrenfest model. Statistics & Probability Letters, 157(??):??, February 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302718.

Yakubovich:2021:SPL

[Yak21] Yuri Yakubovich. A simple proof of the Lévy–Khintchine formula for subordinators. Statistics & Probability Letters, 176(??): ??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221000985.

Yang:2024:EQG

[Yan24] Lijian Yang. Exact quantiles of Gaussian process extremes. Statistics & Probability Letters, 213(??):??, October 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001421.

Yang:2025:HSP

[Yan25a] Li Yang. Homogenization for singularly perturbed stochastic wave equations with Hölder continuous coefficients. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002281.

Yang:2025:CGE

[Yan25b] Lijian Yang. Continuity of Gaussian extreme distributions. Statistics & Probability Letters, 216(??):??, January 2025. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002438.

Yang:2025:SMS

[Yan25c] Lijian Yang. Strict monotonicity of stochastic process extreme distributions. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522400261X.

Yosef:2020:KLE

[YB20a] Arthur Yosef and Amos Baranes. Karhunen–Loève expansion of a set indexed fractional Brownian motion. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302755.

Yu:2020:VEG

[YB20b] Qian Yu and Salwa Bajja. Volatility estimation of general Gaussian Ornstein–Uhlenbeck process. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220300997.

Choi:2023:ETE

[yCLjL23] Jin young Choi, Goeun Lee, and Myoung jae Lee. Endogenous treatment effect for any response conditional on control propensity score. Statistics & Probability Letters, 196(??):??, May 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002607.

Yang:2023:SIT

[YDB23] Xiaoran Yang, Junjie Du, and Fangfang Bai. Semiparametric inference of treatment effects on restricted mean survival time in two sample problems from length-biased samples. Statistics & Probability Letters, 193(??):??, February 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002280.

Ye:2024:LIL

[Ye24] Yinna Ye. From law of the iterated logarithm to Zolotarev distance for supercritical branching processes in random environment. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001639.

Yuan:2020:RTW

[YJT20] Hongmin Yuan, Long Jiang, and Dejian Tian. Representation theorems for WVaR with respect to a capacity. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303013.

Yang:2020:IFI

[YJWY20] Yang Yang, Tao Jiang, Kaiyong Wang, and Kam C. Yuen. Interplay of financial and insurance risks in dependent discrete-time risk models. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300559.

Yoon:2020:LBP

[YKK20] Hyungkuk Yoon, Bara Kim, and Jeongsim Kim. Lower bounds on partial sums of expected hitting times. Statistics & Probability Letters, 160(??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300183.

Yuan:2022:PLD

[YL22] Meng Yuan and Dawei Lu. Precise large deviation for sums of sub-exponential claims with the m -dependent semi-Markov type structure. Statistics & Probability Letters, 185(??):??, June 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522200044X.

Yao:2023:CSO

[YLP23] Jia Yao, Bendong Lou, and Xiaoqing Pan. Characterizations of stochastic ordering for non-negative random variables. Statistics & Probability Letters, 203(??):??, December 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001438.

Yang:2022:CGL

[YLW22] Yaxing Yang, Shiqing Ling, and Qiying Wang. Consistency of global LSE for MA(1) models. Statistics & Probability Letters, 182 (??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002546.

Yu:2024:WHO

[YM24] Lin Yu and Xue Mi. Weighted Hardy–Orlicz-amalgam spaces of martingales. Statistics & Probability Letters, 206(??):??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002195.

Yan:2025:MDC

[YMWW25] Kaixin Yan, Ruixing Ming, Haibin Wang, and Wenyuan Wang. On the moments of dividends and capital injections under a variant type of Parisian ruin. Statistics & Probability Letters, 216(??): ??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224001949.

Yin:2024:STL

[YS24] Xiuwei Yin and Guangjun Shen. On the small time large deviation principles of 1D stochastic Landau–Lifshitz–Bloch equation. Statistics & Probability Letters, 207(??):??, April 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002468.

Yin:2020:HIS

[YSZ20] Xiuwei Yin, Guangjun Shen, and Jinhong Zhang. Harnack inequalities for stochastic heat equation with locally unbounded drift. Statistics & Probability Letters, 163(??):??, August 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300936.

Yu:2020:MLE

[Yu20] Qiqing Yu. The maximum likelihood estimator under the exponential distribution with right-censored linear regression data. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301498.

Yu:2022:UPP

[Yu22] Feng Yu. Uniqueness of principal points with respect to porder distance for a class of univariate continuous distribution.

Statistics & Probability Letters, 183(??):??, April 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002881.

Yuan:2021:LDS

[Yua21] Mingao Yuan. Limiting distribution of short cycles in inhomogeneous random uniform hypergraph. Statistics & Probability Letters, 175(??):??, August 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771522100078X.

Yuan:2024:ACR

[Yua24] Mingao Yuan. Asymptotic coarse Ricci curvature of inhomogeneous random graph. Statistics & Probability Letters, 215(??): ??, December 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002141.

Yuan:2023:APS

[YWY23] Mingxia Yuan, Bingjun Wang, and Zhiyan Yang. On the averaging principle for stochastic differential equations driven by G-Lévy process. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000135.

Yu:2020:MAD

[YY20] Lin Yu and Xiao Yu. Mixed atomic decomposition of martingale weak Hardy–Morrey spaces. Statistics & Probability Letters, 164(??):??, September 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715220301073.

Yang:2022:CTN

[YY22] Fen-Fen Yang and Chenggui Yuan. Comparison theorem for neutral stochastic functional differential equations driven by G-Brownian motion. Statistics & Probability Letters, 184(??):??, May 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000165.

Yu:2020:BSD

[YZ20] Xianye Yu and Mingbo Zhang. Backward stochastic differential equations driven by fractional noise with non-Lipschitz coefficients. Statistics & Probability Letters, 159(??):??, April 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521930327X.

Yang:2022:CPD

[YZ22] Qing Yang and Yi Zhang. Change-point detection for the link function in a single-index model. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222000608.

Yan:2023:OFT

[YZ23a] Zhaohui Yan and Shengli Zhao. Optimal fractions of threelevel factorials under a baseline parameterization. Statistics & Probability Letters, 202(??):??, November 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001268.

Yu:2023:MPF

[YZ23b] Huazhen Yu and Lixin Zhang. On the maximum penalized full likelihood approach for Cox model with extreme value for heavily censored survival data. Statistics & Probability Letters, 201(??): ??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223001049.

Zajkowski:2020:BTP

[Zaj20] Krzysztof Zajkowski. Bounds on tail probabilities for quadratic forms in dependent sub-gaussian random variables. Statistics & Probability Letters, 167(??):??, December 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220302017.

Zachariah:2024:NCC

[ZAP24] Swaroop Georgy Zachariah, Mohd. Arshad, and Ashok Kumar Pathak. A new class of copulas having dependence range larger than FGM-type copulas. Statistics & Probability Letters, 206(??):

??, March 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715223002122.

Zhang:2023:TPO

[ZC23] Jichen Zhang and Zengjing Chen. A transitivity property of Ocone martingales. Statistics & Probability Letters, 193(??):??, February 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222002164.

Zhang:2022:CPC

[ZCL22] Tianfang Zhang, Guoqian Cai, and Zhiming Li. Correction to the paper “Construction of some s-level regular designs with general minimum lower-order confounding” [Statist. Probab. Lett. 167 (2020) 108897]. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715222000554. See [LKA20].

Zhang:2021:PBO

[ZCP21] Guoyi Zhang, Ronald Christensen, and John Pesko. Parametric bootstrap and objective Bayesian testing for heteroscedastic oneway ANOVA. Statistics & Probability Letters, 174(??):??, July 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221000572.

Zhou:2024:IRG

[ZDF24] Guangshuo Zhou, Fengjiao Du, and Shengjun Fan. Invariant representation for generators of general time interval quadratic BSDEs under stochastic growth conditions. Statistics & Probability Letters, 205(??):??, February 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223001852.

Zeifman:2021:BRC

[Zei21] Alexander Zeifman. Bounds on the rate of convergence for Markovian queuing models with catastrophes. Statistics & Probability Letters, 176(??):??, September 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221001127.

Zhang:2024:ARB

[ZF24] Yu Zhang and Long Feng. Adaptive rank-based tests for high dimensional mean problems. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001950.

Zhang:2024:RCM

[ZG24] Jing Zhang and Shaojun Guo. Regularized covariance matrix estimation in high dimensional approximate factor models. Statistics & Probability Letters, 207(??):??, April 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002407.

Zou:2020:GFM

[ZGQC20] Na Zou, Tingxun Gou, Hong Qin, and Kashinath Chatterjee. Generalized foldover method for high-level designs. Statistics & Probability Letters, 164(??):??, September 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300985.

Zhong:2021:ETE

[ZGZF21] Wei Zhong, Yang Gao, Wei Zhou, and Qingliang Fan. Endogenous treatment effect estimation using high-dimensional instruments and double selection. Statistics & Probability Letters, 169(??): ??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302704.

Zhang:2022:QCR

[ZH22] Xiaoyue Zhang and Wenming Hong. Quenched convergence rates for a supercritical branching process in a random environment. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002418.

Zhao:2024:RDS

[ZH24] Huiyan Zhao and Yupei Huang. Random Dirichlet series with α-stable coefficients. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000166.

Zhang:2020:CDW

[Zha20a] Tonglin Zhang. CP decomposition and weighted clique problem. Statistics & Probability Letters, 161(??):??, June 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300262.

Zhao:2020:NNB

[Zha20b] Yunpeng Zhao. A note on new Bernstein-type inequalities for the log-likelihood function of Bernoulli variables. Statistics & Probability Letters, 163(??):??, August 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220300821.

Zhang:2021:SFP

[Zha21a] Hua Zhang. Strong Feller property for one-dimensional Lévy processes driven stochastic differential equations with Hölder continuous coefficients. Statistics & Probability Letters, 169(??):??, February 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302777.

Zhang:2021:HTU

[Zha21b] Li-Xin Zhang. Heyde’s theorem under the sub-linear expectations. Statistics & Probability Letters, 170(??):??, March 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522030290X.

Zhang:2022:MMN

[Zha22a] Li-Xin Zhang. The moments of the maximum of normalized partial sums related to laws of the iterated logarithm under the sublinear expectation. Statistics & Probability Letters, 188(??):??, September 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715222001067.

Zhang:2022:WJM

[Zha22b] Rui-Ray Zhang. When Janson meets McDiarmid: Bounded difference inequalities under graph-dependence. Statistics & Probability Letters, 181(??):??, February 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002340.

Zhao:2022:LTB

[Zha22c] Linjie Zhao. Long-time behavior of SSEP with slow boundary. Statistics & Probability Letters, 185(??):??, June 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000530.

Zhang:2023:AND

[Zha23] Qingyang Zhang. On the asymptotic null distribution of the symmetrized Chatterjee’s correlation coefficient. Statistics & Probability Letters, 194(??):??, March 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222002723.

Zhang:2024:NDC

[Zha24] Qiang Zhang. A novel dual-criterion framework for change point detection. Statistics & Probability Letters, 211(??):??, August 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224001019.

Zhang:2025:SIO

[Zha25] Dingwen Zhang. Statistical inference for Ornstein–Uhlenbeck processes based on low-frequency observations. Statistics & Probability Letters, 216(??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002554.

Zhang:2024:NEP

[ZHD24] Xuekang Zhang, Chengzhe Huang, and Shounian Deng. Nonparametric estimation for periodic stochastic differential equations driven by G-Brownian motion. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001718.

Zhang:2021:SBK

[ZJ21] Wei Zhang and Long Jiang. Solutions of BSDEs with a kind of non-Lipschitz coefficients driven by G-Brownian motion. Statistics & Probability Letters, 171(??):??, April 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220303278.

Zhuo:2020:TSC

[ZJD20] Bin Zhuo, Duo Jiang, and Yanming Di. Test-statistic correlation and data-row correlation. Statistics & Probability Letters, 167(??): ??, December 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220302066.

Zeng:2022:BCS

[ZK22] Xiaoqiang Zeng and Yoshihide Kakizawa. Bias-correction of some estimators in the INAR(1) process. Statistics & Probability Letters, 187(??):??, August 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715222000839.

Zeng:2024:TSC

[ZK24] Xiaoqiang Zeng and Yoshihide Kakizawa. Two-step conditional least squares estimation in ADCINAR(1) process, revisited. Statistics & Probability Letters, 206(??):??, March 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223002274.

Zhao:2020:LDR

[ZL20] Shoujiang Zhao and Qiaojing Liu. A large deviation result for maximum likelihood estimator of non-homogeneous Ornstein– Uhlenbeck processes. Statistics & Probability Letters, 162(??):??, July 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300560.

Zhou:2025:FET

[ZL25] Yinbing Zhou and Dawei Lu. The first exit time of fractional Brownian motion from an unbounded domain. Statistics & Probability Letters, 218(??):??, March 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224002888.

Zhang:2023:PCI

[ZLMT23] Jeffrey Zhang, Wei Li, Wang Miao, and Eric Tchetgen Tchetgen. Proximal causal inference without uniqueness assumptions. Statistics & Probability Letters, 198(??):??, July 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0167715223000603. See correction [ZLMT24].

Zhang:2024:CPC

[ZLMT24] Jeffrey Zhang, Wei Li, Wang Miao, and Eric Tchetgen Tchetgen. Corrigendum to “Proximal causal inference without uniqueness assumptions” [Statistics and Probability Letters 198 (2023) 1–8/109836]. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000245. See [ZLMT23].

Zhao:2024:CTS

[ZLS24] Jiayang Zhao, Jie Liu, and Yuting Su. Consistent two-stage estimation in heterogeneous network autoregressive model. Statistics & Probability Letters, 212(??):??, September 2024. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715224001160.

Zhang:2020:AJE

[ZLWZ20] Xiuzhen Zhang, Zhiping Lu, Yangye Wang, and Riquan Zhang. Adjusted jackknife empirical likelihood for stationary ARMA and ARFIMA models. Statistics & Probability Letters, 165(??):??, October 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301334.

Zheng:2020:ISD

[ZLZK20] Zemin Zheng, Liwan Li, Jia Zhou, and Yinfei Kong. Innovated scalable dynamic learning for time-varying graphical models. Statistics & Probability Letters, 165(??):??, October 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715220301462.

Zhang:2023:DAQ

[ZM23] Hanjun Zhang and Yongxiang Mo. Domain of attraction of quasi-stationary distribution for absorbing Markov processes. Statistics & Probability Letters, 192(??):??, January 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522200205X.

Zou:2023:RMT

[Zou23] Nan Zou. Reweighted madogram-type estimator of Pickands dependence function. Statistics & Probability Letters, 195(??):??, April 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000147.

Zhou:2020:EFP

[ZP20]  Jianjun Zhou and Qingyan Peng.  Estimation for func-

tional partial linear models with missing responses. Statistics & Probability Letters, 156(??):??, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219302445.

Zhang:2021:MEC

[ZPC21] Qihu Zhang, Cheolwoo Park, and Jongik Chung. Minimax estimation of covariance and precision matrices for high-dimensional time series with long-memory. Statistics & Probability Letters, 177(??): ??, October 2021. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221001395.

Zhang:2024:NCD

[ZS24] Jialin Zhang and Jingyi Shi. Nonparametric clustering of discrete probability distributions with generalized Shannon’s entropy and heatmap. Statistics & Probability Letters, 208(??):??, May 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224000397.

Zeifman:2020:OSB

[ZSK20] A. I. Zeifman, Y. A. Satin, and K. M. Kiseleva. On obtaining sharp bounds of the rate of convergence for a class of continuoustime Markov chains. Statistics & Probability Letters, 161(??): ??, June 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771522030033X.

Zuo:2020:LSP

[Zuo20] Yijun Zuo. Large sample properties of the regression depth induced median. Statistics & Probability Letters, 166(??):??, November 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220301826.

Zhang:2023:MPM

[ZW23] Pengcheng Zhang and Xueyuan Wu. Multivariate Poisson model adjusting for unidirectional covariate misrepresentation. Statistics & Probability Letters, 197(??):??, June 2023. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715223000615.

Zhou:2022:GWP

[ZWW22] Guoli Zhou, Lidan Wang, and Jiang-Lun Wu. Global wellposedness of 2D stochastic Burgers equations with multiplicative noise. Statistics & Probability Letters, 182(??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715221002741.

Zhao:2020:SFT

[ZX20] Huiyan Zhao and Siyan Xu. A stochastic Fubini theorem for αstable process. Statistics & Probability Letters, 160(??):??, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715220300031.

Zhao:2021:FPT

[ZX21] Zhenwen Zhao and Yuejuan Xi. The first passage time on the (reflected) Brownian motion with broken drift hitting a random boundary. Statistics & Probability Letters, 171(??):??, April 2021. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771522100002X.

Zhang:2025:MHP

[ZX25] Zhehao Zhang and Ruina Xing. Multivariate Hawkes process allowing for common shocks. Statistics & Probability Letters, 216 (??):??, January 2025. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224002396.

Zhang:2020:NHD

[ZY20] Yue Zhang and Mingao Yuan. Nonreconstruction of highdimensional stochastic block model with bounded degree. Statistics & Probability Letters, 158(??):??, March 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715219303219.

Zhang:2023:CMS

[ZYL23] Bin Zhang, Zhigang Yao, and Junfeng Liu. On a class of mixed stochastic heat equations driven by spatially homogeneous Gaussian noise. Statistics & Probability Letters, 196(??):??, May 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223000317.

Zhou:2022:CSC

[ZYLB22] S. Zhou, K. Yao, W. Liu, and F. Bretz. Construction of simultaneous confidence bands using conditional Monte Carlo. Statistics & Probability Letters, 182(??):??, March 2022. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715221002807.

Zhang:2020:PEN

[ZYS20] Xuekang Zhang, Haoran Yi, and Huisheng Shu. Parameter estimation for non-stationary reflected Ornstein–Uhlenbeck processes driven by α-stable noises. Statistics & Probability Letters, 156(??):

??, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302639.

Zhang:2022:CST

[ZYY22] Zhengcheng Zhang, Luyi Yang, and Yonghong Yang. The conditional spacings and their stochastic properties. Statistics & Probability Letters, 186(??):??, July 2022. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715222000591.

Zhang:2024:NTI

[ZZ24a] Jialin Zhang and Zhiyi Zhang. A normal test for independence via generalized mutual information. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000828.

Zhang:2024:SDR

[ZZ24b] Yawen Zhang and Caibin Zhang. Stackelberg differential reinsurance and investment game for a dependent risk model with Ornstein–Uhlenbeck process. Statistics & Probability Letters, 214(??):??, November 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715224001925.

Zhang:2024:ERI

[ZZ24c] Yuhui Zhang and Yi Zhao. Explicit representation of invariant measures for 2-death processes. Statistics & Probability Letters, 210(??):??, July 2024. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715224000890.

Zhao:2024:LDS

[ZZ24d] Yinxuan Zhao and Mei Zhang. Large deviations for sums associated with supercritical branching process in a random environment. Statistics & Probability Letters, 207(??):??, April 2024. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715223002420.

Zhang:2025:PMI

[ZZ25] Chi Zhang and Danna Zhang. Probability and moment inequalities for quadratic forms in independent random variables with fat tails. Statistics & Probability Letters, 217(??):??, February 2025. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715224002591.

Zhang:2022:RST

[ZZF22] Xiaoxu Zhang, Ping Zhao, and Long Feng. Robust sphericity test in the panel data model. Statistics & Probability Letters, 182 (??):??, March 2022. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715221002650.

Zhou:2021:ISE

[ZZZD21] Jia Zhou, Zemin Zheng, Huiting Zhou, and Ruipeng Dong. Innovated scalable efficient inference for ultra-large graphical models. Statistics & Probability Letters, 173(??):??, June 2021. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771522100047X.